### v2

# Submitted toTopology Proceedings

Abstract.
In this follow-up paper to [10] we investigate the global topology and geometry of dynamical systems $\dot{x}=F(x)$ with entire vector field $F$ by exploiting and extending the local structure of finite elliptic decompositions. We prove topological properties of the basins of centers, nodes, and foci, while excluding isolated equilibria at the boundaries of the latter two. We propose a definition of global elliptic sectors and introduce the notion of a sector-forming orbit allowing global topological analysis. The structure of heteroclinic regions between two equilibria is characterized.
Key words and phrases:
Holomorphic dynamical system, global phase portrait, basin of attraction/stability, global elliptic sector, heteroclinic region
2020 Mathematics Subject Classification:
Primary 30A99, 30C10, 30C15, 30D30; Secondary 32M25, 37F10, 37F75
Acknowledgement: We thank Jörn Dietrich for mathematical discussions and collaborative solution-finding.
1. Introduction
In our previous paper [10], we focused on the local topological properties of both simple and higher-order equilibria to set the stage for investigating the global phase portrait of holomorphic vector fields in this follow-up paper. In order to reveal the relation between local and global dynamics, we utilize the geometric structure of finite elliptic decompositions to explore the "stability regions" of equilibria, global elliptic sectors, and heteroclinic regions. This is the next step towards studying holomorphic and meromorphic flows of vector fields derived from complex functions pertinent to number theory, such as the Riemann $\zeta$- and $\xi$-function, cf. [5, 6, 11, 21]. The boundary-forming orbits (separatrices) of these "stability regions" are particularly interesting since they divide the complex plane into regions that exhibit completely distinct topological behavior, reminiscent of phase transitions. In previous works, we have been analyzing normally attracting/repelling (slow) invariant manifolds of equilibria in the context of multi scale dynamical systems, cf. [8, 12, 13, 14, 15, 16, 17], which, in the 2-dimensional planar case, are usually separatrices dividing the phase space into topological basins.
In the following, we provide an outline of our paper. We consider the special case of entire vector fields in real time parameterization, i.e. a holomorphic dynamical system of the form
(1.1) $\dot{x}=\frac{\mathrm{d}x}{\mathrm{d}t}=F(x),\quad x\in\mathbb{C}{},\;t\in\mathbb{R}$
with $F\in\mathcal{O}(\mathbb{C}{})$. [^1] In the context of dynamical systems, we use the notation of orbits and flows given in [10, Chapter 2]. A definition of limit sets, cf. [2, §4] and [20, Chapter 8.4], as well as some basic results can also be found in [10, Chapter 2]. We start with the analysis of the basin of a center, cf. [4, Theorem 4.1] and [3, Theorem 3.3], and provide detailed proofs of several topological properties. In particular, we detail the proof in [3, Theorem 3.3] for the openness of the basin of a center.
Subsequently, we examine the topological structure of equilibria of order $m\geq 2$. Since a precise definition of global elliptic sectors is missing in the literature222cf. [2, Chapter VIII, §17], [4, Chapter 4], [7, Chapter 1.5] and [19, Chapter 2.11]., we propose an appropriate characterization by introducing the concept of sector-forming orbits. We demonstrate that the intuitive idea of the geometric structure of a global elliptic sector is indeed rigorously valid. We support the proof of the openness with geometric visualization, cf. Figure 1, and point out the conditions under which the local structure of a finite elliptic decomposition can be transferred to the global phase portrait, an issue already addressed in [10, Example 4.7, Example 4.8].
Building on these results, we analyze the topology of the basin of nodes and foci. Notably, the authors of [4, Theorem 4.3] explicitly do not exclude the scenario of an isolated equilibrium on the boundary of the basin – specifically, an equilibrium that is not attached to an orbit lying outside the basin. We address this aspect in detail in Proposition 4.5.
Finally, we define and investigate heteroclinic regions between two equilibria, observing properties similar to the preceding cases. This leads to the exclusion of the existence of a cycle in the phase space formed by equilibria and heteroclinic orbits connecting these equilibria, cf. Corollary 5.7.
2. Basin of centers
In this chapter, we treat the case of a center. We provide more detailed complete proofs of topological results given in [4, Theorem 4.1] and [3, Theorem 3.3]. In all considerations of this chapter we consider an entire function $F\in\mathcal{O}{}(\mathbb{C}{})$, $F\not\equiv 0$, and a center $a\in\mathbb{C}{}$ of (1.1).
Definition 2.1.
We define the basin $\mathcal{V}$ of $F$ in $a$ as
$\mathcal{V}:=\{a\}\cup\left\{x\in\mathbb{C}:\Gamma(x)\text{ is periodic with }a\in\operatorname{Int}(\Gamma)\right\}.$
The boundary of $a$ is defined as the boundary of $\mathcal{V}$.
Proposition 2.2.
$\mathring{\mathcal{V}}$ is nonempty. $\mathcal{V}$ and $\partial\mathcal{V}$ are flow-invariant. It holds
(2.1) $\mathcal{V}=\bigcup_{\mathclap{x\in\mathcal{V}}}\overline{\operatorname{Int}(\Gamma(x))}\text{.}$
Proof.
[10, Definition 3.1] ensures $a\in\mathring{\mathcal{V}}$, i.e. $\mathring{\mathcal{V}}\not=\emptyset$. Let $y\in\mathcal{V}$ and $z\in\Gamma(y)$. Then $\Gamma(y)=\Gamma(z)\subset\mathcal{V}$. Thus $\mathcal{V}$ is flow-invariant. Every $y\in\operatorname{Int}(\Gamma(x))$ either satisfies $y=a$, or, by [10, Theorem 5.1], is a periodic orbit with $a$ in its interior. This proves equation (2.1). By [22, Lemma 6.4], it follows that also $\partial\mathcal{V}$ is flow-invariant. ∎
Proposition 2.3.
It holds $\partial\mathcal{V}\cap F^{-1}(\{0\})=\emptyset$, i.e. there are no equilibria on the boundary of $\mathcal{V}$.
Proof.
The case $\partial\mathcal{V}=\emptyset$ is trivial, so let $x\in\partial\mathcal{V}$ and assume this is an equilibrium. For all $\varepsilon>0$ there exists $y\in\mathcal{B}_{\varepsilon}(x)\cap\mathcal{V}$ such that $\Gamma(y)$ neither has only $x$ in its limit sets nor is a periodic orbit with $x$ in its interior. Note that $x\not=a$, since $a\in\mathring{\mathcal{V}}$. Hence, $x$ cannot be a center, focus or node. By [10, Theorem 4.4], $x$ has a FED. In particular, by [10, Definition 4.1], there exists $r>0$ such that $x\in w_{+}(\Gamma(z))\cup w_{-}(\Gamma(z))$ for all $z\in\mathcal{B}_{r}(x)$. But no orbits in $\mathcal{B}_{r}(x)\cap\mathcal{V}\not=\emptyset$ tend to $x$, i.e. this is a contradiction to [10, Definition 4.1 c) (ii)]. Thus, $x$ cannot be an equilibrium. ∎
Theorem 2.4.
$\mathcal{V}$ is open.
Proof.
Suppose there exists $x\in\mathcal{V}\cap\partial\mathcal{V}\not=\emptyset$, $x\not=a$. Then $\Gamma(x)\subset\mathcal{V}\cap\partial\mathcal{V}$, cf. Proposition 2.2. In particular, $\Gamma(x)$ is a periodic orbit with $a\in\operatorname{Int}(\Gamma(x))\not=\emptyset$. Suppose there also exists $y\in\partial\mathcal{V}\setminus\Gamma(x)\not=\emptyset$. The set $\mathbb{C}\setminus\Gamma(x)=\operatorname{Int}(\Gamma(x))\cup\operatorname{Ext}(\Gamma(x))$ is open. If $y\in\operatorname{Int}(\Gamma(x))$, there exists $r>0$ such that $\mathcal{B}_{r}(y)\subset\operatorname{Int}(\Gamma(x))$. By equation (2.1), we get $\mathcal{B}_{r}(y)\subset\mathcal{V}$, i.e. $y\not\in\partial\mathcal{V}$. This a contradiction. Furthermore, if $y\in\operatorname{Ext}(\Gamma(x))$, there exists again an $r>0$ such that $\mathcal{B}_{r}(y)\subset\operatorname{Ext}(\Gamma(x))$. Since $y$ lies on the boundary of $a$, there exists $z\in\mathcal{B}_{r}(y)\cap\mathcal{V}$, i.e. $\Gamma(z)$ is a periodic orbit with $a\in\operatorname{Int}(\Gamma(z))$ and $\Gamma(z)\not=\Gamma(x)$. Since $\Gamma(z)$ is path-connected and $z\in\operatorname{Ext}(\Gamma(x))$, we must have $x\in\operatorname{Int}(\Gamma(z))$. By equation (2.1) and openness of $\operatorname{Int}(\Gamma(z))$, we derive the contradiction $x\in\mathring{\mathcal{V}}$. Thus, $y\not\in\operatorname{Int}(\Gamma(x))\cup\operatorname{Ext}(\Gamma(x))$, which is also a contradiction. We get $\partial\mathcal{V}\setminus\Gamma(x)=\emptyset$, i.e. the boundary of $a$ consists exactly of the unique outermost periodic orbit $\Gamma(x)$ and $\mathcal{V}=\overline{\operatorname{Int}(\Gamma(x))}$ must be compact.
By equation (2.1), $\Gamma(x)$ is not isolated. Hence we can apply a result on limit cycles of analytic dynamical systems, [2, p. 203, §12.1, Lemma 1], to deduce the existence of $\rho>0$ such that all orbits through the set $U:=\{y\in\mathbb{C}{}:\text{dist}(\Gamma(x),y)<\rho\}\cap\operatorname{Ext}(\Gamma(x))$ are periodic. Reduce $\rho$ such that $\overline{\operatorname{Int}(\Gamma(x))}\cup U$ is simply connected and $U\cap F^{-1}(\{0\})=\emptyset$. By continuity of the flow, cf. [19, Chapter 2.4, Theorem 4], we deduce the existence of at least one periodic orbit $\Gamma\subset U$. By choice of $\rho$, we get $\operatorname{Int}(\Gamma)\subset U\cup\overline{\operatorname{Int}(\Gamma(x))}$ and thus $a\in\operatorname{Int}(\Gamma)$, cf. [2, §11, Corollary 1] and [3, Theorem 2.4]. This is a contradiction to $\Gamma\subset\mathbb{C}{}\setminus\mathcal{V}$. Hence the assumed $x$ cannot exist and $\mathcal{V}$ is open. ∎
Theorem 2.5.
$\mathcal{V}$ is connected, path-connected and simply connected.
Proof.
The interior of any periodic orbit is clearly connected. Hence, by Proposition 2.2, $\mathcal{V}$ can be interpreted as union of connected sets, all having the point $a$ in common. Thus, also the union must be connected, cf. [18, Theorem 23.3]. By [23, Proposition 12.25], $\mathcal{V}$ is even path-connected.
Let $L\subset\mathcal{V}$ be a loop (closed path). By openness of $\mathcal{V}$, cf. Theorem 2.4, for all $x\in L$ there exists a $y_{x}\in\operatorname{Ext}(\Gamma(x))\cap\mathcal{V}\not=\emptyset$ such that $x\in\operatorname{Int}(\Gamma(y_{x}))\subset\mathcal{V}$. Hence, the system $\left\{\operatorname{Int}(\Gamma(y_{x})):x\in L\right\}$ is an open cover of $L$. Since $L$ is compact, there exist finitely many points $x_{1},\ldots,x_{\ell}\in L$, $\ell\in\mathbb{N}{}$, such that
$L\subset\bigcup_{k=1}^{\ell}\operatorname{Int}(\Gamma(y_{x_{k}}))\subset\mathcal{V}\text{.}$
Since $\ell$ is finite, $y\in\left\{y_{x_{k}}:1\leq k\leq\ell\right\}$ must exist such that $\Gamma(y)$ is the outermost periodic orbit having the other $\ell-1$ periodic orbits in its interior. By equation (2.1), we conclude $L\subset\operatorname{Int}(\Gamma(y))\subset\mathcal{V}$ with $y\in\mathcal{V}$. Since $\operatorname{Int}(\Gamma(y))$ is clearly simply connected, $L$ can be continuously (homotopically) deformed to a point in $L\subset\mathcal{V}$. Since $\mathcal{V}$ is already path-connected, it follows that $\mathcal{V}$ is simply connected. ∎
Theorem 2.6.
All orbits on the boundary of $a$ are unbounded. $\mathcal{V}$ is unbounded.
Proof.
If the boundary of $a$ is empty, then $\mathcal{V}$ is open, closed and nonempty, i.e. $\mathcal{V}=\mathbb{C}{}$ and thus unbounded. Assume $\partial\mathcal{V}\not=\emptyset$ and let $x\in\partial\mathcal{V}$. Suppose that $\Gamma(x)\subset\partial\mathcal{V}$ is bounded. By Proposition 2.3, we get $w_{\pm}(\Gamma(x))\cap F^{-1}(\{0\})=\emptyset$. Hence, by [10, Corollary 5.2], $\Gamma(x)$ must be periodic. By [10, Theorem 5.1], $\Gamma(x)$ has exactly one equilibrium, a center $\tilde{a}$, in $\operatorname{Int}(\Gamma(x))$. If $\tilde{a}\not=a$, $x\in\partial\mathcal{V}$ would be a point in the basin in $\tilde{a}$, which is a contradiction to the openness of this basin, cf. Theorem 2.4. If $a=\tilde{a}$, we get $x\in\mathcal{V}$, which is again a contradiction, since $\mathcal{V}\cap\partial\mathcal{V}=\emptyset$. Thus, $\Gamma(x)$ and $\mathcal{V}$ must be unbounded. ∎
3. Global elliptic sectors
In all claims and arguments of this chapter, we consider an entire function $F\in\mathcal{O}{}(\mathbb{C})$, $F\not\equiv 0$, and an equilibrium $a\in\mathbb{C}$ of (1.1) of order $m\in\mathbb{N}\setminus\{1\}$. By applying [10, Proposition 4.3] and [10, Theorem 4.4], (1.1) has a finite elliptic decomposition (FED) of order $22m-$ 2 in $a$, where all local elliptic sectors have adjacent definite directions given by $\mathcal{E}(F,m)$. Moreover, we use the function $\lambda:\mathcal{E}(F,m)\to\{-1,1\}$, $\lambda(\theta):=\cos(\operatorname{arg}(F^{(m)}(a))+\theta(m-1))$ to determine the time direction (positive or negative) in which the orbits tend to $a$ in the FED, cf. [10, Proposition 4.3 (iii)] and [10, calculation of $\lambda_{i}$ on p. 11].
In the following, we will extend the analysis of the local structure described in [10, Definition 4.1 c)] to a global point of view and provide detailed complete proofs of the topological results given in [4, Theorem 4.2]. To do this, we need at first a precise definition of a global elliptic sector.
Definition 3.1.
  • (i)
    Let $\Gamma\subset\mathbb{C}$ be an orbit and $x\in\Gamma$. We call $\Gamma$ $(\theta_{+},\theta_{-})$-elliptic in $a$, if there exists $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ such that $w_{+}(\Gamma)=w_{-}(\Gamma)=\{a\}$, $\arg(\Phi(t,x)-a)=\theta_{+}$ for $t\to\infty$ and $\arg(\Phi(t,x)-a)=\theta_{-}$ for $t\to-\infty$, i.e. $\Gamma$ is homoclinic and tends to $a$ in in the definite direction $\theta_{\pm}$ for $t\to\pm\infty$.
  • (ii)
    Let $\Xi\subset\mathbb{C}$ be an orbit, which is $(\theta_{+},\theta_{-})$-elliptic in $a$. We call $\Xi$ a sector-forming orbit in $a$ with definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, if for all $y\in\operatorname{Int}(\Xi\cup\{a\})$ the orbit $\Gamma(y)$ is also $(\theta_{+},\theta_{-})$-elliptic in $a$.
  • (iii)
[…]
    $\mathcal{S}(\theta_{+},\theta_{-},\Xi):=\underbrace{\Xi\cup\operatorname{Int}(\Xi\cup\{a\})}_{=\overline{\operatorname{Int}(\Xi\cup\{a\})}\setminus\{a\}}\cup\,\mathcal{S}^{\prime}(\theta_{+},\theta_{-},\Xi)$
    with
    $\mathcal{S}^{\prime}(\theta_{+},\theta_{-},\Xi):=\big{\{}x\in\mathbb{C}:\,$ $\Gamma(x)\text{ is }(\theta_{+},\theta_{-})\text{-elliptic in }a,$
    $\Xi\subset\operatorname{Int}(\Gamma(x)\cup\{a\})\big{\}}\text{.}$
Remark 3.2.
Let $\Gamma\subset\mathbb{C}$ be $(\theta_{+},\theta_{-})$-elliptic in $a$. A construction of a parameterization for the closed Jordan curve $\Gamma\cup\{a\}$ with compact time interval can be found in [10, Remark 4.2]. Hence, all sets in Definition 3.1 are indeed well-defined.
The global elliptic sector in the constructive Definition 3.1 (iii) depends on the choice of the sector-forming orbit. We will later see the motivation for this type of construction and demonstrate that the global elliptic sector is in an appropriate sense independent of the choice of a particular sector-forming orbit – as can be intuitively deduced from the idea of the structure of an equilibrium of order $m\geq 2$, cf. [10, Figure 2, Figure 3].
Lemma 3.3.
Let $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ be two adjacent definite directions with $\lambda(\theta_{\pm})=\mp 1$. Then there exists a sector-forming orbit $\Xi$ in $a$ with definite directions $\theta_{+},\theta_{-}$.
Proof.
By [10, Theorem 4.4], there exists a local elliptic sector $S$ between $\theta_{+}$ and $\theta_{-}$. Moreover, the two characteristic orbits of $S$ tend to $a$ in the directions $\theta_{+}$ and $\theta_{-}$, cf. [10, Definition 4.1 c)] and [10, Proposition 4.3 (ii)]. Hence, we can choose the orbit $\Xi:=\Gamma(E_{1})\subset S$, where $E_{1}\in S$ is the point given by [10, Definition 4.1 c)]. In fact, all orbits in $\overline{\operatorname{Int}(\Xi\cup\{a\})}\setminus\{a\}\subset S$ lie between the two characteristic orbits of $S$ and thus are $(\theta_{+},\theta_{-})$-elliptic in $a$, cf. [10, Definition 4.1 c) (ii), (iv)]. ∎
The above Lemma ensures the existence of a sector-forming orbit for adjacent definite directions. In the following propositions and theorems we use this construction to show various topological properties of global elliptic sectors such as invariance, openness, unboundedness, connectedness.
Proposition 3.4.
Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$. $\mathring{\mathcal{S}}$ is nonempty. $\mathcal{S}$ and $\partial\mathcal{S}$ are flow-invariant. It holds
$\mathcal{S}=\bigcup_{\mathclap{x\in\mathcal{S}}}\Gamma(x)\text{.}$
Proof.
By choosing a point in $\operatorname{Int}(\Xi\cup\{a\})$, we have $\mathring{\mathcal{S}}\not=\emptyset$. The rest of the proof is analogous to the proof of Proposition 2.2. ∎
Proposition 3.5.
Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$. It holds $\partial\mathcal{S}\cap F^{-1}(\{0\})=\{a\}$.
Proof.
Let $x\in\partial\mathcal{S}\setminus\{a\}$. Suppose $x$ is an equilibrium. Since the orbits in $\mathcal{S}$ all tend to $a\not=x$ in both time directions and are not periodic, $x$ cannot be a center, focus or node. Hence, as in the proof of Proposition 2.3, $x$ has a FED, cf. [10, Theorem 4.4]. Moreover, by [10, Definition 4.1], there exists $r>0$ such that $a\in w_{+}(\Gamma(x))\cup w_{-}(\Gamma(x))$ for all $x\in\mathcal{B}_{r}(a)$. In particular, we get $\mathcal{B}_{r}(a)\cap\mathcal{S}\not=\emptyset$, which is a contradiction. Thus, $x$ cannot be an equilibrium. This proves $\partial\mathcal{S}\cap F^{-1}(\{0\})\subset\{a\}$. Moreover, $\Xi\subset\mathcal{S}$ tends to $a$, i.e. $a\in\overline{\mathcal{S}}$. Since $a\not\in\mathcal{S}$, we conclude $a\in\overline{\mathcal{S}}\setminus\mathcal{S}\subset\partial\mathcal{S}$ and thus $\partial\mathcal{S}\cap F^{-1}(\{0\})=\{a\}$. ∎
Theorem 3.6.
Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$. Define $\mathcal{S}^{\prime}:=\mathcal{S}^{\prime}(\theta_{+},\theta_{-},\Xi)$. Then $\mathcal{S}$ is open. In particular, for all $x\in\mathcal{S}$ there exists $y\in\mathcal{S}\cap\operatorname{Ext}(\Gamma(x)\cup\{a\})$ such that $\Gamma(x)\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$.
Proof.
Let $x\in\mathcal{S}$, i.e. $\Gamma(x)$ is $(\theta_{+},\theta_{-})$-elliptic in $a$. If $x\in\operatorname{Int}(\Xi\cup\{a\})$, we can choose $y\in\Xi$ to ensure $\Gamma(x)\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$. Moreover, $\operatorname{Int}(\Xi\cup\{a\})$ is open, i.e. nothing has to be shown. Hence, we can assume $x\in\Xi\cup\mathcal{S}^{\prime}$.
Step 1: Application of [10, Proposition 4.3]
By [10, Proposition 4.3 (iii)], there exists $r,\Theta_{\pm}>0$ such that for all $z_{\pm}\in\{z\in\mathbb{C}{}:|z-a|<r,|\arg(z-a)-\theta_{\pm}|<\Theta_{\pm}\}=:A_{\pm}$ the orbit through $z_{\pm}$ tends to $a$ in the definite direction $\theta_{\pm}$ for $t\to\pm\infty$. Moreover, there exists $\tau>0$ such that $M_{\pm}:=\Phi(\pm\tau,x)\in A_{\pm}$. Choose $\varepsilon_{\pm}>0$ small enough such that $\mathcal{B}_{\varepsilon_{\pm}}(M_{\pm})\subset A_{\pm}$. Define $\varepsilon:=\min\{\varepsilon_{+},\varepsilon_{-}\}$. By continuity of the flow, cf. [19, Chapter 2.4, Theorem 4], there exists a $\delta_{1}>0$ such that $|\Phi(t,x)-\Phi(t,\hat{x})|<\varepsilon$ for all $\hat{x}\in\mathcal{B}_{\delta_{1}}(x)$ and $t\in[-\tau,\tau]$. This shows that for all $\hat{x}\in\mathcal{B}_{\delta_{1}}(x)$ the orbit $\Gamma(\hat{x})$ is $(\theta_{+},\theta_{-})$-elliptic in $a$, cf. [10, Proposition 4.3 (iii)].
Step 2: Construction of an elementary topological quadrangle
We choose two paths without contact $l_{\pm}\subset\mathcal{B}_{\varepsilon}(M_{\pm})$ through $M_{\pm}$, cf. [2, §3], parameterized by ±$\ell_{\pm}:[q_{\pm},Q_{\pm}]\to l_{\pm}$ with $q_{\pm}<0<Q_{\pm}$ and $M_{\pm}=\ell_{\pm}(0)$. We denote the curve segment from $M_{-}$ to $M_{+}$ by $L_{0}$. By [2, §3, 6., Lemma 9 and proof of Lemma 10] and [2, §3, 5., Lemma 8], there exists $\sigma\in(0,\min\{|q_{-}|,Q_{-}\})$ as well as two continuously differentiable functions $\chi_{1}:[-\sigma,\sigma]\to(0,\infty)$ and $\chi_{2}:[-\sigma,\sigma]\to[q_{+},Q_{+}]$ with the following properties.
  • (A)
    $\chi_{2}$ is strictly monotonically increasing.
  • (B)
    For all $s\in[-\sigma,\sigma]$ the orbit through $\ell_{-}(s)\in l_{-}$ crosses $l_{+}$ at the point $\ell_{+}(\chi_{2}(s))$ in the time $\chi_{1}(s)$, i.e. $\Gamma(\ell_{-}(s))=\Gamma(\ell_{+}(\chi_{2}(s)))$ and
    $\quad\quad\quad l_{+}\cap\Phi([0,\chi_{1}(s)],\ell_{-}(s))=\{\Phi(\chi_{1}(s),\ell_{-}(s))\}=\{\ell_{+}(\chi_{2}(s))\}\text{.}$
Define the shorter paths without contact $\hat{l}_{-}:=\ell_{-}([-\sigma,\sigma])$ and $\hat{l}_{+}:=\ell_{+}([\chi_{2}(-\sigma),\chi_{2}(\sigma)])$ as well as the compact sets
$R$ $:=\;\bigcup_{\mathclap{s\in[-\sigma,\sigma]}}\;\Phi([0,\chi_{1}(s)],\ell_{-}(s))\text{,}$
$\hat{R}$ :=$:=\left\{(s,t)\in\mathbb{R}^{2}:-\sigma\leq s\leq\sigma,\,0\leq t\leq\chi_{1}(s)\right\}\text{.}$
  • (C)
    The function $\zeta:\hat{R}\to R$, $\zeta(s,t):=\Phi(t,\ell_{-}(s))$, is a homeomorphism, i.e. $R\subset\mathbb{C}{}$ is an elementary topological quadrangle around $L_{0}$, cf. [2, §3, p. 69] and [2, §3, 5. Figure 37].
  • (D)
    For all $s\in[-\sigma,\sigma]$ the set $\zeta(s,[0,\chi_{1}(s)])$ is the part of the orbit $\Gamma(\ell_{-}(s))$ from $\ell_{-}(s)$ to $\ell_{+}(\chi_{2}(s))$.
  • (E)
    It holds
    $\partial R=\hat{l}_{+}\cup\hat{l}_{-}\cup\zeta(\sigma,[0,\chi_{1}(\sigma)])\cup\zeta(-\sigma,[0,\chi_{1}(-\sigma)])\text{.}$
Step 3: Impossibility of mutually exterior orbits (IMEO) in $R$
There holds the following auxiliary result:
  1. (IMEO)
    If two orbits $\Gamma_{1},\Gamma_{2}\subset\mathbb{C}{}$, $\Gamma_{1}\not=\Gamma_{2}$, have nonempty intersection with the elementary topological quadrangle $R$, then either $\Gamma_{1}\subset\operatorname{Int}(\Gamma_{2})\cup\{a\})$ or $\Gamma_{2}\subset\operatorname{Int}(\Gamma_{1}\cup\{a\})$.
In fact, suppose that there exist $-\sigma\leq s_{1}<s_{2}\leq\sigma$ with $L_{j}:=\ell_{-}(s_{j})\in\hat{l}_{-}$, $j\in\{1,2\}$, such that $\Gamma(L_{1})\subset\operatorname{Ext}(\Gamma(L_{2})\cup\{a\})$ and $\Gamma(L_{2})\subset\operatorname{Ext}(\Gamma(L_{1})\cup\{a\})$, i.e. $\Gamma(L_{1})$ and $\Gamma(L_{2})$ are mutually exterior in the sense of [2, §17, 3.]. Then we can define the closed piecewise continuously differentiable Jordan curve
$J:=\ell_{-}([s_{1},s_{2}])\cup\ell_{+}([\chi_{2}(s_{1}),\chi_{2}(s_{2})])\cup\bigcup\limits_{j=1}^{2}\zeta(s_{j},[0,\chi_{1}(s_{j})])\subset R\text{.}$
Since $\hat{R}$ is simply connected and $\zeta$ a homeomorphism, we can apply [18, Corollary 52.5] to conclude that also $R$ is simply connected, i.e. we must have $\operatorname{Int}(J)\subset R$. Since $a\not\in R$, we get $a\not\in\operatorname{Int}(J)$. But this is impossible for two mutually exterior orbits $\Gamma(L_{1})$ and $\Gamma(L_{2})$, cf. Figure 1. Thus, the auxiliary result (IMEO) holds.
[…]
Figure 1. Geometrical Visualization of the construction in Step 3 of the proof of Theorem 3.6. The dashed black curve is $J$. If $\Gamma(L_{1})$ and $\Gamma(L_{2})$ are mutually exterior, then $a\in\operatorname{Int}(J)$ and $\Gamma(L_{2})$ cannot cross $\hat{l}_{-}\subset l_{-}$.
Step 4: Proof of openness of $\mathcal{S}$
Since one part of $\hat{l}_{-}$ must be outside of $\Gamma(x)\cup\{a\}$, we can assume w.l.o.g. $\ell_{-}((0,\sigma])\subset\operatorname{Ext}(\Gamma(x)\cup\{a\})$. By property (E), $x=\zeta(0,\tau)$ is an interior point of $R$, i.e. there exists $\delta_{2}>0$ such that $\mathcal{B}_{\delta_{2}}(x)\subset R$. We choose $\delta:=\min\{\delta_{1},\delta_{2}\}$, cf. Step 1, and proof $\mathcal{B}_{\delta}(x)\subset\mathcal{S}$. Let $\hat{x}\in\mathcal{B}_{\delta}(x)$.
By Definition 3.1 (iii) and Proposition 3.4, the cases $\hat{x}\in\Gamma(x)\subset\mathcal{S}$ and $\hat{x}\in\overline{\operatorname{Int}(\Xi\cup\{a\})}\setminus\{a\}\subset\mathcal{S}$ are trivial and nothing has to be shown. Hence, assume $\Gamma(\hat{x})\subset\operatorname{Ext}(\Xi\cup\{a\})\setminus\Gamma(x)$. We can choose $\Gamma_{1}=\Gamma(\hat{x})$ and $\Gamma_{2}=\Gamma(x)$ in (IMEO) to conclude that $\Gamma(\hat{x})$ and $\Gamma(x)$ are not mutually exterior. If $\Xi=\Gamma(x)$, then $\Xi\subset\operatorname{Int}(\Gamma(\hat{x})\cup\{a\})$ and thus $\hat{x}\in\mathcal{S}$. Hence, we additionally assume $\Xi\not=\Gamma(x)$. Since $\Gamma(x)\subset\mathcal{S}$, $\Xi$ and $\Gamma(x)$ are also not mutually exterior. Suppose that $\Xi\subset\operatorname{Ext}(\Gamma(\hat{x})\cup\{a\})$. Then we have the following situation:
  • $\bullet$
    $\Gamma(x)\subset\operatorname{Int}(\Gamma(\hat{x})\cup\{a\})$ or $\Gamma(\hat{x})\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$.
  • $\bullet$
    $\Gamma(x)\subset\operatorname{Int}(\Xi\cup\{a\})$ or $\Xi\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$.
  • $\bullet$
    $\Gamma(\hat{x})$ and $\Xi$ are mutually exterior.
There is only one way in which these three properties are fulfilled: $\Gamma(\hat{x})\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$ and $\Xi\subset\operatorname{Ext}(\Gamma(\hat{x})\cup\{a\})\cap\operatorname{Int}(\Gamma(x)\cup\{a\})$. Since both orbits, $\Gamma(\hat{x})$ and $\Gamma(x)$, cross $\hat{l}_{-}$, also $\Xi$ must cross $\hat{l}_{-}$, cf. Figure 1. It follows that $\Xi\cap R\not=\emptyset$. Hence, by choosing $\Gamma_{1}=\Gamma(\hat{x})$ and $\Gamma_{2}=\Xi$ in (IMEO), $\Gamma(\hat{x})$ and $\Xi$ cannot be mutually exterior. This is a contradiction. Thus, $\Xi\subset\operatorname{Int}(\Gamma(\hat{x})\cup\{a\})$ and $\hat{x}\in\mathcal{S}$. ∎
Theorem 3.7.
Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$. All orbits in $\partial\mathcal{S}\setminus\{a\}$ are unbounded. $\mathcal{S}$ is unbounded.
Proof.
Let $x\in\partial\mathcal{S}\setminus\{a\}$ and suppose that $\Gamma(x)\subset\partial\mathcal{S}\setminus\{a\}$ is bounded. The set $K:=\overline{\Gamma(x)}$ is compact and $\Gamma(x)\cup w_{\pm}(\Gamma(x))\subset K\subset\partial\mathcal{S}$. By [10, Theorem 5.1], $\Gamma(x)$ either is a periodic orbit with exactly one equilibrium, a center $\tilde{a}$, in $\operatorname{Int}(\Gamma(x))$, or $w_{+}(\Gamma(x))$ and $w_{-}(\Gamma(x))$ each consist of exactly one equilibrium. If the first case occurs, by openness of the basin $\mathcal{V}$ in $\tilde{a}$, cf. Theorem 2.4, there would be a point $y\in\mathcal{S}\cap\mathcal{V}$, $y\not=x$. Thus, $\Gamma(y)$ would be periodic with $a$ in its limit sets. This is impossible and the second case in [10, Theorem 5.1] must occur.
By Proposition 3.5, $\Gamma(x)$ must be a homoclinic orbit with $w_{\pm}(\Gamma(x))=\{a\}$. In particular, by [10, Proposition 4.3 (ii)], $\Gamma(x)$ tends to $a$ in definite directions $\eta_{\pm}\in\mathcal{E}(F,m)$ for $t\to\pm\infty$. Moreover, there exist $r,\Theta_{\pm}>0$ such that for all $z_{\pm}\in\{z\in\mathbb{C}{}:|z-a|<r,|\arg(z-a)-\eta_{\pm}|<\Theta_{\pm}\}$ the orbit through $z_{\pm}$ tends to $a$ in the definite direction $\eta_{\pm}$ for $t\to\pm\infty$. Additionally, there exists $\tau>0$ such that $M_{\pm}:=\Phi(\pm\tau,x)\in A_{\pm}$. Choose $\varepsilon_{\pm}>0$ small enough such that $\mathcal{B}_{\varepsilon_{\pm}}(M_{\pm})\subset A_{\pm}$. Since $M_{\pm}\in\partial\mathcal{S}$, there exists $y_{\pm}\in\mathcal{B}_{\varepsilon_{\pm}}(M_{\pm})\cap\mathcal{S}$. As an orbit cannot tend to $a$ in more than one direction for $t\to\infty$ and $t\to-\infty$, it follows $\theta_{+}=\eta_{+}$ and $\theta_{-}=\eta_{-}$, i.e. $\Gamma(x)$ must be $(\theta_{+},\theta_{-})$-elliptic in $a$.
Define $\varepsilon:=\min\{\varepsilon_{+},\varepsilon_{-}\}$. Let $L_{0}$ be the curve segment of $\Gamma(x)$ from $M_{-}$ to $M_{+}$. As in Step 2 of the proof of Theorem 3.6, we choose paths without contact $l_{\pm}\subset\mathcal{B}_{\varepsilon}(M_{\pm})$ through $M_{\pm}$ and construct an elementary topological quadrangle $R$ around $L_{0}$ defined via the functions $\chi_{1}$, $\chi_{2}$ and $\zeta$. Define $\mathcal{S}^{\prime}:=\mathcal{S}^{\prime}(\theta_{+},\theta_{-},\Xi)$. Since $x\in\partial\mathcal{S}$ is an interior point of $R$ and $\operatorname{Ext}(\Xi\cup\{a\})$, there exists $\hat{y}\in\mathcal{S}^{\prime}\cap R$, $\Gamma(\hat{y})\not=\Gamma(x)$.
If $\hat{y}\in\operatorname{Int}(\Gamma(x)\cup\{a\})$, we would have $\Xi\subset\operatorname{Int}(\Gamma(\hat{y})\cup\{a\})\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$ and thus $x\in\partial\mathcal{S}\cap\mathcal{S}$, which is a contradiction to Theorem 3.6. Hence we must have $\hat{y}\in\operatorname{Ext}(\Gamma(x)\cup\{a\})$. Choose $\Gamma_{1}=\Gamma(x)$ and $\Gamma_{2}=\Gamma(\hat{y})\subset\operatorname{Ext}(\Gamma_{1}\cup\{a\})$ in (IMEO) to conclude $\Gamma(x)\subset\operatorname{Int}(\Gamma(\hat{y})\cup\{a\})$. We get $\Xi\subset\operatorname{Ext}(\Gamma(x)\cup\{a\})\cap\operatorname{Int}(\Gamma(\hat{y})\cup\{a\})$. Additionally, both orbits, $\Gamma(x)$ and $\Gamma(\hat{y})$ cross $l_{-}\subset\partial R$. As in Step 4 of the proof of Theorem 3.6, it follows $\Xi\cap R\not=\emptyset$. By choosing $\Gamma_{1}=\Gamma(x)$ and $\Gamma_{2}=\Xi$ in (IMEO), we conclude again $\Xi\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$ and thus $x\in\partial\mathcal{S}\cap\mathcal{S}$. This is again a contradiction to Theorem 3.6. Thus, $\Gamma(x)$ cannot be homoclinic and must be unbounded. It follows that also $\mathcal{S}$ is unbounded. ∎
Proposition 3.8.
Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$. It holds
(3.1) $\mathcal{S}=\bigcup_{\mathclap{x\in\mathcal{S}}}\overline{\operatorname{Int}(\Gamma(x)\cup\{a\})}\setminus\{a\}=\bigcup_{\mathclap{x\in\mathcal{S}}}\operatorname{Int}(\Gamma(x)\cup\{a\})\text{.}$
Moreover, all orbits in $\mathcal{S}$ are nested, i.e. for all $y_{1},y_{2}\in\mathcal{S}$, $\Gamma(y_{1})\not=\Gamma(y_{2})$, there holds either $\Gamma(y_{1})\subset\operatorname{Int}(\Gamma(y_{2})\cup\{a\})$ or $\Gamma(y_{2})\subset\operatorname{Int}(\Gamma(y_{1})\cup\{a\})$.
Proof.
Let $x\in\mathcal{S}$. Suppose there exists $z_{1}\in\operatorname{Int}(\Gamma(x)\cup\{a\})$ such that $z_{1}\not\in\mathcal{S}$. Then there exists $z_{2}\in\partial\mathcal{S}\cap\operatorname{Int}(\Gamma(x)\cup\{a\})\not=\emptyset$. Since the case $a=z_{2}$ is not possible, $\Gamma(z_{2})$ has to be unbounded, cf. Theorem 3.7. But $\Gamma(z_{2})$ is path-connected and thus must lie completely in $\operatorname{Int}(\Gamma(x)\cup\{a\})$, which is bounded. This is a contradiction and $z_{1}$ cannot exist. Hence we have $\operatorname{Int}(\Gamma(x)\cup\{a\})\subset\mathcal{S}$. Additionally, by applying Proposition 3.4, we get
$\overline{\operatorname{Int}(\Gamma(x)\cup\{a\})}\setminus\{a\}=\operatorname{Int}(\Gamma(x)\cup\{a\})\cup\Gamma(x)\subset\mathcal{S}\quad\forall\,x\in\mathcal{S}\text{.}$
By Theorem 3.6, for all $x\in\mathcal{S}$ we can choose $y_{x}\in\operatorname{Ext}(\Gamma(x)\cup\{a\})\cap\mathcal{S}$ such that $\Gamma(x)\subset\operatorname{Int}(\Gamma(y_{x})\cup\{a\})$. It follows
$\overline{\operatorname{Int}(\Gamma(x)\cup\{a\})}\setminus\{a\}\subset\operatorname{Int}(\Gamma(y_{x})\cup\{a\})\subset\mathcal{S}\quad\forall\,x\in\mathcal{S}\text{.}$
These considerations show both equations in (3.1).
Suppose there exist $y_{1},y_{2}\in\mathcal{S}$, $\Gamma(y_{1})\not=\Gamma(y_{2})$, such that $\Gamma(y_{1})$ and $\Gamma(y_{2})$ are mutually exterior, i.e. $\Gamma(y_{1})\subset\operatorname{Ext}(\Gamma(y_{2})\cup\{a\})$ and $\Gamma(y_{2})\subset\operatorname{Ext}(\Gamma(y_{1})\cup\{a\})$.
At first, assume that $\{y_{1},y_{2}\}\cap\operatorname{Ext}(\Xi\cup\{a\})\not=\emptyset$. To fix ideas, assume w.l.o.g. $y_{1}\in\operatorname{Ext}(\Xi\cup\{a\})$. We have $\Xi\subset\operatorname{Int}(\Gamma(y_{1})\cup\{a\})$. Since $\Gamma(y_{2})\subset\operatorname{Ext}(\Gamma(y_{1})\cup\{a\})$, we get $y_{2}\in\operatorname{Ext}(\Xi\cup\{a\})$ and thus $\Xi\subset\operatorname{Int}(\Gamma(y_{1})\cup\{a\})\subset\operatorname{Ext}(\Gamma(y_{2})\cup\{a\})$. It follows the contradiction $y_{2}\not\in\mathcal{S}$.
Secondly, we assume $\{y_{1},y_{2}\}\subset\operatorname{Int}(\Xi\cup\{a\})$. By using (3.1), we have $\operatorname{Int}(\Gamma(y_{1})\cup\{a\})\subset\mathcal{S}$, i.e. all orbits in $\operatorname{Int}(\Gamma(y_{1})\cup\{a\})$ are $(\theta_{+},\theta_{-})$-elliptic in $a$ and $\hat{\Xi}:=\Gamma(y_{1})$ is another sector-forming orbit in $a$ with definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$. Define $\hat{\mathcal{S}}:=\mathcal{S}(\theta_{+},\theta_{-},\hat{\Xi})$. Since $\Xi$ is $(\theta_{+},\theta_{-})$-elliptic and $\hat{\Xi}\subset\operatorname{Int}(\Xi\cup\{a\})$, we get $\Xi\subset\hat{\mathcal{S}}$. By using (3.1), we conclude $y_{2}\in\operatorname{Int}(\Xi\cup\{a\})\subset\hat{\mathcal{S}}$. Additionally, since $y_{2}\in\operatorname{Ext}(\hat{\Xi}\cup\{a\})$, we must have $\Gamma(y_{1})=\hat{\Xi}\subset\operatorname{Int}(\Gamma(y_{2})\cup\{a\})$, which is a contradiction.
The last possible case is $\{y_{1},y_{2}\}\cap\Xi\not=\emptyset$. To fix ideas, we assume w.l.o.g. $\Xi=\Gamma(y_{1})$. By Definition 3.1 (iii), we must have either $y_{2}\in\operatorname{Int}(\Xi\cup\{a\})=\operatorname{Int}(\Gamma(y_{1})\cup\{a\})$ or $\Gamma(y_{1})=\Xi\subset\operatorname{Int}(\Gamma(y_{2})\cup\{a\})$. Both results lead to a contradiction.
Finally, $y_{1}$ and $y_{2}$ cannot exist and we conclude that all orbits in $\mathcal{S}$ must be nested. ∎
Theorem 3.9.
Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$. $\mathcal{S}$ is connected, path-connected and simply connected.
Proof.
Suppose, there exists a separation $(U,V)$ of $\mathcal{S}$, i.e. $\mathcal{S}\subset U\cup V$, where $U$ and $V$ are open, disjoint and have nonempty intersection with $\mathcal{S}$. Then $a$ can only be in one of these two sets. Assume w.l.o.g. $a\in U$. Choose $y\in V\cap\mathcal{S}$. Since $\Gamma(y)$ is connected, we get $\Gamma(y)\subset V$. Since $a\not\in\overline{V}$, it follows the contradiction $a\not\in w_{+}(\Gamma(y))\cup w_{-}(\Gamma(y))$. Hence, such a separation cannot exist and $\mathcal{S}$ is connected. By [23, Proposition 12.25], $\mathcal{V}$ is even path-connected.
Let $L\subset\mathcal{S}$ be a loop. Define $J_{x}:=\Gamma(x)\cup\{a\}$ for $x\in\mathcal{S}$. Due to the geometry described in Proposition 3.8, we can use the same idea as in the proof of Theorem 2.5. By Theorem 3.6, for all $x\in L$ exists $y_{x}\in\operatorname{Ext}(J_{x})\cap\mathcal{S}$ such that $x\in\operatorname{Int}(J_{y_{x}})$. We can replace the periodic orbits $\Gamma(y_{x})$, $x\in\mathcal{V}$, by the curves $J_{y_{x}}$, $x\in\mathcal{S}$. Thus, we can find a suitable finite open cover of $L$ and an outermost orbit $J\subset\mathcal{S}$ satisfying $a\in J$ and $L\subset\operatorname{Int}(J)$. By (3.1), we have $\operatorname{Int}(J\cup\{a\})\subset\mathcal{S}$, which is clearly simply connected. Hence $L$ can be continuously deformed to a point in $L\subset\mathcal{S}$. Since $\mathcal{S}$ is path-connected, it follows that $\mathcal{S}$ is even simply connected. ∎
We now address the independence of the global elliptic sector from a particular choice of the sector-forming orbit.
Lemma 3.10.
Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$. For all $x\in\mathcal{S}$ the orbit $\Gamma(x)$ is a sector-forming orbit in $a$ with definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ and $\mathcal{S}=\mathcal{S}(\theta_{+},\theta_{-},\Gamma(x))$, i.e. the global elliptic sector $\mathcal{S}$ does not depend on the particular choice of a sector-forming orbit.
Proof.
Let $x\in\mathcal{S}$ be arbitrary. By (3.1), all orbits in $\operatorname{Int}(\Gamma(x)\cup\{a\})$ are $(\theta_{+},\theta_{-})$-elliptic and $\hat{\Xi}:=\Gamma(x)$ is indeed an sector-forming orbit in $a$ with definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$.
Define $\hat{\mathcal{S}}:=\mathcal{S}(\theta_{+},\theta_{-},\hat{\Xi})$. If $\hat{\Xi}=\Xi$, then obviously $\mathcal{S}=\hat{\mathcal{S}}$. We have to show $\mathcal{S}\subset\hat{\mathcal{S}}$ as well as $\hat{\mathcal{S}}\subset\mathcal{S}$ for the two cases $\hat{\Xi}\subset\operatorname{Int}(\Xi\cup\{a\})$ and $\Xi\subset\operatorname{Int}(\hat{\Xi}\cup\{a\})$. As these four proofs are all based on similar arguments, we assume w.l.o.g. $\hat{\Xi}\subset\operatorname{Int}(\Xi\cup\{a\})$ and show only $\mathcal{S}\subset\hat{\mathcal{S}}$. Let $y\in\mathcal{S}$. If $y\in\Xi\cup\hat{\Xi}\cup\operatorname{Int}(\hat{\Xi}\cup\{a\})$, then nothing is to show. Thus, assume $y\in\operatorname{Ext}(\hat{\Xi}\cup\{a\})\setminus\Xi$. If $y\in\operatorname{Ext}(\Xi\cup\{a\})$, then we have $\Xi\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$ and thus $\hat{\Xi}\subset\operatorname{Int}(\Xi\cup\{a\})\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$, i.e. $y\in\hat{\mathcal{S}}$. Additionally, if $y\in\operatorname{Ext}(\hat{\Xi}\cup\{a\})\cap\operatorname{Int}(\Xi\cup\{a\})$, then we can apply Proposition 3.8 with $y_{1}:=x$ and $y_{2}:=y$ to conclude $\hat{\Xi}\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$, i.e. $y\in\hat{\mathcal{S}}$. ∎
Finally, at the end of this chapter, we want to supplement and enhance our topological study with two important corollaries, which confirm our intuitive idea of the geometrical structure of an equilibrium of order $m\geq 2$, cf. [10, Figure 2, Figure 3].
Corollary 3.11.
There exist exactly $22m-$ 2 distinct444in the sense of [2, §17, 3.]. global elliptic sectors in $a$. For all $x\in\mathbb{C}{}$ with $w_{+}(\Gamma(x))=w_{-}(\Gamma(x))=\{a\}$ there exist two adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ such that $\Gamma(x)$ is a sector-forming orbit for $\mathcal{S}(\theta_{+},\theta_{-},\Gamma(x))$.
Proof.
By [10, Theorem 4.4], there is a FED of order $d:=2m-2$ in $a$. We use the geometrical objects $\Gamma$, $\Gamma_{i}$ and $p_{i}$, $i\in\{1,\ldots,d\}$, defined by [10, Definition 4.1 d)]. We order the definite directions $\mathcal{E}(F,m)=\{\theta_{1},\ldots,\theta_{d}\}$ in such a way that for all $i\in\{1,\ldots,d\}$ the local elliptic sector with characteristic orbits $\Gamma_{i}$ and $\Gamma_{i+1}$ lies between $\theta_{i}$ and $\theta_{i+1}$, i.e. $\Gamma_{i}$ tends to $a$ in the definite direction $\theta_{i}$. [^5] If $i=d$, we define $\Gamma_{i+1}:=\Gamma_{1}$ and $\theta_{i+1}:=\theta_{1}$. By using the notation in [10, Definition 4.1 a)] and applying [10, Definition 4.1 c) (iv)], there exist points $E_{i}\in\Gamma(p_{i},p_{i+1})$ such that $\Xi_{i}:=\Gamma(E_{i})$ is a sector-forming orbit with definite directions $\theta_{i},\theta_{i+1}$ for all $i\in\{1,\ldots,d\}$. Moreover, since all local sectors in $a$ have pairwise empty intersection except for the characteristic orbits and $a$, these sector-forming orbits are pairwise mutually exterior. Define the set
$A:=\overline{\operatorname{Int}(\Gamma)}\setminus\bigcup_{i=1}^{d}\overline{\operatorname{Int}(\Xi_{i}\cup\{a\})}$
as well as the $d$ global elliptic sectors
$\mathcal{S}_{i}:=\begin{dcases}\mathcal{S}(\theta_{i},\theta_{i+1},\Xi_{i})&\text{if }\lambda(\theta_{i})=-1\\ \mathcal{S}(\theta_{i+1},\theta_{i},\Xi_{i})&\text{if }\lambda(\theta_{i})=1\end{dcases}\;,\,i\in\{1,\ldots,d\}\text{.}$
Suppose there exist $i,j\in\{1,\ldots,d\}$, $i\not=j$, with $z\in\mathcal{S}_{i}\cap\mathcal{S}_{j}\not=\emptyset$. Since $\Xi_{i}$ and $\Xi_{j}$ are mutually exterior, we must have $z\in\operatorname{Ext}(\Xi_{i}\cup\{a\})\cap\operatorname{Ext}(\Xi_{j}\cup\{a\})$ and thus $\Xi_{i}\cup\Xi_{j}\subset\operatorname{Int}(\Gamma(z)\cup\{a\})$. By (3.1), we have $\operatorname{Int}(\Gamma(z)\cup\{a\})\subset\mathcal{S}_{i}$, i.e. $\Xi_{i}\cup\Xi_{j}\subset\mathcal{S}_{i}$. Since $\Xi_{i}$ and $\Xi_{j}$ are mutually exterior, we can derive a contradiction to the second statement in Proposition 3.8. Thus, these global elliptic sectors are distinct.
Let $x\in\mathbb{C}{}\setminus\{a\}$ be arbitrary with $w_{+}(\Gamma(x))=w_{-}(\Gamma(x))=\{a\}$. By [10, Proposition 4.3 (ii)], $\Gamma(x)$ has to tend to $a$ in definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ for $t\to\infty$ and $t\to-\infty$, respectively.
At first, we assume $\Gamma(x)\subset\overline{\operatorname{Int}(\Gamma)}$. Then, by applying the geometrical structure of the FED and [10, Proposition 4.3 (iii)], all orbits through $A$ leave $\overline{\operatorname{Int}(\Gamma)}$ in finite time. Hence, there exists $i\in\{1,\ldots,d\}$ such that $\Gamma(x)\subset\Xi_{i}\cup\operatorname{Int}(\Xi_{i}\cup\{a\})\subset\mathcal{S}_{i}$. By Definition 3.1 (iii), it follows that $\{\theta_{+},\theta_{-}\}=\{\theta_{i},\theta_{i+1}\}$, i.e. $\theta_{+},\theta_{-}$ are adjacent.
Next, we consider the case $\Gamma(x)\cap\operatorname{Ext}(\Gamma)\not=\emptyset$. As $\Gamma(x)$ must leave and enter $\overline{\operatorname{Int}(\Gamma)}$ again, a part of the closed Jordan curve $\Gamma$ must lie in $\operatorname{Int}(\Gamma(x)\cup\{a\})$. By [10, Definition 4.1 c) (iii)], there must exist $i\in\{1,\ldots,d\}$ with $E_{i}\in\operatorname{Int}(\Gamma(x)\cup\{a\})$, i.e. $\Xi_{i}\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$. Suppose a second sector-forming orbit is in the interior of $\Gamma(x)\cup\{a\}$, i.e. there exists $j\not=i$ with $\Xi_{j}\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$. If $x\in\mathcal{S}_{i}$, then, by applying the second statement in Proposition 3.8, $\Xi_{i}$ and $\Xi_{j}$ cannot be mutually exterior, which is a contradiction. Thus, $x\not\in\mathcal{S}_{i}$. Since $\mathcal{S}_{i}$ is connected, cf. Theorem 3.9, we conclude $\mathcal{S}_{i}\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$. This makes $\mathcal{S}_{i}$ bounded, which is a contradiction to Theorem 3.7. Hence, $\Xi_{i}$ is the only sector-forming orbit in $\operatorname{Int}(\Gamma(x)\cup\{a\})$. By applying the geometrical structure of the FED, we again get $\{\theta_{+},\theta_{-}\}=\{\theta_{i},\theta_{i+1}\}$, i.e. $\theta_{+},\theta_{-}$ are adjacent. In particular, by Definition 3.1 (iii), also in this case it follows $x\in\mathcal{S}_{i}$.
Finally, in both cases we can apply Lemma 3.10 to conclude that $\Gamma(x)$ is a sector forming orbit for the global elliptic sector $\mathcal{S}(\theta_{+},\theta_{-},\Gamma(x))$ and $\mathcal{S}(\theta_{+},\theta_{-},\Gamma(x))=\mathcal{S}_{i}$. This shows the second statement. Moreover, using this result, there cannot exist more then the $d$ distinct global elliptic sectors $\mathcal{S}_{1},\ldots,\mathcal{S}_{d}$. This completes the proof of the first statement. ∎
We have proved Corollary 3.11 by applying the geometry of the local structure in $a$. This structure is defined by a FED, which is not uniquely determined, i.e. there exist several FEDs in $a$. Nevertheless, we can now see that the global structure is indeed unique.
Corollary 3.12.
Let $m\geq 3$. Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$. Then it holds
(3.2) $\mathcal{S}(\theta_{+},\theta_{-},\Xi)=\left\{x\in\mathbb{C}{}:\Gamma(x)\text{ is }(\theta_{+},\theta_{-})\text{-elliptic in }a\right\}\text{.}$
Proof.
By Definition 3.1, $\subset$" is clear. Since each of the $22m-$ 2 distinct global elliptic sectors lies between two definite directions and $m\geq 3$, there exists exactly one global elliptic sector $\mathcal{S}_{0}$ between $\theta_{+}$ and $\theta_{-}$, cf. Corollary 3.11. If there was a second one, there would be only $22<2m-$ 2 global elliptic sectors in $a$. Thus, we must have $\mathcal{S}(\theta_{+},\theta_{-},\Xi)=\mathcal{S}_{0}$. Moreover, by applying the second statement in Corollary 3.11, all $(\theta_{+},\theta_{-})$-elliptic orbits must lie in one of these $22m-$ 2 global elliptic sectors. Since the definite directions on the left hand side coincide with those on the right hand side in (3.2), all these orbits have to lie in $\mathcal{S}_{0}$ and thus in $\mathcal{S}(\theta_{+},\theta_{-},\Xi)$. This shows $\supset$". ∎
This second corollary shows that we need to choose a sector-forming orbit only in the case $m=2$. Here we have only two definite directions, which requires the choice of one arbitrary elliptic orbit to specify one of the two global elliptic sectors. The two sectors cannot be distinguished on the basis of the definite directions alone. If we consider the set on the right hand side of (3.2) in the case $m=2$, then this set is still open and unbounded, but not connected. The set would become connected (but no longer open) only by adding the equilibrium $a$. For this reason, the construction in Definition 3.1 uses a sector-forming orbit.
We observe that the boundaries of two adjacent global elliptic sectors do not have to coincide near $a$, as it is the case for the characteristic orbits in a FED, cf. [10, Theorem 4.4]. However, if they indeed coincide between all adjacent sectors, the local structure of the FED can even be transferred to the global phase portrait. The question therefore arises as to what the global geometric structure looks like in a region between two adjacent global elliptic sectors having no common boundary near $a$. How many unbounded orbits can exist in sum? This question takes up the discussions at the end of [10, Example 4.7] and [10, Example 4.8].
4. Basin of nodes and foci
This chapter provides detailed proofs of the topological results given in [4, Theorem 4.3]. In all claims and arguments of this chapter we consider an entire function $F\in\mathcal{O}{}(\mathbb{C})$, $F\not\equiv 0$, and a node or focus $a\in\mathbb{C}$ of (1.1).
Definition 4.1.
The stable (unstable) basin $\mathcal{N}$ of $F$ in $a$ is defined as
$\mathcal{N}:=\left\{x\in\mathbb{C}:w_{+(-)}(\Gamma(x))=\{a\}\right\}.$
The boundary of $a$ is defined as the boundary of the basin of $F$ in $a$.
Proposition 4.2.
$\mathcal{N}$ and $\partial\mathcal{N}$ are flow-invariant. It holds
$\mathcal{N}=\bigcup_{\mathclap{x\in\mathcal{N}}}\Gamma(x)\text{.}$
Proof.
This proof is analogous to the proof of Proposition 2.2. ∎
Theorem 4.3.
It holds $a\in\mathring{\mathcal{N}}$. $\mathcal{N}$ is open, connected and path-connected.
Proof.
Assume w.l.o.g. that $a$ is stable. By [10, Definition 3.1], there exists $\varepsilon_{1}>0$ such that for all $y\in\mathcal{B}_{\varepsilon_{1}}(a)$ we have $w_{+}(\Gamma(y))=\{a\}$. This proves $\mathcal{B}_{\varepsilon_{1}}(a)\subset\mathcal{N}$ and $a\in\mathring{\mathcal{N}}$. Let $x\in\mathcal{N}\setminus\{a\}$ be arbitrary. Then there exists a $\tau>0$ with $\xi:=\Phi(\tau,x)\in\mathcal{B}_{\varepsilon_{1}}(a)$. Choose $\varepsilon_{2}:=\frac{1}{2}\min\{|\xi-a|,\varepsilon_{1}-|\xi-a|\}>0$ such that $\mathcal{B}_{\varepsilon_{2}}(\xi)\subset\mathcal{B}_{\varepsilon_{1}}(a)$. By continuity of the flow, cf. [19, Chapter 2.4, Theorem 4], there exists a $\delta>0$ such that $|\Phi(\tau,z)-\xi|<\varepsilon_{2}$ for all $z\in\mathcal{B}_{\delta}(x)$. Thus $w_{+}(\Gamma(z))=\{a\}$ for all $z\in\mathcal{B}_{\delta}(x)$ and $\mathcal{N}$ is open.
For $x\in\mathcal{N}$ set $J_{x}:=\Gamma(x)\cup\{a\}$ and define the cover $C:=\{J_{x}:x\in\mathcal{N}\}$ of $\mathcal{N}$. Since $a$ lies on the boundary of each orbit in $\mathcal{N}$, every set in $C$ is connected, cf. [18, Theorem 23.4]. Moreover, for $J,\tilde{J}\in C$ it always holds $J\cap\tilde{J}\supset\{a\}\not=\emptyset$. Thus, by [18, Theorem 23.3], $\mathcal{N}$ is connected. From [23, Proposition 12.25] it follows that $\mathcal{N}$ is even path-connected. ∎
Lemma 4.4.
All centers $\tilde{a}\in F^{-1}(\{0\})\setminus\{a\}$ with its corresponding basin $\mathcal{V}$ satisfy $\mathcal{V}\cap\partial\mathcal{N}=\emptyset$. In particular, there are no centers on $\partial\mathcal{N}$.
Proof.
Assume w.l.o.g. that $a$ is stable. Suppose there exists a center $\tilde{a}\in\mathbb{C}{}\setminus\{a\}$ such that its basin $\mathcal{V}$ satisfies $\mathcal{V}\cap\partial\mathcal{N}\not=\emptyset$. Since $\mathcal{V}$ is open, cf. Theorem 2.4, we can choose $z\in\mathcal{V}\cap\partial\mathcal{N}$ and $r>0$ sufficiently small such that $\mathcal{B}_{r}(z)\subset\mathcal{V}$. Since $z\in\partial\mathcal{N}$, there exists $y\in\mathcal{N}\cap\mathcal{B}_{r}(z)\not=\emptyset$ satisfying $w_{+}(\Gamma(y))=\{a\}$. But this is impossible for a periodic orbit in $\mathcal{V}$. Thus $\mathcal{V}\cap\partial\mathcal{N}=\emptyset$. In particular, there are no centers on $\partial\mathcal{N}$, i.e. all equilibria on the boundary of $a$ have to be nodes, foci or zeros of order $m\geq 2$. ∎
In the proof of [4, Theorem 4.3] the author does not investigate the case of an isolated equilibrium on the boundary of a node or focus. This is a significant gap, which is closed by the following result.
Proposition 4.5.
It holds
$\forall\,\tilde{a}\in\partial\mathcal{N}\cap F^{-1}(\{0\}):\forall\,\rho>0:\left(\mathcal{B}_{\rho}(\tilde{a})\cap\partial\mathcal{N}\right)\setminus\{\tilde{a}\}\not=\emptyset,$
i.e. there are no isolated points with respect to the subspace topology on $\partial\mathcal{N}$. Moreover, for all $\tilde{a}\in\partial\mathcal{N}\cap F^{-1}(\{0\})$ there exists an unbounded orbit $\Gamma\subset\partial\mathcal{N}$ with $\tilde{a}\in w_{+}(\Gamma)\cup w_{-}(\Gamma)$, i.e. all equilibria on the boundary of $a$ are attached to an orbit on the boundary of $a$.
Proof.
Suppose there exists an isolated point $\tilde{a}\in\partial\mathcal{N}\cap F^{-1}(\{0\})$, i.e. there exists $\rho>0$ such that
$\partial\mathcal{N}\cap(\underbrace{\mathcal{B}_{\rho}(\tilde{a})\setminus\{\tilde{a}\}}_{=:A})=\left(\mathcal{B}_{\rho}(\tilde{a})\cap\partial\mathcal{N}\right)\setminus\{\tilde{a}\}=\emptyset\text{.}$
It follows [^6] The simple but technical proof of ^$\hat{\partial}(\mathcal{N}\cap A)\subset\partial\mathcal{N}\cap A=\emptyset$, where $\hat{\partial}$ denotes the boundary with respect to the subspace topology on $A$, is omitted. that $\mathcal{N}\cap A\not=\emptyset$ is open and closed with respect to the subspace topology on $A$. Since the punctured circle $A$ is connected, we must have $\mathcal{N}\cap A=A$. Hence, since $\tilde{a}\not\in\mathcal{N}$, we conclude
(4.1) $\mathcal{N}\cap\mathcal{B}_{\rho}(\tilde{a})=\mathcal{B}_{\rho}(\tilde{a})\setminus\{\tilde{a}\}\text{.}$
If $\tilde{a}$ was of order $m\geq 2$, then, by [10, Theorem 4.4], there would be at least one homoclinic orbit in $A$ tending to $\tilde{a}$ in both time directions, i.e. equation (4.1) would not be fulfilled and $\tilde{a}$ would not be isolated. Thus, $\tilde{a}$ must be a node or focus, cf. Lemma 4.4. Denote by $\tilde{\mathcal{N}}$ the basin of $F$ in $\tilde{a}$. We reduce $\rho$ such that $\mathcal{B}_{\rho}(\tilde{a})\subset\tilde{\mathcal{N}}$ and $\mathcal{B}_{\rho}(a)\subset\mathcal{N}$. By (4.1) and Proposition 4.2, we have $\tilde{\mathcal{N}}\subset\mathcal{N}$.
By applying the theory of circles without contact around nodes and foci, cf. [2, §3, 10.-14., §7, 1.-2. and §18, Lemma 3], we find two continuously differentiable closed paths $C_{1}\subset\mathcal{B}_{\rho}(a)$ and $C_{2}\subset\mathcal{B}_{\rho}(\tilde{a})$, $C_{1}\cap C_{2}=\emptyset$, being nowhere tangential to $F$ and satisfying $\operatorname{Int}(C_{1})\cap F^{-1}(\{0\})=\{a\}$ and $\operatorname{Int}(C_{2})\cap F^{-1}(\{0\})=\{\tilde{a}\}$. [^7] In particular, from the equations (6) and (11) in [2, §7, 1.] and the remarks made in [2, §7, 2.] it follows that $C_{1}$ as well as $C_{2}$ can be chosen as a linear transformed circle or ellipse. It holds $\overline{\operatorname{Int}(C_{1})}\cap\overline{\operatorname{Int}(C_{2})}=\emptyset$. Moreover, every orbit in $\mathcal{N}$ ($\tilde{\mathcal{N}}$) crosses $C_{1}$ ($C_{2}$) exactly once, cf. [2, §3, 10., Figure 54]. We equip $C_{1}$ and $C_{2}$ with the subspace topology. Using this construction, the map $\Psi:\mathcal{N}\setminus\{a\}\to\mathbb{R}$ is well-defined, when $\Psi(x)$ is the transit time888cf. [4, Definition 3.2]. the point $x\in\mathcal{N}\setminus\{a\}$ needs to $C_{1}$ along the path $\Gamma(x)$, i.e. $\{\Phi(\Psi(x),x)\}=\Gamma(x)\cap C_{1}$. Moreover, we define $\eta:C_{2}\to C_{1}$ by $\eta(x):=\Phi(\Psi(x),x)$. By the continuous dependence of the flow on initial conditions, cf. [19, Chapter 2.4, Theorem 4], and the continuous dependence of the transit time integral on the integral limits, $\Psi$ and $\eta$ are both continuous.
One the one hand, since $C_{2}$ is compact, also $\eta(C_{2})$ is compact in $C_{1}$, cf. [18, Theorem 26.5]. Thus, $\eta(C_{2})$ is closed in $C_{1}$, cf. [18, Theorem 26.3]. On the other hand, by Proposition 4.2, we have $\eta(C_{2})\subset\tilde{\mathcal{N}}\cap C_{1}$. Additionally, since all orbits through $\tilde{\mathcal{N}}\cap C_{1}\subset\tilde{\mathcal{N}}$ have to cross $C_{2}$, we also have $\tilde{\mathcal{N}}\cap C_{1}\subset\eta(C_{2})$, i.e. $\eta(C_{2})=\tilde{\mathcal{N}}\cap C_{1}$. Moreover, by Theorem 4.3, $\tilde{\mathcal{N}}\cap C_{1}$ is open in $C_{1}$. We conclude that $\eta(C_{2})\not=\emptyset$ is open and closed in $C_{1}$. Since $C_{1}$ is connected, we must have $C_{1}=\eta(C_{2})$, i.e. $\eta$ is surjective. Let $y_{1},y_{2}\in C_{2}$ be arbitrary. Assume $\eta(y_{1})=\eta(y_{2})$. Then, by definition of $\eta$, $\Gamma(y_{1})=\Gamma(\eta(y_{1}))=\Gamma(\eta(y_{2}))=\Gamma(y_{2})$. Since all orbits in $\tilde{\mathcal{N}}$ cross $C_{2}$ exactly once, we must have $y_{1}=y_{2}$. Hence, $\eta$ is also injective. It follows that $\eta$ is bijective and thus a homeomorphism, cf. [18, Theorem 26.6]. We conclude that every orbit in $\mathcal{N}$ and $\tilde{\mathcal{N}}$ crosses $C_{1}$ as well as $C_{2}$ exactly once, i.e. [^9] The set $\mathcal{H}$ is the so-called heteroclinic region between $a$ and $\tilde{a}$ and is introduced in Chapter 5.
(4.2) $\mathcal{H}:=\mathcal{N}\cap\tilde{\mathcal{N}}=\mathcal{N}\setminus\{a\}=\tilde{\mathcal{N}}\setminus\{\tilde{a}\}=\bigcup\limits_{\mathclap{x_{2}\in C_{2}}}\Gamma(x_{2})=\bigcup\limits_{\mathclap{x_{1}\in C_{1}}}\Gamma(x_{1})\text{.}$
Since $C_{1}\subset\mathcal{H}$ is a Jordan curve, there exists a homeomorphism $h:C_{1}\to S^{1}$. Equip the cylinder surface $B:=\mathbb{R}\times S^{1}\subset\mathbb{R}^{3}$ with the subspace topology and define the map $\zeta:B\to\mathcal{H}$ by $\zeta(t,x):=\Phi(t,h^{-1}(x))$. By (4.2), we have $\zeta(B)=\mathcal{H}$, i.e. $\zeta$ is surjective. Let $(t_{1},x_{1}),(t_{2},x_{2})\in B$ be arbitrary and assume $\zeta(t_{1},x_{1})=\zeta(t_{2},x_{2})$. Then $h^{-1}(x_{1}),h^{-1}(x_{2})\in C_{1}$ and
$\Gamma(h^{-1}(x_{1}))=\Gamma(\zeta(t_{1},x_{1}))=\Gamma(\zeta(t_{2},x_{2}))=\Gamma(h^{-1}(x_{2}))\text{.}$
Since all orbits in $\mathcal{H}$ cross $C_{1}$ exactly once, we get $h^{-1}(x_{1})=h^{-1}(x_{2})$. Since $h$ is bijective, we conclude $x_{1}=x_{2}$. This gives us $\Phi(t_{1},h^{-1}(x_{1}))=\Phi(t_{2},h^{-1}(x_{1}))$ and thus $t_{1}=t_{2}$, since $\Gamma(h^{-1}(x_{1}))$ is not periodic. It follows that $\zeta$ is also injective and thus bijective. Hence, there exists the inverse $\zeta^{-1}$ of $\zeta$, which is given by $\zeta^{-1}(x):=(\Psi(x),h(\Phi(\Psi(x),x)))$, $x\in\mathcal{H}$. As a composition of continuous functions, $\zeta$ as well as $\zeta^{-1}$ are continuous and therefore $\zeta$ is a homeomorphism.
By [18, Corollary 52.5], the fundamental group of $\mathcal{H}$ is isomorphic to $B$. Additionally, as the circle line $S^{1}\subset\mathbb{R}^{2}$ is a deformation retract of the cylinder surface $B$, we conclude that the fundamental group of $\mathcal{H}$ is isomorphic to the additive group $\mathbb{Z}$, cf. [18, Theorem 54.5] and [18, Theorem 58.3]. But $\mathcal{H}$ has the two holes $a$ and $\tilde{a}$, which is impossible for a set having a fundamental group isomorphic to $\mathbb{Z}$. In particular, by using (4.2), we can choose the loop $C_{2}\cup\Phi([0,\Psi(z)],z)\cup C_{1}\subset\mathcal{H}$ with an arbitrary point $z\in C_{2}$. This loop has $a\not\in\mathcal{H}$ as well as $\tilde{a}\not\in\mathcal{H}$ in its interior and thus cannot be transferred to a loop on $S^{1}$. All in all, we derive a contradiction and $\tilde{a}$ cannot be isolated on $\partial\mathcal{N}$.
Moreover, all equilibria $\hat{a}\in\partial\mathcal{N}\cap F^{-1}(\{0\})$ are nodes, foci or equilibria of order $m\geq 2$, cf. Lemma 4.4. Thus, we can choose $\hat{\rho}$ such that for all $\xi\in\mathcal{B}_{\hat{\rho}}(\hat{a})$ there holds $\hat{a}\in w_{+}(\Gamma(\xi))\cup w_{-}(\Gamma(\xi))$, cf. [10, Definition 3.1] and [10, Theorem 4.4]. If we choose $\xi\in(\mathcal{B}_{\hat{\rho}}(\hat{a})\cap\partial\mathcal{N})\setminus\{\hat{a}\}\not=\emptyset$, then $\Gamma:=\Gamma(\xi)\subset\mathcal{N}$ satisfies $\hat{a}\in w_{+}(\Gamma)\cup w_{-}(\Gamma)$, i.e. $\hat{a}$ is attached to $\Gamma$. ∎
Theorem 4.6.
$\partial\mathcal{N}$ is either empty or consists only of equilibria and unbounded orbits. $\mathcal{N}$ is unbounded.
Proof.
If $\partial\mathcal{N}=\emptyset$, then $\mathcal{N}$ is open, closed and nonempty, i.e. $\mathcal{N}=\mathbb{C}{}$ is clearly unbounded. Thus, assume $\partial\mathcal{N}\not=\emptyset$. Additionally, assume w.l.o.g. that $a$ is stable.
Suppose there exists a bounded orbit $\Gamma\subset\partial\mathcal{N}\setminus F^{-1}(\{0\})$. As in the proof of Theorem 2.6, $K:=\overline{\Gamma}$ is compact and $\Gamma\cup w_{\pm}(\Gamma)\subset K\subset\partial\mathcal{N}$. By Lemma 4.4, $\Gamma$ cannot be periodic. Hence, by [10, Theorem 5.1], there exist two equilibria $a_{+},a_{-}\in\partial\mathcal{N}\cap F^{-1}(\{0\})$ such that $w_{\pm}(\Gamma)=\{a_{\pm}\}$, i.e. $\Gamma$ is either heteroclinic or homoclinic. In both cases we want to derive a contradiction.
Suppose $\Gamma$ is homoclinic, i.e. $a_{+}=a_{-}$. By Corollary 3.11, $\Gamma$ is a sector-forming orbit lying in an global elliptic sector $\mathcal{S}_{0}$ in $a_{+}$. Choose $z\in\Gamma\subset\mathcal{S}_{0}\cap\partial\mathcal{N}$. Since $\mathcal{S}_{0}$ is open, cf. Theorem 3.6, there exists $\varepsilon>0$ such that $\mathcal{B}_{\varepsilon}(z)\subset\mathcal{S}_{0}$. Thus, analogous to the proof of Lemma 4.4, we can choose a point $y\in\mathcal{S}_{0}\cap\mathcal{N}\not=\emptyset$, i.e. $\Gamma(y)\subset\mathcal{N}$ is a homoclinic orbit. This is a contradiction.
Suppose, $\Gamma$ is heteroclinic, i.e. $a_{+}\not=a_{-}$. Since $a_{+},a_{-}\in\partial\mathcal{N}$ and $a\in\mathring{\mathcal{N}}$, cf. Theorem 4.3, it holds $a\not\in\{a_{+},a_{-}\}$. Thus, $a_{+}$ cannot be a node or focus, since its basin would have nonempty intersection with $\mathcal{N}$ [^10] If we investigated the case that $a$ is unstable, we would choose $a_{-}$ instead of $a_{+}$ at this point. Note that an orbit in $\mathcal{N}$ cannot have $a$ as well as $a_{+}$ in its limit set for $t\to\infty$. Additionally, by Lemma 4.4, $a_{+}\in\partial\mathcal{N}$ cannot be a center. It follows that $a_{+}$ must have order $m\geq 2$. By [10, Proposition 4.3], there exist $\theta\in\mathcal{E}(F,m)$ and $r,\delta>0$ such that $\Gamma$ as well as the orbit $\Gamma(x_{0})$ through $x_{0}\in\{x\in\mathbb{R}^{2}:|x-a_{+}|<r,|\arg(x-a_{+})-\theta|<\delta\}=:A$ tends to $a_{+}$ for $t\to\infty$. Choose $\hat{z}\in\Gamma\cap A\subset\partial\mathcal{N}\cap A$. Since $A$ is open, we can choose $\hat{\varepsilon}>0$ such that $\mathcal{B}_{\hat{\varepsilon}}(\hat{z})\subset A$. Thus, again analogous to the proof of Lemma 4.4, we can choose $x_{0}\in A\cap\partial\mathcal{N}\not=\emptyset$, i.e. $\Gamma(x_{0})\subset\mathcal{N}$ tends to $a_{+}\not=a$ for $t\to\infty$. This is a contradiction.
We conclude that $\Gamma$ cannot be bounded. All in all, it follows that $\partial\mathcal{N}$ consists only of equilibria (but no centers) and unbounded orbits. Since $\partial\mathcal{N}\not=\emptyset$, we can apply Proposition 4.5 to conclude that there exists at least one unbounded orbit lying on the boundary of $a$. Thus, also $\mathcal{N}$ is unbounded. ∎
With the results we have obtained so far, we can now prove that the basin is even simply connected. The crucial factor here is Proposition 4.5, which excludes the existence of holes in the form of isolated equilibria on the boundary of $a$.
Theorem 4.7.
$\mathcal{N}$ is simply connected.
Proof.
Let $L\subset\mathcal{N}$ be a loop. Since $L$ is compact and $\mathcal{N}$ is open, cf. Theorem 4.3, there exist points $y_{1},\ldots,y_{\ell}\in L$ and radii $r_{1},\ldots,r_{\ell}>0$, $\ell\in\mathbb{N}{}$, such that
$L\subset\bigcup_{k=1}^{\ell}\mathcal{B}_{r_{k}}(y_{k})=:O$
and $\overline{O}\subset\mathcal{N}$. The latter property can be ensured by reducing the radii. In addition, there exists a closed Jordan curve $J\subset\partial O\subset\mathcal{N}$ such that $L\subset\operatorname{Int}(J)$. If $\operatorname{Int}(J)\setminus\mathcal{N}\not=\emptyset$, we can choose $x\in\partial\mathcal{N}\cap\operatorname{Int}(J)\not=\emptyset$. [^11] A detailed topological argumentation for this can be found in the proof of [9, Theorem 5.10]. If $x$ is not an equilibrium, $\Gamma(x)$ must be unbounded, cf. Theorem 4.6. If $x$ is an equilibrium, there exists an orbit $\Gamma\subset\partial\mathcal{N}$ with $x$ in its limit set, cf. Proposition 4.5. Thus, by openness of $\operatorname{Int}(J)$, we have $\Gamma\cap\operatorname{Int}(J)\not=\emptyset$. In both cases we get an unbounded orbit with nonempty intersection with $\operatorname{Int}(J)$. But $\operatorname{Int}(J)$ is bounded, i.e. this is a contradiction. Thus, we have $\operatorname{Int}(J)\subset\mathcal{N}$. Since $\operatorname{Int}(J)$ is simply connected, $L$ can be continuously deformed to a point. Since $L$ is arbitrary and $\mathcal{N}$ is already path-connected, it follows that $\mathcal{N}$ is even simply connected. ∎
5. Heteroclinic regions
In this chapter we consider an entire function $F\in\mathcal{O}{}(\mathbb{C}{})$, $F\not\equiv 0$, and at least two equilibria $a_{+},a_{-}\in\mathbb{C}{}$ of (1.1) with at least one heteroclinic orbit connecting these two equilibria. In particular, we assume that $a_{\pm}$ is not a center. Under these assumptions, it is possible to investigate topologically another characteristic set, the so-called "heteroclinic region". Taking up the questions at the end of chapter 3, a heteroclinic region is one possible structure between two adjacent global elliptic sectors. We analyse it topologically and illustrate it by an example.
Definition 5.1.
We define the heteroclinic region $\mathcal{H}$ of $F$ between $a_{+}$ and $a_{-}$ as
$\mathcal{H}:=\mathcal{H}_{+}\cap\mathcal{H}_{-}$
with
+():=$\mathcal{H}_{+(-)}:=\left\{x\in\mathbb{C}{}:w_{+(-)}(\Gamma(x))=\{a_{+(-)}\}\right\}\text{.}$
Proposition 5.2.
$\mathcal{H}$ and $\partial\mathcal{H}$ are flow-invariant. It holds
$\mathcal{H}=\bigcup_{\mathclap{x\in\mathcal{H}}}\Gamma(x)\text{.}$
Proof.
This proof is analogous to the proof of Proposition 2.2. ∎
Theorem 5.3.
$\mathcal{H}$ is open.
Proof.
Both, $a_{+}$ as well $a_{-}$ is either a focus or node or an equilibrium of order $m\in\mathbb{N}\setminus\{1\}$. Let $x\in A_{+}$ and assume that $a_{+}$ is an equilibrium of order $m\geq 2$. By [10, Proposition 4.3], there exist numbers $r,\theta,\Theta,\tau>0$ such that for all $z\in\{\hat{z}\in\mathbb{C}{}:|\hat{z}-a_{+}|<r,|\arg(\hat{z}-a_{+})-\theta|<\Theta\}=:A$ the orbit through $z$ tends to $a$ for $t\to\infty$ and $M:=\Phi(\tau,x)\in A$. Since $A$ is open, we can choose $\varepsilon>0$ small enough such that $\mathcal{B}_{\varepsilon}(M)\subset A$. By continuity of the flow, cf. [19, Chapter 2.4, Theorem 4], there exists $\delta>0$ such that $|\Phi(\tau,x)-\Phi(\tau,y)|<\varepsilon$ for all $y\in\mathcal{B}_{\delta}(x)$, i.e. $\mathcal{B}_{\delta}(x)\subset\mathcal{H}_{+}$ and $\mathcal{H}_{+}$ is open. Additionally, if $a_{+}$ is a node or focus, then we can apply Theorem 4.3 to conclude again the openness of $\mathcal{H}_{+}$. Since the same argumentation holds for $a_{-}$, $\mathcal{H}_{-}$ is also open. It follows that $\mathcal{H}=\mathcal{H}_{+}\cap\mathcal{H}_{-}$ is open. ∎
Lemma 5.4.
The basins of all centers, foci and nodes $a\in F^{-1}(\{0\})\setminus\{a_{+},a_{-}\}$ have empty intersection with $\mathcal{H}$. In particular, there are no centers, nodes and foci on $\partial\mathcal{H}$.
Proof.
There are no periodic orbits in $\mathcal{H}$. Additionally, $a\not\in\{a_{+},a_{-}\}$. Thus, this proof is analogous to the proof of Lemma 4.4. ∎
Proposition 5.5.
It holds $\partial\mathcal{H}\cap F^{-1}(\{0\})=\{a_{+},a_{-}\}$.
Proof.
Since there exists at least one heteroclinic orbit in $\mathcal{H}$ connecting $a_{+}$ and $a_{-}$, we have $\{a_{+},a_{-}\}\subset\partial\mathcal{H}$. Suppose there exists an equilibrium $a\in\partial\mathcal{H}\setminus\{a_{+},a_{-}\}$. By Lemma 5.4, $a$ is an equilibrium of order $m\geq 2$, i.e. $a$ has a FED, cf. [10, Theorem 4.4]. In particular, by [10, Definition 4.1], there exists $r>0$ such that $a\in w_{+}(\Gamma(x))\cup w_{-}(\Gamma(x))$ for all $x\in\mathcal{B}_{r}(a)$. We get $\mathcal{B}_{r}(a)\cap\mathcal{H}\not=\emptyset$, which is also impossible. Thus, we conclude $\partial\mathcal{H}\cap F^{-1}(\{0\})=\{a_{+},a_{-}\}$. ∎
Theorem 5.6.
All orbits in $\partial\mathcal{H}\setminus\{a_{+},a_{-}\}$ are unbounded. $\mathcal{H}$ is unbounded.
Proof.
If =$\partial\mathcal{H}=\{a_{+},a_{-}\}$, then $\mathcal{H}\not=\emptyset$ is open and closed with respect to the subspace topology in $\mathbb{C}{}\setminus\{a_{+},a_{-}\}$. [^12] This argument is similar to the beginning of the proof of Proposition 4.5. Since $\mathbb{C}{}\setminus\{a_{+},a_{-}\}$ is connected, =$\mathcal{H}=\mathbb{C}{}\setminus\{a_{+},a_{-}\}$ is clearly unbounded. Thus, assume $\partial\mathcal{H}\setminus\{a_{+},a_{-}\}\not=\emptyset$.
Suppose there exists a bounded orbit $\Gamma\subset\partial\mathcal{H}\setminus\{a_{+},a_{-}\}$. As in the proof of Theorem 2.6, $K:=\overline{\Gamma}$ is compact and $\Gamma\cup w_{\pm}(\Gamma)\subset K\subset\partial\mathcal{H}$. By Lemma 5.4, $\Gamma$ cannot be periodic. Hence, by Proposition 5.5 and [10, Theorem 5.1], we must have $w_{+}(\Gamma)\cup w_{-}(\Gamma)=\{a_{+},a_{-}\}$. Using this result, we want to derive a contradiction.
If $w_{+}(\Gamma)=\{a_{+}\}$ and $w_{-}(\Gamma)=\{a_{-}\}$, then we get $\Gamma\subset\mathcal{H}$, which is contraction to Theorem 5.3. If $w_{+}(\Gamma)=\{a_{-}\}$ and $w_{-}(\Gamma)=\{a_{+}\}$, then there exists another heteroclinic region $\tilde{\mathcal{H}}$ between $a_{-}$ and $a_{+}$. By Theorem 5.3, $\tilde{\mathcal{H}}$ is open, i.e. $\tilde{\mathcal{H}}\cap\mathcal{H}\not=\emptyset$, which is again a contradiction. Thus, $\Gamma$ must be homoclinic. Assume w.l.o.g. that $\Gamma$ is homoclinic in $a_{+}$, i.e. $a_{+}$ is an equilibrium of order $m\geq 2$. By Corollary 3.11, there exist two adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ such that $\Gamma(x)$ is a sector-forming orbit in $a_{+}$ with these definite directions. By Theorem 3.6, we get $\mathcal{S}(\theta_{+},\theta_{-},\Gamma(x))\cap\mathcal{H}\not=\emptyset$, i.e. there exists a homoclinic orbit in $\mathcal{H}$. This is a contradiction. We conclude that $\Gamma$ cannot be bounded. Moreover, since $\partial\mathcal{H}\not=\{a_{+},a_{-}\}$, there exists at least one unbounded orbit on $\partial\mathcal{H}$. Therefore, it follows that $\mathcal{H}$ is unbounded. ∎
Corollary 5.7.
Let $p\in\mathbb{N}\setminus\{1\}$. It is not possible to find $p$ equilibria $a_{1},\ldots,a_{p}\in F^{-1}(\{0\})$ and $p$ heteroclinic orbits $\Gamma_{1},\ldots,\Gamma_{p}$ such that $\Gamma_{p}$ is a heteroclinic orbit connecting $a_{p}$ and $a_{1}$ and $\Gamma_{j}$ is a heteroclinic orbit connecting $a_{j}$ and $a_{j+1}$ for all $j\in\{1,\ldots,p\}$.
Proof.
Suppose we find $p$ equilibria and $p$ heteroclinic orbits with these properties. Then $J:=a_{1}\cup\ldots\cup a_{p}\cup\Gamma_{1}\cup\ldots\cup\Gamma_{p}$ is a closed piecewise continuously differentiable Jordan curve. Let $\mathcal{H}_{j}$ be the heteroclinic region defined by $\Gamma_{j}$. Define the two open sets $U:=\mathcal{H}_{1}$ and $V:=\mathcal{H}_{2}\cup\ldots\cup\mathcal{H}_{p}$, cf. Theorem 5.3. Then there holds $U\cap V=\emptyset$ as well as, by Theorem 5.3, $U\cap\operatorname{Int}(J)\not=\emptyset$ and $V\cap\operatorname{Int}(J)\not=\emptyset$. Since $\operatorname{Int}(J)$ is connected, we get $U\cup V\not=\operatorname{Int}(J)$ and thus there exists $x\in\partial U\cap\operatorname{Int}(J)\not=\emptyset$. By Theorem 5.6, $\Gamma(x)\subset\partial\mathcal{H}_{1}\cap\operatorname{Int}(J)$ is unbounded. This is a contradiction. ∎
Theorem 5.8.
$\mathcal{H}$ is connected, path-connected and simply connected.
Proof.
For $x\in\mathcal{H}$ define $A_{x}:=\Gamma(x)\cup\{a_{+}\}$ and the cover $C:=\{A_{x}:x\in\mathcal{H}\}$ of $\mathcal{H}$. Since $a_{+}$ lies on the boundary of each orbit in $\mathcal{H}$, every set in $C$ is connected, cf. [18, Theorem 23.4]. Moreover, for $A,\tilde{A}\in C$ it always holds $A\cap\tilde{A}\supset\{a_{+}\}\not=\emptyset$. Thus, by [18, Theorem 23.3], $\mathcal{H}$ is connected. By [23, Proposition 12.25], $\mathcal{H}$ is even path-connected.
Let $L\subset\mathcal{H}$ be a loop. Analogous to the beginning of the proof of Theorem 4.7, we find a closed Jordan curve $J\subset\mathcal{H}$ such that $L\subset\operatorname{Int}(J)$. Suppose that $\operatorname{Int}(J)\setminus\mathcal{H}\not=\emptyset$, i.e. $\partial\mathcal{H}\cap\operatorname{Int}(J)\not=\emptyset$.
Let $y\in\partial\mathcal{H}\cap\operatorname{Int}(J)$. By Proposition 5.2, we have $\Gamma(y)\subset\partial\mathcal{H}$. Since $\mathcal{H}$ is open, cf. Theorem 5.3, we get $\Gamma(y)\cap J=\emptyset$, i.e. $\Gamma(y)\subset\operatorname{Int}(J)$. By Proposition 5.5 and Theorem 5.6, $\Gamma(y)=\{y\}$ must be an equilibrium and thus $y\in\{a_{+},a_{-}\}$. If $y$ is an equilibrium of order $m\geq 2$, we can apply Corollary 3.11 to conclude the existence of an global elliptic sector in $y$ lying completely in $\operatorname{Int}(J)$, which is a contradiction to Theorem 3.7. Hence, by Lemma 5.4, $y$ must be a node or focus. If the second equilibrium $a\in\{a_{+},a_{-}\}\setminus\{y\}$ is also a node or focus, we can apply Proposition 4.5 and Theorem 4.6 to conclude the existence of an unbounded orbit $\Gamma_{0}\subset\partial\mathcal{N}$ with $y\in w_{+}(\Gamma_{0})\cup w_{-}(\Gamma_{0})$, where $\mathcal{N}$ is the basin in $a$, cf. Definition 4.1. Thus, $\Gamma_{0}\cap\mathcal{H}\supset\Gamma_{0}\cap J\not=\emptyset$, i.e. there exists an unbounded orbit in $\mathcal{H}$, which is impossible. It follows that $a$ must be an equilibrium of order $m\geq 2$ and therefore, as above, $a\in\operatorname{Ext}(J)$.
To summarise, if $\operatorname{Int}(J)\setminus\mathcal{H}\not=\emptyset$, the only possible case is that one of the two equilibria, $a_{+}$ and $a_{-}$, is a node or focus being the only point in $\operatorname{Int}(J)\setminus\mathcal{H}$ and the other is an equilibrium of order $m\geq 2$ in $\operatorname{Ext}(J)$. To fix ideas, assume w.l.o.g. $a_{+}\in\operatorname{Int}(J)$, $a_{-}\in\operatorname{Ext}(J)$ and $\operatorname{Int}(J)\setminus\mathcal{H}=\{a_{+}\}$. Using this and a similar construction as in the proof of Proposition 4.5, we want to derive a contradiction.
First, by [10, Theorem 4.4] and [10, Definition 4.1 c) (iii)], there exists a closed piecewise continuously differentiable Jordan curve $\Gamma$ that is nowhere tangential to $F$ and satisfies $\operatorname{Int}(\Gamma)\cap F^{-1}(\{0\})=\{a_{-}\}$. Moreover, all non-elliptic orbits in $a_{-}$, i.e. all orbits not lying in a global elliptic sector in $a_{-}$, cross $\Gamma$ exactly once. Hence, all orbits in $\mathcal{H}$ cross $C_{1}:=\Gamma\cap\mathcal{H}$ exactly once. We equip $C_{1}$ with the subspace topology. Using this construction, the map $\Psi:\mathcal{H}\to\mathbb{R}$ is well-defined, when $\Psi(x)$ is the transit time131313cf. [4, Definition 3.2]. the point $x\in\mathcal{H}$ needs to $C_{1}$ along the path $\Gamma(x)$, i.e. $\{\Phi(\Psi(x),x)\}=\Gamma(x)\cap C_{1}$.
Secondly, choose $\rho>0$ such that $\mathcal{B}_{\rho}(a_{+})\subset\operatorname{Int}(J)$. By applying the theory of circles without contact around nodes and foci, cf. [2, §3, 10.-14., §7, 1.-2. and §18, Lemma 3], we find a continuously differentiable closed path $C_{2}\subset\mathcal{B}_{\rho}(a)\subset\mathcal{H}$ that is nowhere tangential to $F$ and satisfies $\operatorname{Int}(C_{2})\cap F^{-1}(\{0\})=\{a_{+}\}$. Moreover, every orbit in $\mathcal{H}$ crosses $C_{2}$ exactly once, cf. [2, §3, 10., Figure 54]. We equip $C_{2}$ with the subspace topology and define the map $\eta:C_{2}\to C_{1}$ by $\eta(x):=\Phi(\Psi(x),x)$. Analogous to the proof of Proposition 4.5, $\eta$ is continuous and bijective. Additionally, $C_{1}$ is Hausdorff and $C_{2}$ is compact, i.e. $\eta$ is a homeomorphism, cf. [18, Theorem 26.6]. Furthermore, $C_{2}$ is isomorphic to $S^{1}$. By [18, Corollary 52.5] and [18, Theorem 54.5], it follows that the fundamental group of $C_{1}$ is isomorphic to the additive group $\mathbb{Z}$.
By [10, Definition 4.1 c) (ii)], there exists a homoclinic orbit having nonempty intersection with $\Gamma$. Thus, this homoclinic orbit cannot lie in $\mathcal{H}$. Therefore, it follows that $C_{1}$ cannot be closed, i.e. the fundamental group of $C_{1}$ cannot be isomorphic to $\mathbb{Z}$. More precisely, $C_{1}$ is connected and thus homeomorphic to the simply-connected space $(0,1)\subset\mathbb{R}$, cf. [18, Theorem 23.5]. This is a contradiction. All in all, we conclude $\operatorname{Int}(J)\subset\mathcal{H}$. Since $\operatorname{Int}(J)$ is simply connected, $L$ can be continuously deformed to a point. Since $L$ is arbitrary and $\mathcal{H}$ is already path-connected, it follows that $\mathcal{H}$ is even simply connected. ∎
In [3, 4, 10] many examples are illustrating topological properties of the basins of equilibria and global elliptic sectors. [10, Example 4.8, Figure 3] demonstrates a heteroclinic region between the two equilibria. The following example shows that there can be more than one heteroclinic region between two global elliptic sectors.
Example 5.9.
Consider the vector field $F:\mathbb{C}{}\to\mathbb{C}{}$ with
$F(x):=x^{2}(x-1)(x-\mathrm{i})(x-1-\mathrm{i})\text{.}$
We have $F\in\mathcal{O}(\mathbb{C}{})$ and $F^{-1}(\{0\})=\{0,1,\mathrm{i},1+\mathrm{i}\}$. Moreover, $a_{1}=0$ is an equilibrium of order $2$, $a_{2}=1+\mathrm{i}$ is a stable node and $a_{3}=1$ as well as $a_{4}=\mathrm{i}$ are stable foci.
Figure 2. Local phase portrait of system (1.1) with $F(x)=x^{2}(x-1)(x-\mathrm{i})(x-1-\mathrm{i})$, plotted with Matlab.
We observe three heteroclinic regions between the two global elliptic sectors: the first between $a_{1}$ and $a_{2}$, the second between $a_{1}$ and $a_{3}$, and the third between $a_{1}$ and $a_{4}$. We can verify all the topological properties proven in this chapter as well as Corollary 5.7.
In this context, the question arises as to how many heteroclinic regions can lie between two global elliptic sectors. Finally, we propose the following conjecture, for which a general example is still lacking.
Conjecture 5.10.
For all $n,m\in\mathbb{N}$ there exists a complex polynomial $F_{n,m}$ of degree $n+m$ such that (1.1) has an equilibrium $a$ of order $m$ and $n$ foci or nodes $a_{1},\ldots,a_{n}$ with the following properties:
  • (i)
    Either all equilibria $a_{1},\ldots,a_{n}$ are stable or all are unstable.
  • (ii)
    For all $j\in\{1,\ldots,n\}$ there exists a heteroclinic region between $a$ and $a_{j}$.
  • (iii)
    All heteroclinic orbits of (1.1) tend to $a$ in the same definite direction.

### v3

@@
 Abstract.
-In this follow-up paper to [10] we investigate the global topology and geometry of dynamical systems $\dot{x}=F(x)$ with entire vector field $F$ by exploiting and extending the local structure of finite elliptic decompositions.
-We prove topological properties of the basins of centers, nodes, and foci, while excluding isolated equilibria at the boundaries of the latter two.
-We propose a definition of global elliptic sectors and introduce the notion of a sector-forming orbit allowing global topological analysis.
-The structure of heteroclinic regions between two equilibria is characterized.
+In this follow-up paper, we investigate the global geometry and topology of dynamical systems $\dot{x}=F(x)$ with entire vector field $F$, building on and constructively extending the local structure of simple and higher-order equilibria.
+We provide a step-by-step analysis to reveal topological properties of the basins of centers, nodes, and foci, while excluding isolated equilibria at the boundaries of the latter two.
+We propose a definition of global elliptic sectors and introduce the concept of sector-forming orbits based on the geometry within a finite elliptic decomposition of multiple equilibria.
+Finally, we characterize the structure of heteroclinic regions connecting two equilibria.
 Key words and phrases:
-Holomorphic dynamical system, global phase portrait, basin of attraction/stability, global elliptic sector, heteroclinic region
+Holomorphic dynamical system, entire vector field, center, period annulus, focus, node, basin of attraction/stability, elliptic sector, canonical region
 2020 Mathematics Subject Classification:
 Primary 30A99, 30C10, 30C15, 30D30; Secondary 32M25, 37F10, 37F75
-Acknowledgement: We thank Jörn Dietrich for mathematical discussions and collaborative solution-finding.
+Acknowledgement: We thank Jörn Dietrich for mathematical discussions and collaborative solution-finding. We thank the referee for helpful comments, suggestions, and additional literature references.
 1. Introduction
-In our previous paper [10], we focused on the local topological properties of both simple and higher-order equilibria to set the stage for investigating the global phase portrait of holomorphic vector fields in this follow-up paper.
-In order to reveal the relation between local and global dynamics, we utilize the geometric structure of finite elliptic decompositions to explore the "stability regions" of equilibria, global elliptic sectors, and heteroclinic regions.
-This is the next step towards studying holomorphic and meromorphic flows of vector fields derived from complex functions pertinent to number theory, such as the Riemann $\zeta$- and $\xi$-function, cf.
-[5, 6, 11, 21].
-The boundary-forming orbits (separatrices) of these "stability regions" are particularly interesting since they divide the complex plane into regions that exhibit completely distinct topological behavior, reminiscent of phase transitions.
-In previous works, we have been analyzing normally attracting/repelling (slow) invariant manifolds of equilibria in the context of multi scale dynamical systems, cf.
-[8, 12, 13, 14, 15, 16, 17], which, in the 2-dimensional planar case, are usually separatrices dividing the phase space into topological basins.
-In the following, we provide an outline of our paper. We consider the special case of entire vector fields in real time parameterization, i.e. a holomorphic dynamical system of the form
+In our previous work [19], we examined the local structure of simple and higher-order equilibria in holomorphic vector fields, i.e. in holomorphic dynamical systems of the form
 (1.1) $\dot{x}=\frac{\mathrm{d}x}{\mathrm{d}t}=F(x),\quad x\in\mathbb{C}{},\;t\in\mathbb{R}$
 with $F\in\mathcal{O}(\mathbb{C}{})$.
-[^1] In the context of dynamical systems, we use the notation of orbits and flows given in [10, Chapter 2].
-A definition of limit sets, cf.
-[2, §4] and [20, Chapter 8.4], as well as some basic results can also be found in [10, Chapter 2].
-We start with the analysis of the basin of a center, cf.
-[4, Theorem 4.1] and [3, Theorem 3.3], and provide detailed proofs of several topological properties.
-In particular, we detail the proof in [3, Theorem 3.3] for the openness of the basin of a center.
-Subsequently, we examine the topological structure of equilibria of order $m\geq 2$.
-Since a precise definition of global elliptic sectors is missing in the literature222cf.
-[2, Chapter VIII, §17], [4, Chapter 4], [7, Chapter 1.5] and [19, Chapter 2.11]., we propose an appropriate characterization by introducing the concept of sector-forming orbits.
-We demonstrate that the intuitive idea of the geometric structure of a global elliptic sector is indeed rigorously valid.
-We support the proof of the openness with geometric visualization, cf.
-Figure 1, and point out the conditions under which the local structure of a finite elliptic decomposition can be transferred to the global phase portrait, an issue already addressed in [10, Example 4.7, Example 4.8].
+Building on these local insights, this follow-up paper shifts the perspective to the global phase portrait of entire vector fields and develops an understanding of how local structures influence and shape the global geometry of (1.1).
+Throughout this paper, we use the terminology and notion of orbits and flows in the context of dynamical systems and the Jordan curve Theorem from [19, Chapter 2].
+In particular, if $\Gamma\subset\mathbb{C}{}$ is a closed Jordan curve, we denote the two connecting components resulting from the Jordan curve Theorem by $\operatorname{Int}(\Gamma)$ (the bounded interior of $\Gamma$) and $\operatorname{Ext}(\Gamma)$ (the unbounded exterior of $\Gamma$).
+As examined in [19], each equilibrium of (1.1) falls into one of the following categories:
+  • (i)
+    A center (simple equilibrium), where all orbits in a neighborhood are closed periodic orbits surrounding the center.
+  • (ii)
+    A focus or node (simple equilibrium), which is attracting or repelling such that all nearby orbits tend to the equilibrium either in positive (attracting) or negative (repelling) time.
+  • (iii)
+    A multiple equilibrium of order $m\in\mathbb{N}\setminus\{1\}$, possessing a finite elliptic decomposition of order 2m22m-2.
+This geometric structure is defined in [19, Definition 4.1] and illustrated in Section 3.1.
+With this in mind, we focus on the following two central questions to reveal the relation between local and global dynamics:
+  • (A)
+    How can we globally define the "region of influence" of an equilibrium based on its known local structure, and what are the topological properties of the resulting set?
+  • (B)
+    How can these "regions of influence" and their geometric structure be embedded into the existing theory of canonical (strip, annular, radial, and spiral) regions in general smooth planar dynamical systems, cf.
+[8, 9, 10, 11]?
+For each of the three cases (i)–(iii) described above, we provide a step-by-step analysis to address questions (A) and (B).
+In addition, we emphasize the significance and specific influence of the holomorphy of the vector field in (1.1) on the possible geometric configurations of the global phase space.
+This becomes apparent both in the results stated in theorems and in the arguments within proofs.
+In particular, the central tool is [19, Corollary 5.1], which can be formulated in the case of entire vector fields as follows:
+Theorem 1.1 (Poincaré-Bendixson for Holomorphic Flows).
+Let $F\in\mathcal{O}{}(\mathbb{C})$, $F\not\equiv 0$, be an entire function and $K\subset\mathbb{C}$ compact with $\xi\in K$ such that $\Gamma_{+(-)}(\xi)\subset K$.
+Then either $\Gamma(\xi)$ is a periodic orbit with exactly one equilibrium, a center, in its interior, or the positive (negative) limit set of $\Gamma(\xi)$ consists of exactly one equilibrium.
+Additionally, if $\Gamma(\xi)$ is periodic, the interior of $\Gamma(\xi)$ (except for the center) is filled entirely with periodic orbits all having the center in its interior.
+The analysis of questions (A) and (B) also sets the stage for studying holomorphic and meromorphic flows of vector fields derived from complex functions pertinent to number theory, such as the Riemann $\zeta$- and $\xi$-function [6, 7, 20, 21, 31] and their polynomial approximations.
+Some ideas are motivated by previous works, in which we analyzed normally attracting or repelling (slow) invariant manifolds of equilibria within the context of multiscale dynamical systems [17, 22, 23, 24, 25, 26, 27].
+In the following, we provide an outline of our paper.
+We start with a short chapter about the center basin, also called period annulus.
+We use the well-known theory about closed orbits examined in [2, 9, 11] in order to provide complete proofs of several topological properties that are already stated in [5, Theorem 4.1] and [4, Theorem 3.3].
+Subsequently, as the main focus of this paper, we examine the topological structure of equilibria of order $m\geq 2$.
+Since a precise definition of a global elliptic sector is lacking in the literature, cf.
+[2, 5, 13, 14, 16, 29], we introduce the concept of sector-forming orbits to provide an appropriate global characterization.
+We show that the intuitive notion of the geometric structure of a global elliptic sector can indeed be made rigorous.
+We conclude this section by showing that, in the sense of Markus and Neumann [8, 9], it naturally forms a strip canonical region, thus addressing question (B).
 Building on these results, we analyze the topology of the basin of nodes and foci.
-Notably, the authors of [4, Theorem 4.3] explicitly do not exclude the scenario of an isolated equilibrium on the boundary of the basin – specifically, an equilibrium that is not attached to an orbit lying outside the basin.
-We address this aspect in detail in Proposition 4.5.
-Finally, we define and investigate heteroclinic regions between two equilibria, observing properties similar to the preceding cases.
+The authors of [5, Theorem 4.3] explicitly do not exclude the scenario of an isolated equilibrium on the boundary of the basin.
+We address this aspect in detail in Proposition 4.4 with a rigorous proof in the appendix.
+Finally, we define and investigate heteroclinic regions as one geometric possibility between two adjacent elliptic sectors, observing topological properties similar to the preceding cases.
 This leads to the exclusion of the existence of a cycle in the phase space formed by equilibria and heteroclinic orbits connecting these equilibria, cf.
 Corollary 5.6.
+We conclude this paper with an example that illustrates the occurrence of several elliptic sectors and heteroclinic regions together with their separatrices in the phase space.
+Throughout the paper, we selectively focus on explicit constructions that extend local structures to global ones, without a priori resorting to potentially known or conjectured global results, because we consider constructive proofs to offer a fruitful setting for developing general approaches to questions of type (A) and (B).
+Nevertheless, we aim to contextualize our framework within previous research from different but related perspectives.
 2. Basin of centers
-In this chapter, we treat the case of a center.
-We provide more detailed complete proofs of topological results given in [4, Theorem 4.1] and [3, Theorem 3.3].
-In all considerations of this chapter we consider an entire function $F\in\mathcal{O}{}(\mathbb{C}{})$, $F\not\equiv 0$, and a center $a\in\mathbb{C}{}$ of (1.1).
+Throughout this chapter, we consider an entire function $F\in\mathcal{O}{}(\mathbb{C}{})$, $F\not\equiv 0$, and a center $a\in\mathbb{C}{}$ of (1.1).
 Definition 2.1.
-We define the basin $\mathcal{V}$ of $F$ in $a$ as
+We define the center basin or period annulus $\mathcal{V}$ of $F$ in $a$ as
 $\mathcal{V}:=\{a\}\cup\left\{x\in\mathbb{C}:\Gamma(x)\text{ is periodic with }a\in\operatorname{Int}(\Gamma)\right\}.$
 The boundary of $a$ is defined as the boundary of $\mathcal{V}$.
+A well-established theory exists regarding the local and global geometry of the period annulus $\mathcal{V}$.
+Markus and Neumann, cf.
+[8, 9, 11], provide a thorough analysis of annular (parallel) canonical regions.
+In the following, we highlight which of the known properties extend to the holomorphic setting and examine how holomorphy further refines both the global geometry and the proof techniques.
+This approach leads to complete proofs of the topological results stated in [5, Theorem 4.1] and [4, Theorem 3.3].
 Proposition 2.2.
@@
 Proof.
-[10, Definition 3.1] ensures $a\in\mathring{\mathcal{V}}$, i.e.
-$\mathring{\mathcal{V}}\not=\emptyset$.
-Let $y\in\mathcal{V}$ and $z\in\Gamma(y)$.
-Then $\Gamma(y)=\Gamma(z)\subset\mathcal{V}$.
-Thus $\mathcal{V}$ is flow-invariant.
-Every $y\in\operatorname{Int}(\Gamma(x))$ either satisfies $y=a$, or, by [10, Theorem 5.1], is a periodic orbit with $a$ in its interior.
-This proves equation (2.1).
-By [22, Lemma 6.4], it follows that also $\partial\mathcal{V}$ is flow-invariant.
+Definition 3.1 in [19] ensures $a\in\mathring{\mathcal{V}}\not$$\emptyset$.
+The flow-invariance of $\mathcal{V}$ and $\partial\mathcal{V}=\overline{\mathcal{V}}\cap\overline{\mathbb{C}\setminus\mathcal{V}}$ follow from [32, Lemma 6.4].
+By using the holomorphy of $F$, we can apply Theorem 1.1 to conclude that all orbits in $\operatorname{Int}(\Gamma(x))$, with $x\in\mathcal{V}$, are periodic with $a$ in the interior, which shows (2.1).
 ∎
+This Proposition shows that $\mathcal{V}$ is indeed a parallel annular region, cf.
+[11, p.
+90].
+In general, equation (2.1) does not hold for arbitrary planar dynamical systems.
+In fact, there are examples where one or more limit cycles appear within closed orbits, with spiraling behavior in a neighborhood of the limit cycle.
+However, holomorphy rules out the existence of limit cycles and similar phenomena within periodic orbits, cf.
+[4, Theorem 3.2].
+The following result demonstrates that, once again due to holomorphy of $F$, no equilibria can lie on the boundary of $a$.
 Proposition 2.3.
-It holds $\partial\mathcal{V}\cap F^{-1}(\{0\})=\emptyset$, i.e. there are no equilibria on the boundary of $\mathcal{V}$.
+It holds $\partial\mathcal{V}\cap F^{-1}(\{0\})=\emptyset$.
 Proof.
@@
 Hence, $x$ cannot be a center, focus or node.
-By [10, Theorem 4.4], $x$ has a FED.
-In particular, by [10, Definition 4.1], there exists $r>0$ such that $x\in w_{+}(\Gamma(z))\cup w_{-}(\Gamma(z))$ for all $z\in\mathcal{B}_{r}(x)$.
-But no orbits in $\mathcal{B}_{r}(x)\cap\mathcal{V}\not=\emptyset$ tend to $x$, i.e.
-this is a contradiction to [10, Definition 4.1 c) (ii)].
+By using the holomorphy of $F$, we can apply [19, Theorem 4.4] to conclude that $x$ has a FED.
+In particular, by [19, Definition 4.1], there exists $r>0$ such that $x\in w_{+}(\Gamma(z))\cup w_{-}(\Gamma(z))$ for all $z\in\mathcal{B}_{r}(x)$.
+But the orbits in $\mathcal{B}_{r}(x)\cap\mathcal{V}\not$$\emptyset$ do not tend to $x$, contradicting the property c) (ii) in [19, Definition 4.1].
 Thus, $x$ cannot be an equilibrium.
@@
 Theorem 2.4.
-$\mathcal{V}$ is open.
+$\mathcal{V}$ is open and unbounded. All orbits on $\partial\mathcal{V}$ are unbounded.
 Proof.
-Suppose there exists $x\in\mathcal{V}\cap\partial\mathcal{V}\not=\emptyset$, $x\not=a$.
-Then $\Gamma(x)\subset\mathcal{V}\cap\partial\mathcal{V}$, cf.
-Proposition 2.2.
-In particular, $\Gamma(x)$ is a periodic orbit with $a\in\operatorname{Int}(\Gamma(x))\not=\emptyset$.
-Suppose there also exists $y\in\partial\mathcal{V}\setminus\Gamma(x)\not=\emptyset$.
-The set $\mathbb{C}\setminus\Gamma(x)=\operatorname{Int}(\Gamma(x))\cup\operatorname{Ext}(\Gamma(x))$ is open.
-If $y\in\operatorname{Int}(\Gamma(x))$, there exists $r>0$ such that $\mathcal{B}_{r}(y)\subset\operatorname{Int}(\Gamma(x))$.
-By equation (2.1), we get $\mathcal{B}_{r}(y)\subset\mathcal{V}$, i.e.
-$y\not\in\partial\mathcal{V}$.
-This a contradiction.
-Furthermore, if $y\in\operatorname{Ext}(\Gamma(x))$, there exists again an $r>0$ such that $\mathcal{B}_{r}(y)\subset\operatorname{Ext}(\Gamma(x))$.
-Since $y$ lies on the boundary of $a$, there exists $z\in\mathcal{B}_{r}(y)\cap\mathcal{V}$, i.e.
-$\Gamma(z)$ is a periodic orbit with $a\in\operatorname{Int}(\Gamma(z))$ and $\Gamma(z)\not=\Gamma(x)$.
-Since $\Gamma(z)$ is path-connected and $z\in\operatorname{Ext}(\Gamma(x))$, we must have $x\in\operatorname{Int}(\Gamma(z))$.
-By equation (2.1) and openness of $\operatorname{Int}(\Gamma(z))$, we derive the contradiction $x\in\mathring{\mathcal{V}}$.
-Thus, $y\not\in\operatorname{Int}(\Gamma(x))\cup\operatorname{Ext}(\Gamma(x))$, which is also a contradiction.
-We get $\partial\mathcal{V}\setminus\Gamma(x)=\emptyset$, i.e.
-the boundary of $a$ consists exactly of the unique outermost periodic orbit $\Gamma(x)$ and $\mathcal{V}=\overline{\operatorname{Int}(\Gamma(x))}$ must be compact.
-By equation (2.1), $\Gamma(x)$ is not isolated.
-Hence we can apply a result on limit cycles of analytic dynamical systems, [2, p.
-203, §12.1, Lemma 1], to deduce the existence of $\rho>0$ such that all orbits through the set $U:=\{y\in\mathbb{C}{}:\text{dist}(\Gamma(x),y)<\rho\}\cap\operatorname{Ext}(\Gamma(x))$ are periodic.
-Reduce $\rho$ such that $\overline{\operatorname{Int}(\Gamma(x))}\cup U$ is simply connected and $U\cap F^{-1}(\{0\})=\emptyset$.
-By continuity of the flow, cf.
-[19, Chapter 2.4, Theorem 4], we deduce the existence of at least one periodic orbit $\Gamma\subset U$.
-By choice of $\rho$, we get $\operatorname{Int}(\Gamma)\subset U\cup\overline{\operatorname{Int}(\Gamma(x))}$ and thus $a\in\operatorname{Int}(\Gamma)$, cf.
-[2, §11, Corollary 1] and [3, Theorem 2.4].
-This is a contradiction to $\Gamma\subset\mathbb{C}{}\setminus\mathcal{V}$.
-Hence the assumed $x$ cannot exist and $\mathcal{V}$ is open.
+Since $\mathcal{V}$ consists of closed periodic orbits, we can use the already known theory about annular regions of analytic planar dynamical systems.
+In [2, §12.1, p.
+203], it is shown that $\partial\mathcal{V}$ consists of equilibria as well as separatrices of saddle points.
+However, the holomorphy of $F$ prevents the existence of saddles, cf.
+[19, Theorem 3.2].
+Hence, by Proposition 2.3, $\partial\mathcal{V}$ consists only of unbounded separatrices tending to infinity in both time directions.
+In particular, each of these separatrices connects two saddles at Infinity, cf.
+[12], and is in particular not a periodic orbit, i.e.
+$\mathcal{V}\cap\partial\mathcal{V}=\emptyset$.
 ∎
+We have thus verified that $\mathcal{V}$ is a canonical region whose boundary consists of separatrices, cf.
+[9, Theorem 5.2].
+Furthermore, there exists a global smooth transversal in $\mathcal{V}\setminus\{a\}$, that is, an open smooth arc crossing each closed orbit in $\mathcal{V}$ exactly once.
+The definition of annular canonical regions leads immediately to the following result.
+However, we present a alternative short proof by applying a version of the Seifert–Van Kampen theorem, combined with the results obtained above.
 Theorem 2.5.
@@
 Thus, also the union must be connected, cf.
-[18, Theorem 23.3].
-By [23, Proposition 12.25], $\mathcal{V}$ is even path-connected.
-Let $L\subset\mathcal{V}$ be a loop (closed path).
-By openness of $\mathcal{V}$, cf.
-Theorem 2.4, for all $x\in L$ there exists a $y_{x}\in\operatorname{Ext}(\Gamma(x))\cap\mathcal{V}\not=\emptyset$ such that $x\in\operatorname{Int}(\Gamma(y_{x}))\subset\mathcal{V}$.
-Hence, the system $\left\{\operatorname{Int}(\Gamma(y_{x})):x\in L\right\}$ is an open cover of $L$.
-Since $L$ is compact, there exist finitely many points $x_{1},\ldots,x_{\ell}\in L$, $\ell\in\mathbb{N}{}$, such that
-$L\subset\bigcup_{k=1}^{\ell}\operatorname{Int}(\Gamma(y_{x_{k}}))\subset\mathcal{V}\text{.}$
-Since $\ell$ is finite, $y\in\left\{y_{x_{k}}:1\leq k\leq\ell\right\}$ must exist such that $\Gamma(y)$ is the outermost periodic orbit having the other $\ell-1$ periodic orbits in its interior.
-By equation (2.1), we conclude $L\subset\operatorname{Int}(\Gamma(y))\subset\mathcal{V}$ with $y\in\mathcal{V}$.
-Since $\operatorname{Int}(\Gamma(y))$ is clearly simply connected, $L$ can be continuously (homotopically) deformed to a point in $L\subset\mathcal{V}$.
-Since $\mathcal{V}$ is already path-connected, it follows that $\mathcal{V}$ is simply connected.
+[28, Theorem 23.3].
+By [33, Proposition 12.25], $\mathcal{V}$ is even path-connected.
+By openness of $\mathcal{V}$ we get
+$\mathcal{V}=\bigcup_{\mathclap{x\in\mathcal{V}}}\operatorname{Int}(\Gamma(x))\text{.}$
+All the sets $\operatorname{Int}(\Gamma(x))$, with $x\in\mathcal{V}$, are open and simply connected.
+Hence, by applying the Seifert-van Kampen theorem for infinite open covers, cf.
+[15, Theorem 1.20], we conclude that also $\mathcal{V}$ is simply connected.
 ∎
-Theorem 2.6.
-All orbits on the boundary of $a$ are unbounded. $\mathcal{V}$ is unbounded.
+We conclude this chapter with a lemma that will play an important role at several points in what follows.
+Lemma 2.6.
+Let $A\subset\mathbb{C}$ be a flow-invariant set that admits no closed orbits. Then $\mathcal{V}\cap\partial A=\emptyset$.
 Proof.
-If the boundary of $a$ is empty, then $\mathcal{V}$ is open, closed and nonempty, i.e.
-$\mathcal{V}=\mathbb{C}{}$ and thus unbounded.
-Assume $\partial\mathcal{V}\not=\emptyset$ and let $x\in\partial\mathcal{V}$.
-Suppose that $\Gamma(x)\subset\partial\mathcal{V}$ is bounded.
-By Proposition 2.3, we get $w_{\pm}(\Gamma(x))\cap F^{-1}(\{0\})=\emptyset$.
-Hence, by [10, Corollary 5.2], $\Gamma(x)$ must be periodic.
-By [10, Theorem 5.1], $\Gamma(x)$ has exactly one equilibrium, a center $\tilde{a}$, in $\operatorname{Int}(\Gamma(x))$.
-If $\tilde{a}\not=a$, $x\in\partial\mathcal{V}$ would be a point in the basin in $\tilde{a}$, which is a contradiction to the openness of this basin, cf.
-Theorem 2.4.
-If $a=\tilde{a}$, we get $x\in\mathcal{V}$, which is again a contradiction, since $\mathcal{V}\cap\partial\mathcal{V}=\emptyset$.
-Thus, $\Gamma(x)$ and $\mathcal{V}$ must be unbounded.
-∎
+If the set $\mathcal{V}\cap\partial A$ was not empty, it would contradict the openness of $\mathcal{V}$, cf. Theorem 2.4. ∎
 3. Global elliptic sectors
-In all claims and arguments of this chapter, we consider an entire function $F\in\mathcal{O}{}(\mathbb{C})$, $F\not\equiv 0$, and an equilibrium $a\in\mathbb{C}$ of (1.1) of order $m\in\mathbb{N}\setminus\{1\}$.
-By applying [10, Proposition 4.3] and [10, Theorem 4.4], (1.1) has a finite elliptic decomposition (FED) of order $22m-$ 2 in $a$, where all local elliptic sectors have adjacent definite directions given by $\mathcal{E}(F,m)$.
-Moreover, we use the function $\lambda:\mathcal{E}(F,m)\to\{-1,1\}$, $\lambda(\theta):=\cos(\operatorname{arg}(F^{(m)}(a))+\theta(m-1))$ to determine the time direction (positive or negative) in which the orbits tend to $a$ in the FED, cf.
-[10, Proposition 4.3 (iii)] and [10, calculation of $\lambda_{i}$ on p.
-11].
-In the following, we will extend the analysis of the local structure described in [10, Definition 4.1 c)] to a global point of view and provide detailed complete proofs of the topological results given in [4, Theorem 4.2].
-To do this, we need at first a precise definition of a global elliptic sector.
+In this chapter, we consider an entire function $F\in\mathcal{O}{}(\mathbb{C})$, $F\not\equiv 0$, and an equilibrium $a\in\mathbb{C}$ of (1.1) of order $m\in\mathbb{N}\setminus\{1\}$.
+Our aim is to define and understand the global structure of the "areas of influence" associated with such an equilibrium.
+We begin by providing a brief overview of earlier research on this topic.
+3.1. Survey of related research
+In [2, §17] and [16, Chapter 1.5], the authors describe and analyze several types of curvilinear sectors in general planar dynamical systems, providing detailed proofs of their geometric properties.
+A global analysis of planar dynamical systems is given in [8, 9, 10, 11], where canonical regions and global transversals are classified.
+The structure of multiple equilibria is also discussed there.
+However, these works do not specifically address the special case of holomorphic vector fields in (1.1).
+In [13], the authors classify holomorphic flows up to conformal equivalence in a neighborhood of an equilibrium, providing representatives for each equivalence class.
+This yields several model flows together with their corresponding local phase portraits.
+A related investigation is carried out in [3], where the authors describe various normal forms depending on the order and residue of the equilibrium.
+For example, the local structure of an equilibrium $a$ of order $m\in\mathbb{N}\setminus\{1\}$ is conformally equivalent to (1.1) with
+$F(x):=x^{n}-cx^{2n-1}\text{,}$
+where $c\in\mathbb{C}{}$ is the residuum of $\frac{1}{F}$ at $a$. However, all these analyses concern the local phase portrait.
+We have likewise analyzed the local structure of multiple equilibria in our recent paper [19], where we established the existence of a finite elliptic decomposition (FED) of order 2m22m-2 in the multiple equilibrium $a$, cf.
+[19, Proposition 4.3] and [19, Theorem 4.4].
+In particular, we showed that each local elliptic sector of this decomposition has adjacent definite directions given by
+$\mathcal{E}(F,m)=\left\{\frac{\ell\pi-\arg(F^{(m)}(a))}{m-1}\mod 2\pi:\ell\in\mathbb{Z}\right\}\subset[0,2\pi)\text{.}$
+The geometry of a local elliptic sector is defined in [19, Definition 4.1 c)] and illustrated in Figure 1.
+A finite elliptic decomposition can be obtained by cyclically copying the geometry in Figure 1 around the equilibrium.
+In the proof of [19, Theorem 4.4], we established step by step the various geometric objects required for a finite elliptic decomposition without invoking the theory of topological/conformal equivalence.
+[…]
+Figure 1.
+Geometrical objects of a local elliptic sector (blue) in a multiple equilibrium $a$.
+$\Gamma_{1}$ and $\Gamma_{2}$ (black) are the characteristic orbits.
+$\Lambda_{1}$ and $\Lambda_{2}$ (red) are the transversals or paths without contact.
+$\Xi=\Gamma(E_{1})$ (green) is the homoclinic orbit.
+The arrows indicate the direction of the vector field.
+With this in mind, we now extend the analysis of the local structure to a global perspective by introducing the notion of a global elliptic sector.
+We provide detailed complete proofs of the topological results given in [5, Theorem 4.2].
+Furthermore, we embed this global elliptic sector as a canonical region into the global theory of Markus and Neumann, cf.
+[9, 10, 11].
+3.2. The geometry of global elliptic sectors
+In this chapter, we provide a step-by-step analysis of the global elliptic sector to obtain a complete topological and geometrical characterization.
+We introduce the notion of the global elliptic sector in an intuitive manner by "blowing up" the local structure in an appropriate way.
 Definition 3.1.
   • (i)
-    Let $\Gamma\subset\mathbb{C}$ be an orbit and $x\in\Gamma$.
-We call $\Gamma$ $(\theta_{+},\theta_{-})$-elliptic in $a$, if there exists $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ such that $w_{+}(\Gamma)=w_{-}(\Gamma)=\{a\}$, $\arg(\Phi(t,x)-a)=\theta_{+}$ for $t\to\infty$ and $\arg(\Phi(t,x)-a)=\theta_{-}$ for $t\to-\infty$, i.e.
-$\Gamma$ is homoclinic and tends to $a$ in in the definite direction $\theta_{\pm}$ for $t\to\pm\infty$.
+    Let $\Xi\subset\mathbb{C}\setminus\{a\}$ be a homoclinic orbit in $a$, i.e.
+$w_{+}(\Gamma)=w_{-}(\Gamma)=\{a\}$.
+$\Xi$ is a sector-forming orbit in $a$, if for all $z\in\operatorname{Int}(\Xi\cup\{a\})$ the orbit $\Gamma(z)$ is also homoclinic in aa.
+[^1] A construction of a parameterization for the closed Jordan curve $\Xi\cup\{a\}$ with compact time interval can be found in [19, Remark 4.2].
   • (ii)
-    Let $\Xi\subset\mathbb{C}$ be an orbit, which is $(\theta_{+},\theta_{-})$-elliptic in $a$.
-We call $\Xi$ a sector-forming orbit in $a$ with definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, if for all $y\in\operatorname{Int}(\Xi\cup\{a\})$ the orbit $\Gamma(y)$ is also $(\theta_{+},\theta_{-})$-elliptic in $a$.
+    Let $\Xi$ be a sector-forming orbit in $a$. The global elliptic sector $\mathcal{S}(\Xi)$ of $F$ in $a$ with respect to $\Xi$ is
+    $\mathcal{S}(\Xi):=\underbrace{\Xi\cup\operatorname{Int}(\Xi\cup\{a\})}_{=\overline{\operatorname{Int}(\Xi\cup\{a\})}\setminus\{a\}}\cup\,\mathcal{S}^{\prime}(\Xi)$
+    with
+    $\mathcal{S}^{\prime}(\Xi):=\big{\{}x\in\mathbb{C}:\,$ $\Gamma(x)\text{ is homoclinic in }a,$
+    $\Xi\subset\operatorname{Int}(\Gamma(x)\cup\{a\})\big{\}}\text{.}$
   • (iii)
-[…]
-    $\mathcal{S}(\theta_{+},\theta_{-},\Xi):=\underbrace{\Xi\cup\operatorname{Int}(\Xi\cup\{a\})}_{=\overline{\operatorname{Int}(\Xi\cup\{a\})}\setminus\{a\}}\cup\,\mathcal{S}^{\prime}(\theta_{+},\theta_{-},\Xi)$
-    with
-    $\mathcal{S}^{\prime}(\theta_{+},\theta_{-},\Xi):=\big{\{}x\in\mathbb{C}:\,$ $\Gamma(x)\text{ is }(\theta_{+},\theta_{-})\text{-elliptic in }a,$
-    $\Xi\subset\operatorname{Int}(\Gamma(x)\cup\{a\})\big{\}}\text{.}$
-Remark 3.2.
-Let $\Gamma\subset\mathbb{C}$ be $(\theta_{+},\theta_{-})$-elliptic in $a$.
-A construction of a parameterization for the closed Jordan curve $\Gamma\cup\{a\}$ with compact time interval can be found in [10, Remark 4.2].
-Hence, all sets in Definition 3.1 are indeed well-defined.
-The global elliptic sector in the constructive Definition 3.1 (iii) depends on the choice of the sector-forming orbit.
+    Let $\Gamma_{1},\Gamma_{2}$, be two homoclinic orbits in $a$ and $\Gamma_{1}\not=\Gamma_{2}$.
+$\Gamma_{1}$ and $\Gamma_{2}$ are called nested if either $\Gamma_{1}\subset\operatorname{Int}(\Gamma_{2}\cup\{a\})$ or $\Gamma_{2}\subset\operatorname{Int}(\Gamma_{1}\cup\{a\})$.
+$\Gamma_{1}$ and $\Gamma_{2}$ are called mutually exterior if they are not nested, i.e.
+$\Gamma_{1}\subset\operatorname{Ext}(\Gamma_{2}\cup\{a\})$ and $\Gamma_{2}\subset\operatorname{Ext}(\Gamma_{1}\cup\{a\})$.
+[^2] This terminology is in the sense of [2, §17, 3.].
+Apriori the global elliptic sector in this constructive Definition depends on the choice of the sector-forming orbit.
 We will later see the motivation for this type of construction and demonstrate that the global elliptic sector is in an appropriate sense independent of the choice of a particular sector-forming orbit – as can be intuitively deduced from the idea of the structure of an equilibrium of order $m\geq 2$, cf.
-[10, Figure 2, Figure 3].
+Figures 2 and 3 in [19].
 Lemma 3.2.
-Let $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ be two adjacent definite directions with $\lambda(\theta_{\pm})=\mp 1$.
-Then there exists a sector-forming orbit $\Xi$ in $a$ with definite directions $\theta_{+},\theta_{-}$.
+There exists a sector-forming orbit $\Xi$ in $a$.
 Proof.
-By [10, Theorem 4.4], there exists a local elliptic sector $S$ between $\theta_{+}$ and $\theta_{-}$.
-Moreover, the two characteristic orbits of $S$ tend to $a$ in the directions $\theta_{+}$ and $\theta_{-}$, cf.
-[10, Definition 4.1 c)] and [10, Proposition 4.3 (ii)].
-Hence, we can choose the orbit $\Xi:=\Gamma(E_{1})\subset S$, where $E_{1}\in S$ is the point given by [10, Definition 4.1 c)].
-In fact, all orbits in $\overline{\operatorname{Int}(\Xi\cup\{a\})}\setminus\{a\}\subset S$ lie between the two characteristic orbits of $S$ and thus are $(\theta_{+},\theta_{-})$-elliptic in $a$, cf.
-[10, Definition 4.1 c) (ii), (iv)].
+Since there exists a FED in $a$, we can choose two adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ and the orbit $\Xi:=\Gamma(E_{1})$, where $E_{1}$ is the point given by [19, Definition 4.1 c)], cf.
+Figure 1.
+Then $w_{+}(\Xi)=w_{-}(\Xi)=\{a\}$, $\arg(\Phi(t,E_{1})-a)=\theta_{+}$ for $t\to\infty$ and $\arg(\Phi(t,E_{1})-a)=\theta_{-}$ for $t\to$-$\infty$, i.e.
+$\Xi$ is homoclinic and tends to $a$ in the definite direction $\theta_{\pm}$ for $t\to\pm\infty$.
+This orbit is part of a local elliptic sector $S$ between $\theta_{+}$ and $\theta_{-}$.
+In addition, all orbits in $\overline{\operatorname{Int}(\Xi\cup\{a\})}\setminus\{a\}\subset S$ are homoclinic in $a$, cf.
+[19, Definition 4.1 c)].
 ∎
-The above Lemma ensures the existence of a sector-forming orbit for adjacent definite directions.
-In the following propositions and theorems we use this construction to show various topological properties of global elliptic sectors such as invariance, openness, unboundedness, connectedness.
+The above Lemma ensures the existence of a sector-forming orbit between two arbitrarily chosen adjacent definite directions.
+In the following Propositions and Theorems we present various topological properties of the global elliptic sector.
 Proposition 3.3.
-Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$.
+Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$.
 $\mathring{\mathcal{S}}$ is nonempty.
@@
 Proof.
-By choosing a point in $\operatorname{Int}(\Xi\cup\{a\})$, we have $\mathring{\mathcal{S}}\not=\emptyset$. The rest of the proof is analogous to the proof of Proposition 2.2. ∎
+By choosing a point in $\operatorname{Int}(\Xi\cup\{a\})$, we have $\mathring{\mathcal{S}}\not$$\emptyset$. The rest of the proof is analogous to the proof of Proposition 2.2. ∎
 Proposition 3.4.
-Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$.
-It holds $\partial\mathcal{S}\cap F^{-1}(\{0\})=\{a\}$.
+Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$. It holds $\partial\mathcal{S}\cap F^{-1}(\{0\})=\{a\}$.
 Proof.
@@
 Since the orbits in $\mathcal{S}$ all tend to $a\not=x$ in both time directions and are not periodic, $x$ cannot be a center, focus or node.
-Hence, as in the proof of Proposition 2.3, $x$ has a FED, cf.
-[10, Theorem 4.4].
-Moreover, by [10, Definition 4.1], there exists $r>0$ such that $a\in w_{+}(\Gamma(x))\cup w_{-}(\Gamma(x))$ for all $x\in\mathcal{B}_{r}(a)$.
-In particular, we get $\mathcal{B}_{r}(a)\cap\mathcal{S}\not=\emptyset$, which is a contradiction.
+Hence, as in the proof of Proposition 2.3, $x$ has a FED and there exists $r>0$ such that $x\in w_{+}(\Gamma(z))\cup w_{-}(\Gamma(z))$ for all $z\in\mathcal{B}_{r}(x)$ due to the holomorphy of $F$, cf.
+[19, Theorem 4.4].
+Since $\mathcal{B}_{r}(x)\cap\mathcal{S}\not$$\emptyset$, we get the contradiction $x=a$.
 Thus, $x$ cannot be an equilibrium.
@@
 ∎
+Lemma 3.5.
+Let $\Gamma$ be a homoclinic orbit in $a$ and $x\in\Gamma$.
+Let $\varepsilon>0$.
+Then there exists $\delta>0$ such that all orbits $\hat{\Gamma}$ through $\mathcal{B}_{\delta}(x)$ satisfy the following properties:
+  • (i)
+    $\hat{\Gamma}$ is homoclinic in $a$.
+  • (ii)
+    There exists a set $H\subset\mathcal{U}_{\varepsilon}(\Gamma)$ such that $\partial H=\Gamma\cup\hat{\Gamma}\cup\{a\}$. [^3] The set $H$ looks similar to a crescent moon ☾.
+  • (iii)
+    If $\Gamma\not=\hat{\Gamma}$, then $\Gamma$ and $\hat{\Gamma}$ are nested.
+Proof.
+Due to the holomorphy of $F$, we can apply [19, Proposition 4.3 (ii)] to conclude the existence of two angles $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ such that $\Gamma$ tends to $a$ in the definite directions $\theta_{+}$ and $\theta_{-}$.
+By [19, Proposition 4.3 (iii)] there exist $r,\Theta_{\pm}>0$ such that all orbits through the open set $\{z\in\mathbb{C}{}:|z-a|<r,|\arg(z-a)-\theta_{\pm}|<\Theta_{\pm}\}=:A_{\pm}$ tend to $a$ in the definite direction $\theta_{\pm}$ for $t\to\pm\infty$, i.e.
+are in particular homoclinic.
+Moreover, there exists $\tau>0$ such that $\Phi(\pm\tau,x)\in A_{\pm}$.
+Choose $\rho>0$ small enough such that $\mathcal{B}_{\rho}(\Phi(\pm\tau,x))\subset A_{\pm}$.
+By continuity of the flow, cf.
+[29, Chapter 2.4, Theorem 4], there exists a $\delta>0$ such that $|\Phi(\pm\tau,x)-\Phi(\pm\tau,y)|<\rho$ for all $y\in\mathcal{B}_{\delta}(x)$.
+This shows that all orbits $\hat{\Gamma}$ passing the circle $\mathcal{B}_{\delta}(x)$ are in particular homoclinic in $a$.
+Furthermore, by [2, §16.7, Lemma 15], we can reduce $\delta$ such that there exists a region $H\subset\mathcal{U}_{\varepsilon}(\Gamma)$, whose boundary is $\partial H=\Gamma\cup\hat{\Gamma}\cup\{a\}$.
+By following the argument in the proof of [2, §16.10, Lemma 18], we conclude that $\Gamma$ and $\hat{\Gamma}$ are nested if $\Gamma\not=\hat{\Gamma}$.
+Otherwise $H$ would be unbounded and could not lie within the bounded set $\mathcal{U}_{\varepsilon}(\Gamma)$.
+∎
 Theorem 3.6.
-Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$.
-Define $\mathcal{S}^{\prime}:=\mathcal{S}^{\prime}(\theta_{+},\theta_{-},\Xi)$.
+Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$ and $\mathcal{S}^{\prime}:=\mathcal{S}^{\prime}(\Xi)$.
 Then $\mathcal{S}$ is open.
@@
 Proof.
-Let $x\in\mathcal{S}$, i.e.
-$\Gamma(x)$ is $(\theta_{+},\theta_{-})$-elliptic in $a$.
-If $x\in\operatorname{Int}(\Xi\cup\{a\})$, we can choose $y\in\Xi$ to ensure $\Gamma(x)\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$.
-Moreover, $\operatorname{Int}(\Xi\cup\{a\})$ is open, i.e.
-nothing has to be shown.
-Hence, we can assume $x\in\Xi\cup\mathcal{S}^{\prime}$.
-Step 1: Application of [10, Proposition 4.3]
-By [10, Proposition 4.3 (iii)], there exists $r,\Theta_{\pm}>0$ such that for all $z_{\pm}\in\{z\in\mathbb{C}{}:|z-a|<r,|\arg(z-a)-\theta_{\pm}|<\Theta_{\pm}\}=:A_{\pm}$ the orbit through $z_{\pm}$ tends to $a$ in the definite direction $\theta_{\pm}$ for $t\to\pm\infty$.
-Moreover, there exists $\tau>0$ such that $M_{\pm}:=\Phi(\pm\tau,x)\in A_{\pm}$.
-Choose $\varepsilon_{\pm}>0$ small enough such that $\mathcal{B}_{\varepsilon_{\pm}}(M_{\pm})\subset A_{\pm}$.
-Define $\varepsilon:=\min\{\varepsilon_{+},\varepsilon_{-}\}$.
-By continuity of the flow, cf.
-[19, Chapter 2.4, Theorem 4], there exists a $\delta_{1}>0$ such that $|\Phi(t,x)-\Phi(t,\hat{x})|<\varepsilon$ for all $\hat{x}\in\mathcal{B}_{\delta_{1}}(x)$ and $t\in[-\tau,\tau]$.
-This shows that for all $\hat{x}\in\mathcal{B}_{\delta_{1}}(x)$ the orbit $\Gamma(\hat{x})$ is $(\theta_{+},\theta_{-})$-elliptic in $a$, cf.
-[10, Proposition 4.3 (iii)].
-Step 2: Construction of an elementary topological quadrangle
-We choose two paths without contact $l_{\pm}\subset\mathcal{B}_{\varepsilon}(M_{\pm})$ through $M_{\pm}$, cf.
-[2, §3], parameterized by ±$\ell_{\pm}:[q_{\pm},Q_{\pm}]\to l_{\pm}$ with $q_{\pm}<0<Q_{\pm}$ and $M_{\pm}=\ell_{\pm}(0)$.
-We denote the curve segment from $M_{-}$ to $M_{+}$ by $L_{0}$.
-By [2, §3, 6., Lemma 9 and proof of Lemma 10] and [2, §3, 5., Lemma 8], there exists $\sigma\in(0,\min\{|q_{-}|,Q_{-}\})$ as well as two continuously differentiable functions $\chi_{1}:[-\sigma,\sigma]\to(0,\infty)$ and $\chi_{2}:[-\sigma,\sigma]\to[q_{+},Q_{+}]$ with the following properties.
-  • (A)
-    $\chi_{2}$ is strictly monotonically increasing.
-  • (B)
-    For all $s\in[-\sigma,\sigma]$ the orbit through $\ell_{-}(s)\in l_{-}$ crosses $l_{+}$ at the point $\ell_{+}(\chi_{2}(s))$ in the time $\chi_{1}(s)$, i.e.
-$\Gamma(\ell_{-}(s))=\Gamma(\ell_{+}(\chi_{2}(s)))$ and
-    $\quad\quad\quad l_{+}\cap\Phi([0,\chi_{1}(s)],\ell_{-}(s))=\{\Phi(\chi_{1}(s),\ell_{-}(s))\}=\{\ell_{+}(\chi_{2}(s))\}\text{.}$
-Define the shorter paths without contact $\hat{l}_{-}:=\ell_{-}([-\sigma,\sigma])$ and $\hat{l}_{+}:=\ell_{+}([\chi_{2}(-\sigma),\chi_{2}(\sigma)])$ as well as the compact sets
-$R$ $:=\;\bigcup_{\mathclap{s\in[-\sigma,\sigma]}}\;\Phi([0,\chi_{1}(s)],\ell_{-}(s))\text{,}$
-$\hat{R}$ :=$:=\left\{(s,t)\in\mathbb{R}^{2}:-\sigma\leq s\leq\sigma,\,0\leq t\leq\chi_{1}(s)\right\}\text{.}$
-  • (C)
-    The function $\zeta:\hat{R}\to R$, $\zeta(s,t):=\Phi(t,\ell_{-}(s))$, is a homeomorphism, i.e.
-$R\subset\mathbb{C}{}$ is an elementary topological quadrangle around $L_{0}$, cf.
-[2, §3, p.
-69] and [2, §3, 5.
-Figure 37].
-  • (D)
-    For all $s\in[-\sigma,\sigma]$ the set $\zeta(s,[0,\chi_{1}(s)])$ is the part of the orbit $\Gamma(\ell_{-}(s))$ from $\ell_{-}(s)$ to $\ell_{+}(\chi_{2}(s))$.
-  • (E)
-    It holds
-    $\partial R=\hat{l}_{+}\cup\hat{l}_{-}\cup\zeta(\sigma,[0,\chi_{1}(\sigma)])\cup\zeta(-\sigma,[0,\chi_{1}(-\sigma)])\text{.}$
-Step 3: Impossibility of mutually exterior orbits (IMEO) in $R$
-There holds the following auxiliary result:
-  1. (IMEO)
-    If two orbits $\Gamma_{1},\Gamma_{2}\subset\mathbb{C}{}$, $\Gamma_{1}\not=\Gamma_{2}$, have nonempty intersection with the elementary topological quadrangle $R$, then either $\Gamma_{1}\subset\operatorname{Int}(\Gamma_{2})\cup\{a\})$ or $\Gamma_{2}\subset\operatorname{Int}(\Gamma_{1}\cup\{a\})$.
-In fact, suppose that there exist $-\sigma\leq s_{1}<s_{2}\leq\sigma$ with $L_{j}:=\ell_{-}(s_{j})\in\hat{l}_{-}$, $j\in\{1,2\}$, such that $\Gamma(L_{1})\subset\operatorname{Ext}(\Gamma(L_{2})\cup\{a\})$ and $\Gamma(L_{2})\subset\operatorname{Ext}(\Gamma(L_{1})\cup\{a\})$, i.e.
-$\Gamma(L_{1})$ and $\Gamma(L_{2})$ are mutually exterior in the sense of [2, §17, 3.].
-Then we can define the closed piecewise continuously differentiable Jordan curve
-$J:=\ell_{-}([s_{1},s_{2}])\cup\ell_{+}([\chi_{2}(s_{1}),\chi_{2}(s_{2})])\cup\bigcup\limits_{j=1}^{2}\zeta(s_{j},[0,\chi_{1}(s_{j})])\subset R\text{.}$
-Since $\hat{R}$ is simply connected and $\zeta$ a homeomorphism, we can apply [18, Corollary 52.5] to conclude that also $R$ is simply connected, i.e.
-we must have $\operatorname{Int}(J)\subset R$.
-Since $a\not\in R$, we get $a\not\in\operatorname{Int}(J)$.
-But this is impossible for two mutually exterior orbits $\Gamma(L_{1})$ and $\Gamma(L_{2})$, cf.
-Figure 1.
-Thus, the auxiliary result (IMEO) holds.
-[…]
-Figure 1.
-Geometrical Visualization of the construction in Step 3 of the proof of Theorem 3.6.
-The dashed black curve is $J$.
-If $\Gamma(L_{1})$ and $\Gamma(L_{2})$ are mutually exterior, then $a\in\operatorname{Int}(J)$ and $\Gamma(L_{2})$ cannot cross $\hat{l}_{-}\subset l_{-}$.
-Step 4: Proof of openness of $\mathcal{S}$
-Since one part of $\hat{l}_{-}$ must be outside of $\Gamma(x)\cup\{a\}$, we can assume w.l.o.g.
-$\ell_{-}((0,\sigma])\subset\operatorname{Ext}(\Gamma(x)\cup\{a\})$.
-By property (E), $x=\zeta(0,\tau)$ is an interior point of $R$, i.e.
-there exists $\delta_{2}>0$ such that $\mathcal{B}_{\delta_{2}}(x)\subset R$.
-We choose $\delta:=\min\{\delta_{1},\delta_{2}\}$, cf.
-Step 1, and proof $\mathcal{B}_{\delta}(x)\subset\mathcal{S}$.
-Let $\hat{x}\in\mathcal{B}_{\delta}(x)$.
-By Definition 3.1 (iii) and Proposition 3.4, the cases $\hat{x}\in\Gamma(x)\subset\mathcal{S}$ and $\hat{x}\in\overline{\operatorname{Int}(\Xi\cup\{a\})}\setminus\{a\}\subset\mathcal{S}$ are trivial and nothing has to be shown.
-Hence, assume $\Gamma(\hat{x})\subset\operatorname{Ext}(\Xi\cup\{a\})\setminus\Gamma(x)$.
-We can choose $\Gamma_{1}=\Gamma(\hat{x})$ and $\Gamma_{2}=\Gamma(x)$ in (IMEO) to conclude that $\Gamma(\hat{x})$ and $\Gamma(x)$ are not mutually exterior.
-If $\Xi=\Gamma(x)$, then $\Xi\subset\operatorname{Int}(\Gamma(\hat{x})\cup\{a\})$ and thus $\hat{x}\in\mathcal{S}$.
-Hence, we additionally assume $\Xi\not=\Gamma(x)$.
-Since $\Gamma(x)\subset\mathcal{S}$, $\Xi$ and $\Gamma(x)$ are also not mutually exterior.
-Suppose that $\Xi\subset\operatorname{Ext}(\Gamma(\hat{x})\cup\{a\})$.
-Then we have the following situation:
-  • $\bullet$
-    $\Gamma(x)\subset\operatorname{Int}(\Gamma(\hat{x})\cup\{a\})$ or $\Gamma(\hat{x})\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$.
-  • $\bullet$
-    $\Gamma(x)\subset\operatorname{Int}(\Xi\cup\{a\})$ or $\Xi\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$.
-  • $\bullet$
-    $\Gamma(\hat{x})$ and $\Xi$ are mutually exterior.
-There is only one way in which these three properties are fulfilled: $\Gamma(\hat{x})\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$ and $\Xi\subset\operatorname{Ext}(\Gamma(\hat{x})\cup\{a\})\cap\operatorname{Int}(\Gamma(x)\cup\{a\})$.
-Since both orbits, $\Gamma(\hat{x})$ and $\Gamma(x)$, cross $\hat{l}_{-}$, also $\Xi$ must cross $\hat{l}_{-}$, cf.
-Figure 1.
-It follows that $\Xi\cap R\not=\emptyset$.
-Hence, by choosing $\Gamma_{1}=\Gamma(\hat{x})$ and $\Gamma_{2}=\Xi$ in (IMEO), $\Gamma(\hat{x})$ and $\Xi$ cannot be mutually exterior.
-This is a contradiction.
-Thus, $\Xi\subset\operatorname{Int}(\Gamma(\hat{x})\cup\{a\})$ and $\hat{x}\in\mathcal{S}$.
+Let $x\in\mathcal{S}$.
+If $x$ lies in the open set $\operatorname{Int}(\Xi\cup\{a\})$, we are done.
+If $x\in\Xi$, we can choose $\varepsilon>0$ such that $\mathcal{B}_{\varepsilon}(x)\cap F^{-1}(\{0\})=\emptyset$ and apply Lemma 3.5.
+Hence, it remains to proof the Theorem for the case $x\in\mathcal{S}^{\prime}$, which is open.
+We have $\Xi\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$ and choose $\varepsilon>0$ sufficiently small such that $\mathcal{B}_{\varepsilon}(x)\subset\mathcal{S}^{\prime}$.
+By applying Lemma 3.5, we get $\delta>0$ such that the orbits $\Gamma(x)$ and $\Gamma(y)$, with $y\in\mathcal{B}_{\delta}(x)$, are nested.
+The case $y\in\mathcal{B}_{\delta}(x)\cap\operatorname{Ext}(\Gamma(x)\cup\{a\})$ leads to $\Xi\subset\operatorname{Int}(\Gamma(x)\cup\{a\})\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$.
+Let $y\in\mathcal{B}_{\delta}(x)\cap\operatorname{Int}(\Gamma(x)\cup\{a\})$.
+We have to show $\Xi\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$.
+If $\Xi\subset\operatorname{Ext}(\Gamma(y)\cup\{a\}$, then $\Xi$ and $\Gamma(y)$ are mutually exterior and $\Xi\subset\mathcal{U}_{\varepsilon}(\Gamma(x))$.
+This implies in particular $\text{dist}(\Xi,x)<\varepsilon$, which is a contradiction to our choice of $\varepsilon$.
 ∎
+The openness of the global elliptic sector now allows us to establish its unboundedness.
+At this stage, the holomorphy of $F$ plays a crucial role.
+For general vector fields, a global elliptic sector may in fact be bounded – particularly when its boundary contains a saddle whose two separatrices both connect to the multiple equilibrium $a$, cf.
+[13, Chapter 2].
 Theorem 3.7.
-Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$.
-All orbits in $\partial\mathcal{S}\setminus\{a\}$ are unbounded.
+Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$.
+All orbits on $\partial\mathcal{S}\setminus\{a\}$ are unbounded.
 $\mathcal{S}$ is unbounded.
@@
 The set $K:=\overline{\Gamma(x)}$ is compact and $\Gamma(x)\cup w_{\pm}(\Gamma(x))\subset K\subset\partial\mathcal{S}$.
-By [10, Theorem 5.1], $\Gamma(x)$ either is a periodic orbit with exactly one equilibrium, a center $\tilde{a}$, in $\operatorname{Int}(\Gamma(x))$, or $w_{+}(\Gamma(x))$ and $w_{-}(\Gamma(x))$ each consist of exactly one equilibrium.
+By using the holomorphy of $F$, we can apply Theorem 1.1 to conclude that $\Gamma(x)$ either is a periodic orbit with exactly one equilibrium, a center $\tilde{a}$, in $\operatorname{Int}(\Gamma(x))$, or $w_{+}(\Gamma(x))$ and $w_{-}(\Gamma(x))$ each consist of exactly one equilibrium.
 If the first case occurs, by openness of the basin $\mathcal{V}$ in $\tilde{a}$, cf.
 Theorem 2.4, there would be a point $y\in\mathcal{S}\cap\mathcal{V}$, $y\not=x$.
-Thus, $\Gamma(y)$ would be periodic with $a$ in its limit sets.
-This is impossible and the second case in [10, Theorem 5.1] must occur.
-By Proposition 3.5, $\Gamma(x)$ must be a homoclinic orbit with $w_{\pm}(\Gamma(x))=\{a\}$.
-In particular, by [10, Proposition 4.3 (ii)], $\Gamma(x)$ tends to $a$ in definite directions $\eta_{\pm}\in\mathcal{E}(F,m)$ for $t\to\pm\infty$.
-Moreover, there exist $r,\Theta_{\pm}>0$ such that for all $z_{\pm}\in\{z\in\mathbb{C}{}:|z-a|<r,|\arg(z-a)-\eta_{\pm}|<\Theta_{\pm}\}$ the orbit through $z_{\pm}$ tends to $a$ in the definite direction $\eta_{\pm}$ for $t\to\pm\infty$.
-Additionally, there exists $\tau>0$ such that $M_{\pm}:=\Phi(\pm\tau,x)\in A_{\pm}$.
-Choose $\varepsilon_{\pm}>0$ small enough such that $\mathcal{B}_{\varepsilon_{\pm}}(M_{\pm})\subset A_{\pm}$.
-Since $M_{\pm}\in\partial\mathcal{S}$, there exists $y_{\pm}\in\mathcal{B}_{\varepsilon_{\pm}}(M_{\pm})\cap\mathcal{S}$.
-As an orbit cannot tend to $a$ in more than one direction for $t\to\infty$ and $t\to-\infty$, it follows $\theta_{+}=\eta_{+}$ and $\theta_{-}=\eta_{-}$, i.e.
-$\Gamma(x)$ must be $(\theta_{+},\theta_{-})$-elliptic in $a$.
-Define $\varepsilon:=\min\{\varepsilon_{+},\varepsilon_{-}\}$.
-Let $L_{0}$ be the curve segment of $\Gamma(x)$ from $M_{-}$ to $M_{+}$.
-As in Step 2 of the proof of Theorem 3.6, we choose paths without contact $l_{\pm}\subset\mathcal{B}_{\varepsilon}(M_{\pm})$ through $M_{\pm}$ and construct an elementary topological quadrangle $R$ around $L_{0}$ defined via the functions $\chi_{1}$, $\chi_{2}$ and $\zeta$.
-Define $\mathcal{S}^{\prime}:=\mathcal{S}^{\prime}(\theta_{+},\theta_{-},\Xi)$.
-Since $x\in\partial\mathcal{S}$ is an interior point of $R$ and $\operatorname{Ext}(\Xi\cup\{a\})$, there exists $\hat{y}\in\mathcal{S}^{\prime}\cap R$, $\Gamma(\hat{y})\not=\Gamma(x)$.
-If $\hat{y}\in\operatorname{Int}(\Gamma(x)\cup\{a\})$, we would have $\Xi\subset\operatorname{Int}(\Gamma(\hat{y})\cup\{a\})\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$ and thus $x\in\partial\mathcal{S}\cap\mathcal{S}$, which is a contradiction to Theorem 3.6.
-Hence we must have $\hat{y}\in\operatorname{Ext}(\Gamma(x)\cup\{a\})$.
-Choose $\Gamma_{1}=\Gamma(x)$ and $\Gamma_{2}=\Gamma(\hat{y})\subset\operatorname{Ext}(\Gamma_{1}\cup\{a\})$ in (IMEO) to conclude $\Gamma(x)\subset\operatorname{Int}(\Gamma(\hat{y})\cup\{a\})$.
-We get $\Xi\subset\operatorname{Ext}(\Gamma(x)\cup\{a\})\cap\operatorname{Int}(\Gamma(\hat{y})\cup\{a\})$.
-Additionally, both orbits, $\Gamma(x)$ and $\Gamma(\hat{y})$ cross $l_{-}\subset\partial R$.
-As in Step 4 of the proof of Theorem 3.6, it follows $\Xi\cap R\not=\emptyset$.
-By choosing $\Gamma_{1}=\Gamma(x)$ and $\Gamma_{2}=\Xi$ in (IMEO), we conclude again $\Xi\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$ and thus $x\in\partial\mathcal{S}\cap\mathcal{S}$.
-This is again a contradiction to Theorem 3.6.
+This is impossible and the second case in Theorem 1.1 must occur.
+Hence, by Proposition 3.4, $\Gamma(x)$ must be a homoclinic orbit with $w_{\pm}(\Gamma(x))=\{a\}$.
+Lemma 3.5 gives us a homoclinic orbit $\hat{\Gamma}$ in $\mathcal{S}$ such that $\Gamma(x)$ and $\hat{\Gamma}$ are nested.
+However, by openness of $\mathcal{S}$, this is impossible.
 Thus, $\Gamma(x)$ cannot be homoclinic and must be unbounded.
@@
 Proposition 3.8.
-Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$.
-It holds
+Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$. It holds
 (3.1) $\mathcal{S}=\bigcup_{\mathclap{x\in\mathcal{S}}}\overline{\operatorname{Int}(\Gamma(x)\cup\{a\})}\setminus\{a\}=\bigcup_{\mathclap{x\in\mathcal{S}}}\operatorname{Int}(\Gamma(x)\cup\{a\})\text{.}$
@@
 Proof.
-Let $x\in\mathcal{S}$.
-Suppose there exists $z_{1}\in\operatorname{Int}(\Gamma(x)\cup\{a\})$ such that $z_{1}\not\in\mathcal{S}$.
-Then there exists $z_{2}\in\partial\mathcal{S}\cap\operatorname{Int}(\Gamma(x)\cup\{a\})\not=\emptyset$.
-Since the case $a=z_{2}$ is not possible, $\Gamma(z_{2})$ has to be unbounded, cf.
+This proof requires several case distinctions and is quite technical. A proof can be found in the appendix. ∎
 Theorem 3.9.
-But $\Gamma(z_{2})$ is path-connected and thus must lie completely in $\operatorname{Int}(\Gamma(x)\cup\{a\})$, which is bounded.
-This is a contradiction and $z_{1}$ cannot exist.
-Hence we have $\operatorname{Int}(\Gamma(x)\cup\{a\})\subset\mathcal{S}$.
-Additionally, by applying Proposition 3.4, we get
-$\overline{\operatorname{Int}(\Gamma(x)\cup\{a\})}\setminus\{a\}=\operatorname{Int}(\Gamma(x)\cup\{a\})\cup\Gamma(x)\subset\mathcal{S}\quad\forall\,x\in\mathcal{S}\text{.}$
-By Theorem 3.6, for all $x\in\mathcal{S}$ we can choose $y_{x}\in\operatorname{Ext}(\Gamma(x)\cup\{a\})\cap\mathcal{S}$ such that $\Gamma(x)\subset\operatorname{Int}(\Gamma(y_{x})\cup\{a\})$.
-It follows
-$\overline{\operatorname{Int}(\Gamma(x)\cup\{a\})}\setminus\{a\}\subset\operatorname{Int}(\Gamma(y_{x})\cup\{a\})\subset\mathcal{S}\quad\forall\,x\in\mathcal{S}\text{.}$
-These considerations show both equations in (3.1).
-Suppose there exist $y_{1},y_{2}\in\mathcal{S}$, $\Gamma(y_{1})\not=\Gamma(y_{2})$, such that $\Gamma(y_{1})$ and $\Gamma(y_{2})$ are mutually exterior, i.e.
-$\Gamma(y_{1})\subset\operatorname{Ext}(\Gamma(y_{2})\cup\{a\})$ and $\Gamma(y_{2})\subset\operatorname{Ext}(\Gamma(y_{1})\cup\{a\})$.
-At first, assume that $\{y_{1},y_{2}\}\cap\operatorname{Ext}(\Xi\cup\{a\})\not=\emptyset$.
-To fix ideas, assume w.l.o.g.
-$y_{1}\in\operatorname{Ext}(\Xi\cup\{a\})$.
-We have $\Xi\subset\operatorname{Int}(\Gamma(y_{1})\cup\{a\})$.
-Since $\Gamma(y_{2})\subset\operatorname{Ext}(\Gamma(y_{1})\cup\{a\})$, we get $y_{2}\in\operatorname{Ext}(\Xi\cup\{a\})$ and thus $\Xi\subset\operatorname{Int}(\Gamma(y_{1})\cup\{a\})\subset\operatorname{Ext}(\Gamma(y_{2})\cup\{a\})$.
-It follows the contradiction $y_{2}\not\in\mathcal{S}$.
-Secondly, we assume $\{y_{1},y_{2}\}\subset\operatorname{Int}(\Xi\cup\{a\})$.
-By using (3.1), we have $\operatorname{Int}(\Gamma(y_{1})\cup\{a\})\subset\mathcal{S}$, i.e.
-all orbits in $\operatorname{Int}(\Gamma(y_{1})\cup\{a\})$ are $(\theta_{+},\theta_{-})$-elliptic in $a$ and $\hat{\Xi}:=\Gamma(y_{1})$ is another sector-forming orbit in $a$ with definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$.
-Define $\hat{\mathcal{S}}:=\mathcal{S}(\theta_{+},\theta_{-},\hat{\Xi})$.
-Since $\Xi$ is $(\theta_{+},\theta_{-})$-elliptic and $\hat{\Xi}\subset\operatorname{Int}(\Xi\cup\{a\})$, we get $\Xi\subset\hat{\mathcal{S}}$.
-By using (3.1), we conclude $y_{2}\in\operatorname{Int}(\Xi\cup\{a\})\subset\hat{\mathcal{S}}$.
-Additionally, since $y_{2}\in\operatorname{Ext}(\hat{\Xi}\cup\{a\})$, we must have $\Gamma(y_{1})=\hat{\Xi}\subset\operatorname{Int}(\Gamma(y_{2})\cup\{a\})$, which is a contradiction.
-The last possible case is $\{y_{1},y_{2}\}\cap\Xi\not=\emptyset$.
-To fix ideas, we assume w.l.o.g.
-$\Xi=\Gamma(y_{1})$.
-By Definition 3.1 (iii), we must have either $y_{2}\in\operatorname{Int}(\Xi\cup\{a\})=\operatorname{Int}(\Gamma(y_{1})\cup\{a\})$ or $\Gamma(y_{1})=\Xi\subset\operatorname{Int}(\Gamma(y_{2})\cup\{a\})$.
-Both results lead to a contradiction.
-Finally, $y_{1}$ and $y_{2}$ cannot exist and we conclude that all orbits in $\mathcal{S}$ must be nested. ∎
-Theorem 3.9.
-Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$.
-$\mathcal{S}$ is connected, path-connected and simply connected.
+Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$. $\mathcal{S}$ is connected, path-connected and simply connected.
 Proof.
-Suppose, there exists a separation $(U,V)$ of $\mathcal{S}$, i.e.
-$\mathcal{S}\subset U\cup V$, where $U$ and $V$ are open, disjoint and have nonempty intersection with $\mathcal{S}$.
-Then $a$ can only be in one of these two sets.
-Assume w.l.o.g.
-$a\in U$.
-Choose $y\in V\cap\mathcal{S}$.
-Since $\Gamma(y)$ is connected, we get $\Gamma(y)\subset V$.
-Since $a\not\in\overline{V}$, it follows the contradiction $a\not\in w_{+}(\Gamma(y))\cup w_{-}(\Gamma(y))$.
-Hence, such a separation cannot exist and $\mathcal{S}$ is connected.
-By [23, Proposition 12.25], $\mathcal{V}$ is even path-connected.
-Let $L\subset\mathcal{S}$ be a loop.
-Define $J_{x}:=\Gamma(x)\cup\{a\}$ for $x\in\mathcal{S}$.
-Due to the geometry described in Proposition 3.8, we can use the same idea as in the proof of Theorem 2.5.
-By Theorem 3.6, for all $x\in L$ exists $y_{x}\in\operatorname{Ext}(J_{x})\cap\mathcal{S}$ such that $x\in\operatorname{Int}(J_{y_{x}})$.
-We can replace the periodic orbits $\Gamma(y_{x})$, $x\in\mathcal{V}$, by the curves $J_{y_{x}}$, $x\in\mathcal{S}$.
-Thus, we can find a suitable finite open cover of $L$ and an outermost orbit $J\subset\mathcal{S}$ satisfying $a\in J$ and $L\subset\operatorname{Int}(J)$.
-By (3.1), we have $\operatorname{Int}(J\cup\{a\})\subset\mathcal{S}$, which is clearly simply connected.
-Hence $L$ can be continuously deformed to a point in $L\subset\mathcal{S}$.
-Since $\mathcal{S}$ is path-connected, it follows that $\mathcal{S}$ is even simply connected.
-∎
-We now address the independence of the global elliptic sector from a particular choice of the sector-forming orbit.
+Analogous to the proof of Theorem 2.5, this follows directly from (3.1) together with the Seifert-van Kampen theorem for infinite open covers, cf. [15, Theorem 1.20]. ∎
+We now address the independence of the global elliptic sector from a particular choice of the sector-forming orbit ensuring in some sense the well-definedness of the global elliptic sector.
 Lemma 3.10.
-Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$.
-For all $x\in\mathcal{S}$ the orbit $\Gamma(x)$ is a sector-forming orbit in $a$ with definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ and $\mathcal{S}=\mathcal{S}(\theta_{+},\theta_{-},\Gamma(x))$, i.e.
+Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$.
+For all $x\in\mathcal{S}$ the orbit $\Gamma(x)$ is a sector-forming orbit in $a$ and $\mathcal{S}=\mathcal{S}(\Gamma(x))$, i.e.
 the global elliptic sector $\mathcal{S}$ does not depend on the particular choice of a sector-forming orbit.
 Proof.
-Let $x\in\mathcal{S}$ be arbitrary.
-By (3.1), all orbits in $\operatorname{Int}(\Gamma(x)\cup\{a\})$ are $(\theta_{+},\theta_{-})$-elliptic and $\hat{\Xi}:=\Gamma(x)$ is indeed an sector-forming orbit in $a$ with definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$.
-Define $\hat{\mathcal{S}}:=\mathcal{S}(\theta_{+},\theta_{-},\hat{\Xi})$.
-If $\hat{\Xi}=\Xi$, then obviously $\mathcal{S}=\hat{\mathcal{S}}$.
-We have to show $\mathcal{S}\subset\hat{\mathcal{S}}$ as well as $\hat{\mathcal{S}}\subset\mathcal{S}$ for the two cases $\hat{\Xi}\subset\operatorname{Int}(\Xi\cup\{a\})$ and $\Xi\subset\operatorname{Int}(\hat{\Xi}\cup\{a\})$.
-As these four proofs are all based on similar arguments, we assume w.l.o.g.
-$\hat{\Xi}\subset\operatorname{Int}(\Xi\cup\{a\})$ and show only $\mathcal{S}\subset\hat{\mathcal{S}}$.
-Let $y\in\mathcal{S}$.
-If $y\in\Xi\cup\hat{\Xi}\cup\operatorname{Int}(\hat{\Xi}\cup\{a\})$, then nothing is to show.
-Thus, assume $y\in\operatorname{Ext}(\hat{\Xi}\cup\{a\})\setminus\Xi$.
-If $y\in\operatorname{Ext}(\Xi\cup\{a\})$, then we have $\Xi\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$ and thus $\hat{\Xi}\subset\operatorname{Int}(\Xi\cup\{a\})\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$, i.e.
-$y\in\hat{\mathcal{S}}$.
-Additionally, if $y\in\operatorname{Ext}(\hat{\Xi}\cup\{a\})\cap\operatorname{Int}(\Xi\cup\{a\})$, then we can apply Proposition 3.8 with $y_{1}:=x$ and $y_{2}:=y$ to conclude $\hat{\Xi}\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$, i.e.
-$y\in\hat{\mathcal{S}}$.
-∎
-Finally, at the end of this chapter, we want to supplement and enhance our topological study with two important corollaries, which confirm our intuitive idea of the geometrical structure of an equilibrium of order $m\geq 2$, cf.
-[10, Figure 2, Figure 3].
+Let $x\in\mathcal{S}$. By (3.1), all orbits in $\operatorname{Int}(\Gamma(x)\cup\{a\})$ are homoclinic and $\hat{\Xi}:=\Gamma(x)$ is indeed a sector-forming orbit in $a$.
+[…]
+3.3. Final Corollaries
+Finally, at the end of this chapter, we supplement and enhance our topological study with several important corollaries, which confirm the intuitive idea of the geometrical structure of an equilibrium of order $m\geq 2$, cf.
+[19, Figure 2, Figure 3].
 Corollary 3.11.
-There exist exactly $22m-$ 2 distinct444in the sense of [2, §17, 3.].
-global elliptic sectors in $a$.
-For all $x\in\mathbb{C}{}$ with $w_{+}(\Gamma(x))=w_{-}(\Gamma(x))=\{a\}$ there exist two adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ such that $\Gamma(x)$ is a sector-forming orbit for $\mathcal{S}(\theta_{+},\theta_{-},\Gamma(x))$.
+There exist exactly 2m22m-2 distinct [^4] This means that $\Gamma_{1}$ and $\Gamma_{2}$ are mutually exterior, if if they belong to two different global elliptic sectors.
+global elliptic sectors in $a$, each located between two adjacent definite directions given by $\mathcal{E}(F,m)$.
+All homoclinic orbits are sector-forming orbits.
 Proof.
-By [10, Theorem 4.4], there is a FED of order $d:=2m-2$ in $a$.
-We use the geometrical objects $\Gamma$, $\Gamma_{i}$ and $p_{i}$, $i\in\{1,\ldots,d\}$, defined by [10, Definition 4.1 d)].
+By [19, Theorem 4.4], there is a FED of order d:=2m2d:=2m-2 in $a$.
+We use the geometrical objects $\Gamma$, $\Gamma_{i}$ and $p_{i}$, $i\in\{1,\ldots,d\}$, defined by [19, Definition 4.1 d)], cf Figure 1.
 We order the definite directions $\mathcal{E}(F,m)=\{\theta_{1},\ldots,\theta_{d}\}$ in such a way that for all $i\in\{1,\ldots,d\}$ the local elliptic sector with characteristic orbits $\Gamma_{i}$ and $\Gamma_{i+1}$ lies between $\theta_{i}$ and $\theta_{i+1}$, i.e.
 $\Gamma_{i}$ tends to $a$ in the definite direction $\theta_{i}$.
-[^5] If $i=d$, we define $\Gamma_{i+1}:=\Gamma_{1}$ and $\theta_{i+1}:=\theta_{1}$.
-By using the notation in [10, Definition 4.1 a)] and applying [10, Definition 4.1 c) (iv)], there exist points $E_{i}\in\Gamma(p_{i},p_{i+1})$ such that $\Xi_{i}:=\Gamma(E_{i})$ is a sector-forming orbit with definite directions $\theta_{i},\theta_{i+1}$ for all $i\in\{1,\ldots,d\}$.
+By using [19, Definition 4.1 c) (iv)], there exist points $E_{i}\in\Gamma(p_{i},p_{i+1})$ such that $\Xi_{i}:=\Gamma(E_{i})$ is a sector-forming orbit with definite directions $\theta_{i},\theta_{i+1}$ for all $i\in\{1,\ldots,d\}$.
 Moreover, since all local sectors in $a$ have pairwise empty intersection except for the characteristic orbits and $a$, these sector-forming orbits are pairwise mutually exterior.
-Define the set
-$A:=\overline{\operatorname{Int}(\Gamma)}\setminus\bigcup_{i=1}^{d}\overline{\operatorname{Int}(\Xi_{i}\cup\{a\})}$
-as well as the $d$ global elliptic sectors
-$\mathcal{S}_{i}:=\begin{dcases}\mathcal{S}(\theta_{i},\theta_{i+1},\Xi_{i})&\text{if }\lambda(\theta_{i})=-1\\ \mathcal{S}(\theta_{i+1},\theta_{i},\Xi_{i})&\text{if }\lambda(\theta_{i})=1\end{dcases}\;,\,i\in\{1,\ldots,d\}\text{.}$
-Suppose there exist $i,j\in\{1,\ldots,d\}$, $i\not=j$, with $z\in\mathcal{S}_{i}\cap\mathcal{S}_{j}\not=\emptyset$.
-Since $\Xi_{i}$ and $\Xi_{j}$ are mutually exterior, we must have $z\in\operatorname{Ext}(\Xi_{i}\cup\{a\})\cap\operatorname{Ext}(\Xi_{j}\cup\{a\})$ and thus $\Xi_{i}\cup\Xi_{j}\subset\operatorname{Int}(\Gamma(z)\cup\{a\})$.
-By (3.1), we have $\operatorname{Int}(\Gamma(z)\cup\{a\})\subset\mathcal{S}_{i}$, i.e.
-$\Xi_{i}\cup\Xi_{j}\subset\mathcal{S}_{i}$.
-Since $\Xi_{i}$ and $\Xi_{j}$ are mutually exterior, we can derive a contradiction to the second statement in Proposition 3.8.
-Thus, these global elliptic sectors are distinct.
-Let $x\in\mathbb{C}{}\setminus\{a\}$ be arbitrary with $w_{+}(\Gamma(x))=w_{-}(\Gamma(x))=\{a\}$.
-By [10, Proposition 4.3 (ii)], $\Gamma(x)$ has to tend to $a$ in definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ for $t\to\infty$ and $t\to-\infty$, respectively.
-At first, we assume $\Gamma(x)\subset\overline{\operatorname{Int}(\Gamma)}$.
-Then, by applying the geometrical structure of the FED and [10, Proposition 4.3 (iii)], all orbits through $A$ leave $\overline{\operatorname{Int}(\Gamma)}$ in finite time.
-Hence, there exists $i\in\{1,\ldots,d\}$ such that $\Gamma(x)\subset\Xi_{i}\cup\operatorname{Int}(\Xi_{i}\cup\{a\})\subset\mathcal{S}_{i}$.
-By Definition 3.1 (iii), it follows that $\{\theta_{+},\theta_{-}\}=\{\theta_{i},\theta_{i+1}\}$, i.e.
-$\theta_{+},\theta_{-}$ are adjacent.
-Next, we consider the case $\Gamma(x)\cap\operatorname{Ext}(\Gamma)\not=\emptyset$.
-As $\Gamma(x)$ must leave and enter $\overline{\operatorname{Int}(\Gamma)}$ again, a part of the closed Jordan curve $\Gamma$ must lie in $\operatorname{Int}(\Gamma(x)\cup\{a\})$.
-By [10, Definition 4.1 c) (iii)], there must exist $i\in\{1,\ldots,d\}$ with $E_{i}\in\operatorname{Int}(\Gamma(x)\cup\{a\})$, i.e.
-$\Xi_{i}\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$.
-Suppose a second sector-forming orbit is in the interior of $\Gamma(x)\cup\{a\}$, i.e.
-there exists $j\not=i$ with $\Xi_{j}\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$.
-If $x\in\mathcal{S}_{i}$, then, by applying the second statement in Proposition 3.8, $\Xi_{i}$ and $\Xi_{j}$ cannot be mutually exterior, which is a contradiction.
-Thus, $x\not\in\mathcal{S}_{i}$.
-Since $\mathcal{S}_{i}$ is connected, cf.
-Theorem 3.9, we conclude $\mathcal{S}_{i}\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$.
-This makes $\mathcal{S}_{i}$ bounded, which is a contradiction to Theorem 3.7.
-Hence, $\Xi_{i}$ is the only sector-forming orbit in $\operatorname{Int}(\Gamma(x)\cup\{a\})$.
-By applying the geometrical structure of the FED, we again get $\{\theta_{+},\theta_{-}\}=\{\theta_{i},\theta_{i+1}\}$, i.e.
-$\theta_{+},\theta_{-}$ are adjacent.
-In particular, by Definition 3.1 (iii), also in this case it follows $x\in\mathcal{S}_{i}$.
-Finally, in both cases we can apply Lemma 3.10 to conclude that $\Gamma(x)$ is a sector forming orbit for the global elliptic sector $\mathcal{S}(\theta_{+},\theta_{-},\Gamma(x))$ and $\mathcal{S}(\theta_{+},\theta_{-},\Gamma(x))=\mathcal{S}_{i}$.
-This shows the second statement.
-Moreover, using this result, there cannot exist more then the $d$ distinct global elliptic sectors $\mathcal{S}_{1},\ldots,\mathcal{S}_{d}$.
-This completes the proof of the first statement.
+This leads to $d$ global elliptic sectors, generated by the $d$ pairwise mutually exterior sector-forming orbits $\Xi_{1},\ldots,\Xi_{d}$.
+On the one hand, due to the local geometry of $a$ it is clear that each global elliptic sector admits at least one of the orbits $\Xi_{1},\ldots,\Xi_{d}$.
+On the other hand, since all orbits in a global elliptic sector are nested, it is not possible to find two orbits $\Xi_{i},\Xi_{j}$, $i\not=j$, lying in the same global elliptic sector.
+Hence, we get indeed exactly $d$ distinct global elliptic sectors in $a$.
+The rest follows from the geometry described in [19, Proposition 4.3].
 ∎
-We have proved Corollary 3.11 by applying the geometry of the local structure in $a$.
+We have proved Corollary 3.11 by using the geometry of the local structure in $a$.
 This structure is defined by a FED, which is not uniquely determined, i.e.
 there exist several FEDs in $a$.
-Nevertheless, we can now see that the global structure is indeed unique.
+However, we can now see that the global structure is the unique globalization of this local structure.
+In the following, we consider the case $m\geq 3$.
+In this setting, the global elliptic sector can be defined independently of sector-forming orbits, solely by means of the chosen adjacent definite directions.
+For this purpose, we require the function $\lambda:\mathcal{E}(F,m)\to\{-1,1\}$,
+$\lambda(\theta):=\cos(\operatorname{arg}(F^{(m)}(a))+\theta(m-1))$
+to determine the time direction (positive or negative) in which the orbits tend to $a$ in the FED, cf.
+[19, Proposition 4.3 (iii)] as well as the calculation of $\lambda_{i}$ on page 11 in this reference.
+Definition 3.12.
+Let $m\geq 3$.
+  • (i)
+    Let $\Gamma\subset\mathbb{C}$ be an orbit and $x\in\Gamma$.
+We call $\Gamma$ $(\theta_{+},\theta_{-})$-elliptic in $a$, if there exist $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ such that $w_{+}(\Gamma)=w_{-}(\Gamma)=\{a\}$, $\arg(\Phi(t,x)-a)=\theta_{+}$ for $t\to\infty$ and $\arg(\Phi(t,x)-a)=\theta_{-}$ for $t\to$-$\infty$, that is, $\Gamma$ is homoclinic and tends to $a$ in the definite direction $\theta_{\pm}$ for $t\to\pm\infty$.
+  • (ii)
+    Let $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ be two adjacent definite directions satisfying $\lambda(\theta_{\pm})=\mp 1$.
+[^5] If the corresponding local elliptic sector between $\theta_{+}$ and $\theta_{-}$ has clockwise (counterclockwise) direction, then $\theta_{-(+)}=\theta_{+(-)}+\frac{\pi}{m-1}$.
+We define the set
+    (3.2) $\mathcal{S}(\theta_{+},\theta_{-}):=\left\{x\in\mathbb{C}{}:\Gamma(x)\text{ is }(\theta_{+},\theta_{-})\text{-elliptic in }a\right\}\text{.}$
 Corollary 3.13.
 Let $m\geq 3$.
-Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$.
-Then it holds
-(3.2) $\mathcal{S}(\theta_{+},\theta_{-},\Xi)=\left\{x\in\mathbb{C}{}:\Gamma(x)\text{ is }(\theta_{+},\theta_{-})\text{-elliptic in }a\right\}\text{.}$
+  • (i)
+    Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$. Then there exist $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ such that
+    (3.3) $\mathcal{S}(\Xi)=\mathcal{S}(\theta_{+},\theta_{-})\text{.}$
+  • (ii)
+    Let $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ be two adjacent definite directions satisfying $\lambda(\theta_{\pm})=\mp 1$. Then there exists a sector forming orbit $\Xi$ such that
+    (3.4) $\mathcal{S}(\theta_{+},\theta_{-})=\mathcal{S}(\Xi)\text{.}$
 Proof.
-By Definition 3.1, $\subset$" is clear.
-Since each of the $22m-$ 2 distinct global elliptic sectors lies between two definite directions and $m\geq 3$, there exists exactly one global elliptic sector $\mathcal{S}_{0}$ between $\theta_{+}$ and $\theta_{-}$, cf.
+The key point underlying this result is that, for $m\geq 3$, there exist more than two definite directions.
+Since each of the 2m22m-2 distinct unbounded global elliptic sectors lies between two adjacent definite directions, there always exists exactly one global elliptic sector $\mathcal{S}_{0}$ between two adjacent arbitrarily chosen angles $\theta_{+}$ and $\theta_{-}$, cf.
 Corollary 3.11.
-If there was a second one, there would be only $22<2m-$ 2 global elliptic sectors in $a$.
-Thus, we must have $\mathcal{S}(\theta_{+},\theta_{-},\Xi)=\mathcal{S}_{0}$.
-Moreover, by applying the second statement in Corollary 3.11, all $(\theta_{+},\theta_{-})$-elliptic orbits must lie in one of these $22m-$ 2 global elliptic sectors.
-Since the definite directions on the left hand side coincide with those on the right hand side in (3.2), all these orbits have to lie in $\mathcal{S}_{0}$ and thus in $\mathcal{S}(\theta_{+},\theta_{-},\Xi)$.
-This shows $\supset$".
+Moreover, by applying the second statement in Corollary 3.11, all $(\theta_{+},\theta_{-})$-elliptic orbits must lie in one of these 2m22m-2 global elliptic sectors.
+This shows (3.3) and (3.4).
 ∎
 This second corollary shows that we need to choose a sector-forming orbit only in the case $m=2$.
-Here we have only two definite directions, which requires the choice of one arbitrary elliptic orbit to specify one of the two global elliptic sectors.
+Here we have only two definite directions, which requires the choice of one arbitrary homoclinic orbit to specify one of the two global elliptic sectors.
 The two sectors cannot be distinguished on the basis of the definite directions alone.
@@
 The set would become connected (but no longer open) only by adding the equilibrium $a$.
-For this reason, the construction in Definition 3.1 uses a sector-forming orbit.
-We observe that the boundaries of two adjacent global elliptic sectors do not have to coincide near $a$, as it is the case for the characteristic orbits in a FED, cf.
-[10, Theorem 4.4].
+For this reason, the construction in Definition 3.1 is based on the notion of a sector-forming orbit, which is defined independently of the choice of definite directions.
+We observe that the boundaries of two adjacent global elliptic sectors in general do not have to coincide near $a$, as it is the case for the characteristic orbits in a FED.
 However, if they indeed coincide between all adjacent sectors, the local structure of the FED can even be transferred to the global phase portrait.
 The question therefore arises as to what the global geometric structure looks like in a region between two adjacent global elliptic sectors having no common boundary near $a$.
-How many unbounded orbits can exist in sum?
-This question takes up the discussions at the end of [10, Example 4.7] and [10, Example 4.8].
+How many separatrices can tend to a multiple equilibrium (either in positive or negative time)?
+This question takes up the discussions at the end of [19, Example 4.7] and [19, Example 4.8].
+We will discuss these issues in Chapter 5.
+Finally, we conclude by addressing the question of how the global elliptic sector fits into the theory of canonical regions.
+Corollary 3.14.
+Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$.
+$\mathcal{S}$ is a canonical region of strip type.
+There exists a global smooth transversal through $\mathcal{S}$, i.e.
+an open smooth arc crossing each orbit in $\mathcal{S}$ exactly once.
+Proof.
+By openness of $\mathcal{S}$, all orbits in $\mathcal{S}$ are part of a strip canonical region $R$.
+Thus, there are no separatrices in $\mathcal{S}$.
+Since $\partial\mathcal{S}$ does not have any homoclinic orbits, we have $\partial\mathcal{S}\cap R=\emptyset$, i.e.
+$R=\mathcal{S}$.
+The rest follows directly from [9, Theorem 5.1] and [10, Theorem I].
+∎
 4. Basin of nodes and foci
-This chapter provides detailed proofs of the topological results given in [4, Theorem 4.3].
-In all claims and arguments of this chapter we consider an entire function $F\in\mathcal{O}{}(\mathbb{C})$, $F\not\equiv 0$, and a node or focus $a\in\mathbb{C}$ of (1.1).
+In this chapter, we consider an entire function $F\in\mathcal{O}{}(\mathbb{C})$, $F\not\equiv 0$, and a node or focus $a\in\mathbb{C}$ of (1.1).
+We provide detailed proofs of the topological results given in [5, Theorem 4.3].
 Definition 4.1.
-The stable (unstable) basin $\mathcal{N}$ of $F$ in $a$ is defined as
+The stable (unstable) basin of attraction (repulsion) $\mathcal{N}$ of $F$ in $a$ is defined as
 $\mathcal{N}:=\left\{x\in\mathbb{C}:w_{+(-)}(\Gamma(x))=\{a\}\right\}.$
-The boundary of $a$ is defined as the boundary of the basin of $F$ in $a$.
+The boundary of $a$ is defined as the boundary of $\mathcal{N}$.
 Proposition 4.2.
@@
 that $a$ is stable.
-By [10, Definition 3.1], there exists $\varepsilon_{1}>0$ such that for all $y\in\mathcal{B}_{\varepsilon_{1}}(a)$ we have $w_{+}(\Gamma(y))=\{a\}$.
+By [19, Definition 3.1], there exists $\varepsilon_{1}>0$ such that for all $y\in\mathcal{B}_{\varepsilon_{1}}(a)$ we have $w_{+}(\Gamma(y))=\{a\}$.
 This proves $\mathcal{B}_{\varepsilon_{1}}(a)\subset\mathcal{N}$ and $a\in\mathring{\mathcal{N}}$.
@@
 By continuity of the flow, cf.
-[19, Chapter 2.4, Theorem 4], there exists a $\delta>0$ such that $|\Phi(\tau,z)-\xi|<\varepsilon_{2}$ for all $z\in\mathcal{B}_{\delta}(x)$.
+[29, Chapter 2.4, Theorem 4], there exists a $\delta>0$ such that $|\Phi(\tau,z)-\xi|<\varepsilon_{2}$ for all $z\in\mathcal{B}_{\delta}(x)$.
 Thus $w_{+}(\Gamma(z))=\{a\}$ for all $z\in\mathcal{B}_{\delta}(x)$ and $\mathcal{N}$ is open.
@@
 Since $a$ lies on the boundary of each orbit in $\mathcal{N}$, every set in $C$ is connected, cf.
-[18, Theorem 23.4].
-Moreover, for $J,\tilde{J}\in C$ it always holds $J\cap\tilde{J}\supset\{a\}\not=\emptyset$.
-Thus, by [18, Theorem 23.3], $\mathcal{N}$ is connected.
-From [23, Proposition 12.25] it follows that $\mathcal{N}$ is even path-connected.
+[28, Theorem 23.4].
+Moreover, for $J,\tilde{J}\in C$ it always holds $J\cap\tilde{J}\supset\{a\}\not$$\emptyset$.
+Thus, by [28, Theorem 23.3], $\mathcal{N}$ is connected.
+From [33, Proposition 12.25] it follows that $\mathcal{N}$ is even path-connected.
 ∎
-Lemma 4.4.
-All centers $\tilde{a}\in F^{-1}(\{0\})\setminus\{a\}$ with its corresponding basin $\mathcal{V}$ satisfy $\mathcal{V}\cap\partial\mathcal{N}=\emptyset$.
-In particular, there are no centers on $\partial\mathcal{N}$.
-Proof.
-Assume w.l.o.g.
-that $a$ is stable.
-Suppose there exists a center $\tilde{a}\in\mathbb{C}{}\setminus\{a\}$ such that its basin $\mathcal{V}$ satisfies $\mathcal{V}\cap\partial\mathcal{N}\not=\emptyset$.
-Since $\mathcal{V}$ is open, cf.
-Theorem 2.4, we can choose $z\in\mathcal{V}\cap\partial\mathcal{N}$ and $r>0$ sufficiently small such that $\mathcal{B}_{r}(z)\subset\mathcal{V}$.
-Since $z\in\partial\mathcal{N}$, there exists $y\in\mathcal{N}\cap\mathcal{B}_{r}(z)\not=\emptyset$ satisfying $w_{+}(\Gamma(y))=\{a\}$.
-But this is impossible for a periodic orbit in $\mathcal{V}$.
-Thus $\mathcal{V}\cap\partial\mathcal{N}=\emptyset$.
-In particular, there are no centers on $\partial\mathcal{N}$, i.e.
-all equilibria on the boundary of $a$ have to be nodes, foci or zeros of order $m\geq 2$.
-∎
-In the proof of [4, Theorem 4.3] the author does not investigate the case of an isolated equilibrium on the boundary of a node or focus.
+In the proof of [5, Theorem 4.3] the author does not investigate the case of an isolated equilibrium on the boundary of a node or focus.
 This is a significant gap, which is closed by the following result.
@@
 It holds
-$\forall\,\tilde{a}\in\partial\mathcal{N}\cap F^{-1}(\{0\}):\forall\,\rho>0:\left(\mathcal{B}_{\rho}(\tilde{a})\cap\partial\mathcal{N}\right)\setminus\{\tilde{a}\}\not=\emptyset,$
+$\forall\,\tilde{a}\in\partial\mathcal{N}\cap F^{-1}(\{0\}):\forall\,\rho>0:\left(\mathcal{B}_{\rho}(\tilde{a})\cap\partial\mathcal{N}\right)\setminus\{\tilde{a}\}\not$$\emptyset$,
 i.e.
@@
 Proof.
-Suppose there exists an isolated point $\tilde{a}\in\partial\mathcal{N}\cap F^{-1}(\{0\})$, i.e. there exists $\rho>0$ such that
-$\partial\mathcal{N}\cap(\underbrace{\mathcal{B}_{\rho}(\tilde{a})\setminus\{\tilde{a}\}}_{=:A})=\left(\mathcal{B}_{\rho}(\tilde{a})\cap\partial\mathcal{N}\right)\setminus\{\tilde{a}\}=\emptyset\text{.}$
-It follows [^6] The simple but technical proof of ^$\hat{\partial}(\mathcal{N}\cap A)\subset\partial\mathcal{N}\cap A=\emptyset$, where $\hat{\partial}$ denotes the boundary with respect to the subspace topology on $A$, is omitted.
-that $\mathcal{N}\cap A\not=\emptyset$ is open and closed with respect to the subspace topology on $A$.
-Since the punctured circle $A$ is connected, we must have $\mathcal{N}\cap A=A$.
-Hence, since $\tilde{a}\not\in\mathcal{N}$, we conclude
-(4.1) $\mathcal{N}\cap\mathcal{B}_{\rho}(\tilde{a})=\mathcal{B}_{\rho}(\tilde{a})\setminus\{\tilde{a}\}\text{.}$
-If $\tilde{a}$ was of order $m\geq 2$, then, by [10, Theorem 4.4], there would be at least one homoclinic orbit in $A$ tending to $\tilde{a}$ in both time directions, i.e.
-equation (4.1) would not be fulfilled and $\tilde{a}$ would not be isolated.
-Thus, $\tilde{a}$ must be a node or focus, cf.
-Lemma 4.4.
-Denote by $\tilde{\mathcal{N}}$ the basin of $F$ in $\tilde{a}$.
-We reduce $\rho$ such that $\mathcal{B}_{\rho}(\tilde{a})\subset\tilde{\mathcal{N}}$ and $\mathcal{B}_{\rho}(a)\subset\mathcal{N}$.
-By (4.1) and Proposition 4.2, we have $\tilde{\mathcal{N}}\subset\mathcal{N}$.
-By applying the theory of circles without contact around nodes and foci, cf.
-[2, §3, 10.-14., §7, 1.-2.
-and §18, Lemma 3], we find two continuously differentiable closed paths $C_{1}\subset\mathcal{B}_{\rho}(a)$ and $C_{2}\subset\mathcal{B}_{\rho}(\tilde{a})$, $C_{1}\cap C_{2}=\emptyset$, being nowhere tangential to $F$ and satisfying $\operatorname{Int}(C_{1})\cap F^{-1}(\{0\})=\{a\}$ and $\operatorname{Int}(C_{2})\cap F^{-1}(\{0\})=\{\tilde{a}\}$.
-[^7] In particular, from the equations (6) and (11) in [2, §7, 1.] and the remarks made in [2, §7, 2.] it follows that $C_{1}$ as well as $C_{2}$ can be chosen as a linear transformed circle or ellipse.
-It holds $\overline{\operatorname{Int}(C_{1})}\cap\overline{\operatorname{Int}(C_{2})}=\emptyset$.
-Moreover, every orbit in $\mathcal{N}$ ($\tilde{\mathcal{N}}$) crosses $C_{1}$ ($C_{2}$) exactly once, cf.
-[2, §3, 10., Figure 54].
-We equip $C_{1}$ and $C_{2}$ with the subspace topology.
-Using this construction, the map $\Psi:\mathcal{N}\setminus\{a\}\to\mathbb{R}$ is well-defined, when $\Psi(x)$ is the transit time888cf.
-[4, Definition 3.2].
-the point $x\in\mathcal{N}\setminus\{a\}$ needs to $C_{1}$ along the path $\Gamma(x)$, i.e.
-$\{\Phi(\Psi(x),x)\}=\Gamma(x)\cap C_{1}$.
-Moreover, we define $\eta:C_{2}\to C_{1}$ by $\eta(x):=\Phi(\Psi(x),x)$.
-By the continuous dependence of the flow on initial conditions, cf.
-[19, Chapter 2.4, Theorem 4], and the continuous dependence of the transit time integral on the integral limits, $\Psi$ and $\eta$ are both continuous.
-One the one hand, since $C_{2}$ is compact, also $\eta(C_{2})$ is compact in $C_{1}$, cf.
-[18, Theorem 26.5].
-Thus, $\eta(C_{2})$ is closed in $C_{1}$, cf.
-[18, Theorem 26.3].
-On the other hand, by Proposition 4.2, we have $\eta(C_{2})\subset\tilde{\mathcal{N}}\cap C_{1}$.
-Additionally, since all orbits through $\tilde{\mathcal{N}}\cap C_{1}\subset\tilde{\mathcal{N}}$ have to cross $C_{2}$, we also have $\tilde{\mathcal{N}}\cap C_{1}\subset\eta(C_{2})$, i.e.
-$\eta(C_{2})=\tilde{\mathcal{N}}\cap C_{1}$.
-Moreover, by Theorem 4.3, $\tilde{\mathcal{N}}\cap C_{1}$ is open in $C_{1}$.
-We conclude that $\eta(C_{2})\not=\emptyset$ is open and closed in $C_{1}$.
-Since $C_{1}$ is connected, we must have $C_{1}=\eta(C_{2})$, i.e.
-$\eta$ is surjective.
-Let $y_{1},y_{2}\in C_{2}$ be arbitrary.
-Assume $\eta(y_{1})=\eta(y_{2})$.
-Then, by definition of $\eta$, $\Gamma(y_{1})=\Gamma(\eta(y_{1}))=\Gamma(\eta(y_{2}))=\Gamma(y_{2})$.
-Since all orbits in $\tilde{\mathcal{N}}$ cross $C_{2}$ exactly once, we must have $y_{1}=y_{2}$.
-Hence, $\eta$ is also injective.
-It follows that $\eta$ is bijective and thus a homeomorphism, cf.
-[18, Theorem 26.6].
-We conclude that every orbit in $\mathcal{N}$ and $\tilde{\mathcal{N}}$ crosses $C_{1}$ as well as $C_{2}$ exactly once, i.e.
-[^9] The set $\mathcal{H}$ is the so-called heteroclinic region between $a$ and $\tilde{a}$ and is introduced in Chapter 5.
-(4.2) $\mathcal{H}:=\mathcal{N}\cap\tilde{\mathcal{N}}=\mathcal{N}\setminus\{a\}=\tilde{\mathcal{N}}\setminus\{\tilde{a}\}=\bigcup\limits_{\mathclap{x_{2}\in C_{2}}}\Gamma(x_{2})=\bigcup\limits_{\mathclap{x_{1}\in C_{1}}}\Gamma(x_{1})\text{.}$
-Since $C_{1}\subset\mathcal{H}$ is a Jordan curve, there exists a homeomorphism $h:C_{1}\to S^{1}$.
-Equip the cylinder surface $B:=\mathbb{R}\times S^{1}\subset\mathbb{R}^{3}$ with the subspace topology and define the map $\zeta:B\to\mathcal{H}$ by $\zeta(t,x):=\Phi(t,h^{-1}(x))$.
-By (4.2), we have $\zeta(B)=\mathcal{H}$, i.e.
-$\zeta$ is surjective.
-Let $(t_{1},x_{1}),(t_{2},x_{2})\in B$ be arbitrary and assume $\zeta(t_{1},x_{1})=\zeta(t_{2},x_{2})$.
-Then $h^{-1}(x_{1}),h^{-1}(x_{2})\in C_{1}$ and
-$\Gamma(h^{-1}(x_{1}))=\Gamma(\zeta(t_{1},x_{1}))=\Gamma(\zeta(t_{2},x_{2}))=\Gamma(h^{-1}(x_{2}))\text{.}$
-Since all orbits in $\mathcal{H}$ cross $C_{1}$ exactly once, we get $h^{-1}(x_{1})=h^{-1}(x_{2})$.
-Since $h$ is bijective, we conclude $x_{1}=x_{2}$.
-This gives us $\Phi(t_{1},h^{-1}(x_{1}))=\Phi(t_{2},h^{-1}(x_{1}))$ and thus $t_{1}=t_{2}$, since $\Gamma(h^{-1}(x_{1}))$ is not periodic.
-It follows that $\zeta$ is also injective and thus bijective.
-Hence, there exists the inverse $\zeta^{-1}$ of $\zeta$, which is given by $\zeta^{-1}(x):=(\Psi(x),h(\Phi(\Psi(x),x)))$, $x\in\mathcal{H}$.
-As a composition of continuous functions, $\zeta$ as well as $\zeta^{-1}$ are continuous and therefore $\zeta$ is a homeomorphism.
-By [18, Corollary 52.5], the fundamental group of $\mathcal{H}$ is isomorphic to $B$.
-Additionally, as the circle line $S^{1}\subset\mathbb{R}^{2}$ is a deformation retract of the cylinder surface $B$, we conclude that the fundamental group of $\mathcal{H}$ is isomorphic to the additive group $\mathbb{Z}$, cf.
-[18, Theorem 54.5] and [18, Theorem 58.3].
-But $\mathcal{H}$ has the two holes $a$ and $\tilde{a}$, which is impossible for a set having a fundamental group isomorphic to $\mathbb{Z}$.
-In particular, by using (4.2), we can choose the loop $C_{2}\cup\Phi([0,\Psi(z)],z)\cup C_{1}\subset\mathcal{H}$ with an arbitrary point $z\in C_{2}$.
-This loop has $a\not\in\mathcal{H}$ as well as $\tilde{a}\not\in\mathcal{H}$ in its interior and thus cannot be transferred to a loop on $S^{1}$.
-All in all, we derive a contradiction and $\tilde{a}$ cannot be isolated on $\partial\mathcal{N}$.
-Moreover, all equilibria $\hat{a}\in\partial\mathcal{N}\cap F^{-1}(\{0\})$ are nodes, foci or equilibria of order $m\geq 2$, cf.
-Lemma 4.4.
-Thus, we can choose $\hat{\rho}$ such that for all $\xi\in\mathcal{B}_{\hat{\rho}}(\hat{a})$ there holds $\hat{a}\in w_{+}(\Gamma(\xi))\cup w_{-}(\Gamma(\xi))$, cf.
-[10, Definition 3.1] and [10, Theorem 4.4].
-If we choose $\xi\in(\mathcal{B}_{\hat{\rho}}(\hat{a})\cap\partial\mathcal{N})\setminus\{\hat{a}\}\not=\emptyset$, then $\Gamma:=\Gamma(\xi)\subset\mathcal{N}$ satisfies $\hat{a}\in w_{+}(\Gamma)\cup w_{-}(\Gamma)$, i.e.
-$\hat{a}$ is attached to $\Gamma$.
+This proof, carried out by contradiction, relies on constructing a homeomorphism between two appropriately chosen circles without contact, each encircling a focus/node.
+Since this is highly technical, a very detailed step-by-step proof can be found in the appendix.
 ∎
@@
 $\mathcal{N}=\mathbb{C}{}$ is clearly unbounded.
-Thus, assume $\partial\mathcal{N}\not=\emptyset$.
+Thus, assume $\partial\mathcal{N}\not$$\emptyset$.
 Additionally, assume w.l.o.g.
@@
 By Lemma 2.6, $\Gamma$ cannot be periodic.
-Hence, by [10, Theorem 5.1], there exist two equilibria $a_{+},a_{-}\in\partial\mathcal{N}\cap F^{-1}(\{0\})$ such that $w_{\pm}(\Gamma)=\{a_{\pm}\}$, i.e.
+Hence, by Theorem 1.1, there exist two equilibria $a_{+},a_{-}\in\partial\mathcal{N}\cap F^{-1}(\{0\})$ such that $w_{\pm}(\Gamma)=\{a_{\pm}\}$, i.e.
 $\Gamma$ is either heteroclinic or homoclinic.
@@
 Choose $z\in\Gamma\subset\mathcal{S}_{0}\cap\partial\mathcal{N}$.
-Since $\mathcal{S}_{0}$ is open, cf.
-Theorem 3.6, there exists $\varepsilon>0$ such that $\mathcal{B}_{\varepsilon}(z)\subset\mathcal{S}_{0}$.
-Thus, analogous to the proof of Lemma 4.4, we can choose a point $y\in\mathcal{S}_{0}\cap\mathcal{N}\not=\emptyset$, i.e.
-$\Gamma(y)\subset\mathcal{N}$ is a homoclinic orbit.
-This is a contradiction.
+By openness of $\mathcal{S}_{0}$, cf.
+Theorem 3.6, we get $\mathcal{S}_{0}\cap\mathcal{N}\not$$\emptyset$, contradicting the fact that all orbits in $\mathcal{N}$ tend to $a\not\in\{a_{+},a_{-}\}$.
 Suppose, $\Gamma$ is heteroclinic, i.e.
 $a_{+}\not=a_{-}$.
-Since $a_{+},a_{-}\in\partial\mathcal{N}$ and $a\in\mathring{\mathcal{N}}$, cf.
-Theorem 4.3, it holds $a\not\in\{a_{+},a_{-}\}$.
-Thus, $a_{+}$ cannot be a node or focus, since its basin would have nonempty intersection with $\mathcal{N}$ [^10] If we investigated the case that $a$ is unstable, we would choose $a_{-}$ instead of $a_{+}$ at this point.
-Note that an orbit in $\mathcal{N}$ cannot have $a$ as well as $a_{+}$ in its limit set for $t\to\infty$.
+Since $\not\in\{a_{+},a_{-}\}, a+a_{+}$ cannot be a node or focus, since its basin would have nonempty intersection with $\mathcal{N}$.
+[^6] If we investigated the case that $a$ is unstable, we would choose $a_{-}$ instead of $a_{+}$ at this point.
 Additionally, by Lemma 2.6, $a_{+}\in\partial\mathcal{N}$ cannot be a center.
 It follows that $a_{+}$ must have order $m\geq 2$.
-By [10, Proposition 4.3], there exist $\theta\in\mathcal{E}(F,m)$ and $r,\delta>0$ such that $\Gamma$ as well as the orbit $\Gamma(x_{0})$ through $x_{0}\in\{x\in\mathbb{R}^{2}:|x-a_{+}|<r,|\arg(x-a_{+})-\theta|<\delta\}=:A$ tends to $a_{+}$ for $t\to\infty$.
-Choose $\hat{z}\in\Gamma\cap A\subset\partial\mathcal{N}\cap A$.
-Since $A$ is open, we can choose $\hat{\varepsilon}>0$ such that $\mathcal{B}_{\hat{\varepsilon}}(\hat{z})\subset A$.
-Thus, again analogous to the proof of Lemma 4.4, we can choose $x_{0}\in A\cap\partial\mathcal{N}\not=\emptyset$, i.e.
+By [19, Proposition 4.3], there exist $\theta\in\mathcal{E}(F,m)$ and $r,\delta>0$ such that $\Gamma$ as well as the orbit $\Gamma(x_{0})$ through $x_{0}\in\{x\in\mathbb{R}^{2}:|x-a_{+}|<r,|\arg(x-a_{+})-\theta|<\delta\}=:A$ tend to $a_{+}$ for $t\to\infty$.
+We have $\Gamma\cap A\not$$\emptyset$.
+Since $A$ is open, we can choose $x_{0}\in A\cap\mathcal{N}\not$$\emptyset$, i.e.
 $\Gamma(x_{0})\subset\mathcal{N}$ tends to $a_{+}\not=a$ for $t\to\infty$.
@@
 All in all, it follows that $\partial\mathcal{N}$ consists only of equilibria (but no centers) and unbounded orbits.
-Since $\partial\mathcal{N}\not=\emptyset$, we can apply Proposition 4.5 to conclude that there exists at least one unbounded orbit lying on the boundary of $a$.
+Since $\partial\mathcal{N}\not$$\emptyset$, we can apply Proposition 4.4 to conclude that there exists at least one unbounded orbit lying on the boundary of $a$.
 Thus, also $\mathcal{N}$ is unbounded.
@@
 Proof.
-Let $L\subset\mathcal{N}$ be a loop.
-Since $L$ is compact and $\mathcal{N}$ is open, cf.
-Theorem 4.3, there exist points $y_{1},\ldots,y_{\ell}\in L$ and radii $r_{1},\ldots,r_{\ell}>0$, $\ell\in\mathbb{N}{}$, such that
-$L\subset\bigcup_{k=1}^{\ell}\mathcal{B}_{r_{k}}(y_{k})=:O$
-and $\overline{O}\subset\mathcal{N}$.
-The latter property can be ensured by reducing the radii.
-In addition, there exists a closed Jordan curve $J\subset\partial O\subset\mathcal{N}$ such that $L\subset\operatorname{Int}(J)$.
-If $\operatorname{Int}(J)\setminus\mathcal{N}\not=\emptyset$, we can choose $x\in\partial\mathcal{N}\cap\operatorname{Int}(J)\not=\emptyset$.
-[^11] A detailed topological argumentation for this can be found in the proof of [9, Theorem 5.10].
+Let $J\subset\mathcal{N}$ be a loop and suppose that $\operatorname{Int}(J)\setminus\mathcal{N}\not$$\emptyset$.
+Then we can choose $x\in\partial\mathcal{N}\cap\operatorname{Int}(J)\not$$\emptyset$.
 If $x$ is not an equilibrium, $\Gamma(x)$ must be unbounded, cf.
 Theorem 4.5.
-If $x$ is an equilibrium, there exists an orbit $\Gamma\subset\partial\mathcal{N}$ with $x$ in its limit set, cf.
+If $x$ is an equilibrium, there exists an orbit $\Gamma\subset\partial\mathcal{N}$ with $x$ in its limit sets, cf.
 Proposition 4.4.
-Thus, by openness of $\operatorname{Int}(J)$, we have $\Gamma\cap\operatorname{Int}(J)\not=\emptyset$.
-In both cases we get an unbounded orbit with nonempty intersection with $\operatorname{Int}(J)$.
-But $\operatorname{Int}(J)$ is bounded, i.e.
-this is a contradiction.
-Thus, we have $\operatorname{Int}(J)\subset\mathcal{N}$.
-Since $\operatorname{Int}(J)$ is simply connected, $L$ can be continuously deformed to a point.
-Since $L$ is arbitrary and $\mathcal{N}$ is already path-connected, it follows that $\mathcal{N}$ is even simply connected.
+In both cases we get an unbounded orbit on the boundary of $a$ having nonempty intersection with $\operatorname{Int}(J)$.
+Since $\operatorname{Int}(J)$ is bounded, this orbit must intersect $J$.
+This is a contradiction to the openness of $\mathcal{N}$.
+Thus, we conclude $\operatorname{Int}(J)\subset\mathcal{N}$ and $J$ can be continuously deformed to a point in $\mathcal{N}$.
+Since $J$ is arbitrary and $\mathcal{N}$ is already path-connected, it follows that $\mathcal{N}$ is even simply connected.
 ∎
+To put the basin $\mathcal{N}$ into the framework of the theory of canonical regions and global transversals, we consider the following example of a polynomial vector field.
+Example 4.7.
+We consider the entire vector field $F:\mathbb{C}{}\to\mathbb{C}{}$ with
+$F(x):=x^{3}-x=x(x-1)(x+1)\text{.}$
+Figure 2. Local phase portrait of system (1.1) with $F(x)=x^{3}-x$, plotted with Wolfram Mathematica.
+All equilibria $F^{-1}(\{0\})=\{0,1,-1\}$ are nodes.
+The nodes at $\pm 1$ are repelling, while the node at $a=0$ is attracting.
+By using the notation $\mathrm{i}\mathbb{R}:=\{x\in\mathbb{C}:\Re(x)=0\}$, we obtain the following two properties of the phase space by a straightforward calculation:
+  • $F(\alpha)\in\mathbb{R}$ for all $\alpha\in\mathbb{R}$, i.e. $\mathbb{R}$ is flow-invariant.
+  • $F(\mathrm{i}\beta)\in\mathrm{i}\mathbb{R}$ for all $\beta\in\mathbb{R}$, i.e. $\mathrm{i}\mathbb{R}$ is flow-invariant.
+Thus, (1,0)(-1,0) as well as $(0,1)$ are orbits with distinct negative limit sets ($\{1\}$ and $\{-1\}$, respectively).
+Moreover, the upper (denoted $A_{1}$) and lower (denoted $A_{2}$) parts of the imaginary axis are orbits, both having empty negative limit sets.
+All four of these orbits lie in the basin of the node at $a=0$.
+We therefore conclude that these orbits cannot belong to the same canonical region, cf.
+[9, Chapter II.3].
+In particular, since the orbits through arbitrarily close neighborhoods of $A_{1}$ and $A_{2}$ have nonempty negative limit sets, $A_{1}$ and $A_{2}$ must be separatrices in the sense of [9].
+A similar observation applies to foci if we consider the rotated vector field $\hat{F}(x):=(1+\mathrm{i})F(x)$, $x\in\mathbb{C}$.
+Example 4.7 demonstrates that the basins of nodes and foci do not necessarily have to form canonical regions.
+Applying Proposition 4.4, we obtain the following result: if $\mathcal{N}$ is a canonical region, then there cannot be any equilibria on $\partial\mathcal{N}$.
+In general, nodal and spiral canonical regions can be constructed with the simple linear vector field $x\mapsto\lambda x$ with $\Re(\lambda)\not=0$, cf. [19, Definition 3.1].
 5. Heteroclinic regions
-In this chapter we consider an entire function $F\in\mathcal{O}{}(\mathbb{C}{})$, $F\not\equiv 0$, and at least two equilibria $a_{+},a_{-}\in\mathbb{C}{}$ of (1.1) with at least one heteroclinic orbit connecting these two equilibria.
-In particular, we assume that $a_{\pm}$ is not a center.
+In this chapter we consider an entire function $F\in\mathcal{O}{}(\mathbb{C}{})$, $F\not\equiv 0$, with two equilibria $a_{+},a_{-}\in\mathbb{C}{}$ of (1.1) and at least one heteroclinic orbit connecting these two equilibria.
+By the holomorphy of $F$, the two equilibria have to be nodes, foci or multiple equilibria.
 Under these assumptions, it is possible to investigate topologically another characteristic set, the so-called "heteroclinic region".
 Taking up the questions at the end of chapter 3, a heteroclinic region is one possible structure between two adjacent global elliptic sectors.
-We analyse it topologically and illustrate it by an example.
+We verify that heteroclinic regions are strip canonical regions in the sense of Markus and Neumann, cf.
+[9, 11].
+5.1. Geometry of heteroclinic regions
 Definition 5.1.
@@
 with
-+():=$\mathcal{H}_{+(-)}:=\left\{x\in\mathbb{C}{}:w_{+(-)}(\Gamma(x))=\{a_{+(-)}\}\right\}\text{.}$
+$\mathcal{H}_{+(-)}:=\left\{x\in\mathbb{C}{}:w_{+(-)}(\Gamma(x))=\{a_{+(-)}\}\right\}\text{.}$
 Proposition 5.2.
@@
 Both, $a_{+}$ as well $a_{-}$ is either a focus or node or an equilibrium of order $m\in\mathbb{N}\setminus\{1\}$.
-Let $x\in A_{+}$ and assume that $a_{+}$ is an equilibrium of order $m\geq 2$.
-By [10, Proposition 4.3], there exist numbers $r,\theta,\Theta,\tau>0$ such that for all $z\in\{\hat{z}\in\mathbb{C}{}:|\hat{z}-a_{+}|<r,|\arg(\hat{z}-a_{+})-\theta|<\Theta\}=:A$ the orbit through $z$ tends to $a$ for $t\to\infty$ and $M:=\Phi(\tau,x)\in A$.
+Let $x\in\mathcal{H}_{+}$ and assume that $a_{+}$ is an equilibrium of order $m\geq 2$.
+By [19, Proposition 4.3], there exist numbers $r,\theta,\Theta,\tau>0$ such that for all $z\in\{\hat{z}\in\mathbb{C}{}:|\hat{z}-a_{+}|<r,|\arg(\hat{z}-a_{+})-\theta|<\Theta\}=:A$ the orbit through $z$ tends to $a$ for $t\to\infty$ and $M:=\Phi(\tau,x)\in A$.
 Since $A$ is open, we can choose $\varepsilon>0$ small enough such that $\mathcal{B}_{\varepsilon}(M)\subset A$.
 By continuity of the flow, cf.
-[19, Chapter 2.4, Theorem 4], there exists $\delta>0$ such that $|\Phi(\tau,x)-\Phi(\tau,y)|<\varepsilon$ for all $y\in\mathcal{B}_{\delta}(x)$, i.e.
+[29, Chapter 2.4, Theorem 4], there exists $\delta>0$ such that $|\Phi(\tau,x)-\Phi(\tau,y)|<\varepsilon$ for all $y\in\mathcal{B}_{\delta}(x)$, i.e.
 $\mathcal{B}_{\delta}(x)\subset\mathcal{H}_{+}$ and $\mathcal{H}_{+}$ is open.
@@
 ∎
-Lemma 5.4.
-The basins of all centers, foci and nodes $a\in F^{-1}(\{0\})\setminus\{a_{+},a_{-}\}$ have empty intersection with $\mathcal{H}$.
-In particular, there are no centers, nodes and foci on $\partial\mathcal{H}$.
-Proof.
-There are no periodic orbits in $\mathcal{H}$. Additionally, $a\not\in\{a_{+},a_{-}\}$. Thus, this proof is analogous to the proof of Lemma 4.4. ∎
 Proposition 5.4.
@@
 Suppose there exists an equilibrium $a\in\partial\mathcal{H}\setminus\{a_{+},a_{-}\}$.
-By Lemma 5.4, $a$ is an equilibrium of order $m\geq 2$, i.e.
-$a$ has a FED, cf.
-[10, Theorem 4.4].
-In particular, by [10, Definition 4.1], there exists $r>0$ such that $a\in w_{+}(\Gamma(x))\cup w_{-}(\Gamma(x))$ for all $x\in\mathcal{B}_{r}(a)$.
-We get $\mathcal{B}_{r}(a)\cap\mathcal{H}\not=\emptyset$, which is also impossible.
+By Lemma 2.6, $a$ cannot be a center.
+Additionally, since all orbits in $\mathcal{H}$ connect $a_{+}$ to $_{-}, aa$ cannot be a node a focus.
+Hence, by using the holomorphy of $F$, we can apply [19, Theorem 4.4] to conclude that $a$ has a FED.
+In particular, there exists $r>0$ such that $a\in w_{+}(\Gamma(x))\cup w_{-}(\Gamma(x))$ for all $x\in\mathcal{B}_{r}(a)$.
+We get $\mathcal{B}_{r}(a)\cap\mathcal{H}\not$$\emptyset$, which is also impossible.
 Thus, we conclude $\partial\mathcal{H}\cap F^{-1}(\{0\})=\{a_{+},a_{-}\}$.
@@
 Theorem 5.5.
-All orbits in $\partial\mathcal{H}\setminus\{a_{+},a_{-}\}$ are unbounded. $\mathcal{H}$ is unbounded.
+All orbits on $\partial\mathcal{H}\setminus\{a_{+},a_{-}\}$ are unbounded. $\mathcal{H}$ is unbounded.
 Proof.
-If =$\partial\mathcal{H}=\{a_{+},a_{-}\}$, then $\mathcal{H}\not=\emptyset$ is open and closed with respect to the subspace topology in $\mathbb{C}{}\setminus\{a_{+},a_{-}\}$.
-[^12] This argument is similar to the beginning of the proof of Proposition 4.5.
-Since $\mathbb{C}{}\setminus\{a_{+},a_{-}\}$ is connected, =$\mathcal{H}=\mathbb{C}{}\setminus\{a_{+},a_{-}\}$ is clearly unbounded.
-Thus, assume $\partial\mathcal{H}\setminus\{a_{+},a_{-}\}\not=\emptyset$.
+If $\partial\mathcal{H}=\{a_{+},a_{-}\}$, then $\mathcal{H}\not$$\emptyset$ is open and closed in the connected set $\mathbb{C}{}\setminus\{a_{+},a_{-}\}$, i.e.
+$\mathcal{H}=\mathbb{C}{}\setminus\{a_{+},a_{-}\}$ is clearly unbounded.
+Thus, we assume $\partial\mathcal{H}\setminus\{a_{+},a_{-}\}\not$$\emptyset$.
 Suppose there exists a bounded orbit $\Gamma\subset\partial\mathcal{H}\setminus\{a_{+},a_{-}\}$.
-As in the proof of Theorem 2.6, $K:=\overline{\Gamma}$ is compact and $\Gamma\cup w_{\pm}(\Gamma)\subset K\subset\partial\mathcal{H}$.
 By Lemma 2.6, $\Gamma$ cannot be periodic.
-Hence, by Proposition 5.5 and [10, Theorem 5.1], we must have $w_{+}(\Gamma)\cup w_{-}(\Gamma)=\{a_{+},a_{-}\}$.
-Using this result, we want to derive a contradiction.
-If $w_{+}(\Gamma)=\{a_{+}\}$ and $w_{-}(\Gamma)=\{a_{-}\}$, then we get $\Gamma\subset\mathcal{H}$, which is contraction to Theorem 5.3.
-If $w_{+}(\Gamma)=\{a_{-}\}$ and $w_{-}(\Gamma)=\{a_{+}\}$, then there exists another heteroclinic region $\tilde{\mathcal{H}}$ between $a_{-}$ and $a_{+}$.
-By Theorem 5.3, $\tilde{\mathcal{H}}$ is open, i.e.
-$\tilde{\mathcal{H}}\cap\mathcal{H}\not=\emptyset$, which is again a contradiction.
-Thus, $\Gamma$ must be homoclinic.
-Assume w.l.o.g.
-that $\Gamma$ is homoclinic in $a_{+}$, i.e.
-$a_{+}$ is an equilibrium of order $m\geq 2$.
-By Corollary 3.11, there exist two adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ such that $\Gamma(x)$ is a sector-forming orbit in $a_{+}$ with these definite directions.
-By Theorem 3.6, we get $\mathcal{S}(\theta_{+},\theta_{-},\Gamma(x))\cap\mathcal{H}\not=\emptyset$, i.e.
-there exists a homoclinic orbit in $\mathcal{H}$.
-This is a contradiction.
+Hence, by Proposition 5.4 and Theorem 1.1, we must have $w_{+}(\Gamma)\cup w_{-}(\Gamma)\subset\{a_{+},a_{-}\}$.
+Additionally, by openness of $\mathcal{H}$, $\Gamma$ cannot be heteroclinic.
+It follows that $\Gamma$ must be homoclinic, w.l.o.g.
+in $a_{+}$, i.e.
+$w_{+}(\Gamma)=w_{-}(\Gamma)=\{a_{+}\}$ and $a_{+}$ is an equilibrium of order $m\geq 2$.
+We choose a point $x\in\Gamma\subset\partial\mathcal{H}\setminus\{a_{+},a_{-}\}$.
+By Lemma 3.5, there exists $\delta>0$ such that all orbits $\hat{\Gamma}$ through $\mathcal{B}_{\delta}(x)$ are also homoclinic in $a_{+}$.
+Choosing $\hat{\Gamma}\in\mathcal{H}$ leads to a contradiction.
 We conclude that $\Gamma$ cannot be bounded.
@@
 Theorem 5.3.
-Then there holds $U\cap V=\emptyset$ as well as, by Theorem 5.3, $U\cap\operatorname{Int}(J)\not=\emptyset$ and $V\cap\operatorname{Int}(J)\not=\emptyset$.
-Since $\operatorname{Int}(J)$ is connected, we get $U\cup V\not=\operatorname{Int}(J)$ and thus there exists $x\in\partial U\cap\operatorname{Int}(J)\not=\emptyset$.
-By Theorem 5.6, $\Gamma(x)\subset\partial\mathcal{H}_{1}\cap\operatorname{Int}(J)$ is unbounded.
-This is a contradiction.
+Then there holds $U\cap V=\emptyset$ as well as, by Theorem 5.3, $U\cap\operatorname{Int}(J)\not$$\emptyset$ and $V\cap\operatorname{Int}(J)\not$$\emptyset$.
+Since $\operatorname{Int}(J)$ is connected, we get $U\cup V\not=\operatorname{Int}(J)$ and thus there exists $x\in\partial U\cap\operatorname{Int}(J)\not$$\emptyset$.
+By Theorem 5.5, $\Gamma(x)\subset\partial\mathcal{H}_{1}\cap\operatorname{Int}(J)$ is unbounded, which is a contradiction.
 ∎
-Theorem 5.8.
-$\mathcal{H}$ is connected, path-connected and simply connected.
-Proof.
-For $x\in\mathcal{H}$ define $A_{x}:=\Gamma(x)\cup\{a_{+}\}$ and the cover $C:=\{A_{x}:x\in\mathcal{H}\}$ of $\mathcal{H}$.
-Since $a_{+}$ lies on the boundary of each orbit in $\mathcal{H}$, every set in $C$ is connected, cf.
-[18, Theorem 23.4].
-Moreover, for $A,\tilde{A}\in C$ it always holds $A\cap\tilde{A}\supset\{a_{+}\}\not=\emptyset$.
-Thus, by [18, Theorem 23.3], $\mathcal{H}$ is connected.
-By [23, Proposition 12.25], $\mathcal{H}$ is even path-connected.
-Let $L\subset\mathcal{H}$ be a loop.
-Analogous to the beginning of the proof of Theorem 4.7, we find a closed Jordan curve $J\subset\mathcal{H}$ such that $L\subset\operatorname{Int}(J)$.
-Suppose that $\operatorname{Int}(J)\setminus\mathcal{H}\not=\emptyset$, i.e.
-$\partial\mathcal{H}\cap\operatorname{Int}(J)\not=\emptyset$.
-Let $y\in\partial\mathcal{H}\cap\operatorname{Int}(J)$.
-By Proposition 5.2, we have $\Gamma(y)\subset\partial\mathcal{H}$.
-Since $\mathcal{H}$ is open, cf.
-Theorem 5.3, we get $\Gamma(y)\cap J=\emptyset$, i.e.
-$\Gamma(y)\subset\operatorname{Int}(J)$.
-By Proposition 5.5 and Theorem 5.6, $\Gamma(y)=\{y\}$ must be an equilibrium and thus $y\in\{a_{+},a_{-}\}$.
-If $y$ is an equilibrium of order $m\geq 2$, we can apply Corollary 3.11 to conclude the existence of an global elliptic sector in $y$ lying completely in $\operatorname{Int}(J)$, which is a contradiction to Theorem 3.7.
-Hence, by Lemma 5.4, $y$ must be a node or focus.
-If the second equilibrium $a\in\{a_{+},a_{-}\}\setminus\{y\}$ is also a node or focus, we can apply Proposition 4.5 and Theorem 4.6 to conclude the existence of an unbounded orbit $\Gamma_{0}\subset\partial\mathcal{N}$ with $y\in w_{+}(\Gamma_{0})\cup w_{-}(\Gamma_{0})$, where $\mathcal{N}$ is the basin in $a$, cf.
-Definition 4.1.
-Thus, $\Gamma_{0}\cap\mathcal{H}\supset\Gamma_{0}\cap J\not=\emptyset$, i.e.
-there exists an unbounded orbit in $\mathcal{H}$, which is impossible.
-It follows that $a$ must be an equilibrium of order $m\geq 2$ and therefore, as above, $a\in\operatorname{Ext}(J)$.
-To summarise, if $\operatorname{Int}(J)\setminus\mathcal{H}\not=\emptyset$, the only possible case is that one of the two equilibria, $a_{+}$ and $a_{-}$, is a node or focus being the only point in $\operatorname{Int}(J)\setminus\mathcal{H}$ and the other is an equilibrium of order $m\geq 2$ in $\operatorname{Ext}(J)$.
-To fix ideas, assume w.l.o.g.
-$a_{+}\in\operatorname{Int}(J)$, $a_{-}\in\operatorname{Ext}(J)$ and $\operatorname{Int}(J)\setminus\mathcal{H}=\{a_{+}\}$.
-Using this and a similar construction as in the proof of Proposition 4.5, we want to derive a contradiction.
-First, by [10, Theorem 4.4] and [10, Definition 4.1 c) (iii)], there exists a closed piecewise continuously differentiable Jordan curve $\Gamma$ that is nowhere tangential to $F$ and satisfies $\operatorname{Int}(\Gamma)\cap F^{-1}(\{0\})=\{a_{-}\}$.
-Moreover, all non-elliptic orbits in $a_{-}$, i.e.
-all orbits not lying in a global elliptic sector in $a_{-}$, cross $\Gamma$ exactly once.
-Hence, all orbits in $\mathcal{H}$ cross $C_{1}:=\Gamma\cap\mathcal{H}$ exactly once.
-We equip $C_{1}$ with the subspace topology.
-Using this construction, the map $\Psi:\mathcal{H}\to\mathbb{R}$ is well-defined, when $\Psi(x)$ is the transit time131313cf.
-[4, Definition 3.2].
-the point $x\in\mathcal{H}$ needs to $C_{1}$ along the path $\Gamma(x)$, i.e.
-$\{\Phi(\Psi(x),x)\}=\Gamma(x)\cap C_{1}$.
-Secondly, choose $\rho>0$ such that $\mathcal{B}_{\rho}(a_{+})\subset\operatorname{Int}(J)$.
+This illustrates again that holomorphy severely constrains the possible geometric configurations of the phase space.
+In fact, Corollary 5.6 does not hold if $F$ is not holomorphic.
+This happens in particular when one of the equilibria in Corollary 5.6 is a saddle, cf.
+e.g.
+[13, Chapter 2].
+Having established all of the above topological properties, we can now assemble them to reveal the full geometric structure of the heteroclinic region.
+The following argumentation shows the existence of a global transversal for a heteroclinic region.
+  • (I)
+    By using the local FED geometry in a multiple equilibrium, we deduce the existence of a transversal $T$ connecting two adjacent global elliptic sectors that have no common boundary near the multiple equilibrium, cf.
+[19, Definition 4.1] and Figure 1.
+Every orbit intersecting $T$, crosses it exactly once.
+Assume that a heteroclinic region $\mathcal{H}$ lies between these two sectors, and define the set $A=\mathcal{H}\cap T$.
+Let $q_{1},q_{2}\in A$, and consider the closed Jordan-curve $J$ formed by $\Gamma(q_{1})$, $\Gamma(q_{2})$, $a_{+}$ and $a_{-}$.
+Connecting $q_{1}$ and $a_{2}$ along the segment of $T$, one may ask whether this segment could leave $\mathcal{H}$?
+The answer is no.
+In fact, since the segment lies entirely in the interior of $J$, this would contradict Corollary 5.6 (with $p=2$).
+It follows that all points in $A$ are connected along the transversal $T$.
+Taking now the longest possible segment of the transversal $T$ contained in $\mathcal{H}$, we obtain an open smooth arc crossing each closed orbit in $\mathcal{H}$ exactly once.
+This is illustrated in Figure 3.
+  • (II)
+    Let $a$ be a focus or node on the boundary of a heteroclinic region $\mathcal{H}$.
 By applying the theory of circles without contact around nodes and foci, cf.
 [2, §3, 10.-14., §7, 1.-2.
-and §18, Lemma 3], we find a continuously differentiable closed path $C_{2}\subset\mathcal{B}_{\rho}(a)\subset\mathcal{H}$ that is nowhere tangential to $F$ and satisfies $\operatorname{Int}(C_{2})\cap F^{-1}(\{0\})=\{a_{+}\}$.
-Moreover, every orbit in $\mathcal{H}$ crosses $C_{2}$ exactly once, cf.
-[2, §3, 10., Figure 54].
-We equip $C_{2}$ with the subspace topology and define the map $\eta:C_{2}\to C_{1}$ by $\eta(x):=\Phi(\Psi(x),x)$.
-Analogous to the proof of Proposition 4.5, $\eta$ is continuous and bijective.
-Additionally, $C_{1}$ is Hausdorff and $C_{2}$ is compact, i.e.
-$\eta$ is a homeomorphism, cf.
-[18, Theorem 26.6].
-Furthermore, $C_{2}$ is isomorphic to $S^{1}$.
-By [18, Corollary 52.5] and [18, Theorem 54.5], it follows that the fundamental group of $C_{1}$ is isomorphic to the additive group $\mathbb{Z}$.
-By [10, Definition 4.1 c) (ii)], there exists a homoclinic orbit having nonempty intersection with $\Gamma$.
-Thus, this homoclinic orbit cannot lie in $\mathcal{H}$.
-Therefore, it follows that $C_{1}$ cannot be closed, i.e.
-the fundamental group of $C_{1}$ cannot be isomorphic to $\mathbb{Z}$.
-More precisely, $C_{1}$ is connected and thus homeomorphic to the simply-connected space $(0,1)\subset\mathbb{R}$, cf.
-[18, Theorem 23.5].
-This is a contradiction.
-All in all, we conclude $\operatorname{Int}(J)\subset\mathcal{H}$.
-Since $\operatorname{Int}(J)$ is simply connected, $L$ can be continuously deformed to a point.
-Since $L$ is arbitrary and $\mathcal{H}$ is already path-connected, it follows that $\mathcal{H}$ is even simply connected.
+and §18, Lemma 3], we find always a continuously differentiable closed path $C\subset\mathcal{H}$ being nowhere tangential to $F$ and satisfying $\operatorname{Int}(C)\cap F^{-1}(\{0\})=\{a\}$.
+All orbits in the basin of $a$ cross $C$ exactly once.
+This is illustrated in Figure 4.
+  • (III)
+    If $a_{+}$ or $a_{-}$ is a multiple equilibrium, we can apply (I) to deduce the existence of global transversal.
+If both, $a_{+}$ and $a_{-}$ are foci/nodes, then we can apply Proposition 4.4 to show the existence of at least one unbounded orbit $\Gamma$ with $a_{+}$ in its limit set.
+Since this orbit intersects $C$ at exactly one point, cf.
+(II), $C\cap\mathcal{H}$ cannot be a closed segment of $C$, cf.
+Figure 4.
+Analogous to the argument in (I), $C\cap\mathcal{H}$ has to be path-connected and is thus a global transversal for $\mathcal{H}$.
+Figure 3. A complete transversal (blue) for $\mathcal{H}$ near a multiple equilibrium (green). Separatrices are colored red.
+Figure 4. A complete transversal (blue) for $\mathcal{H}$ near a node (green). The unbounded separatrix is colored red, cf. step (III).
+Theorem 5.7.
+$\mathcal{H}$ is connected, path-connected and simply connected. $\mathcal{H}$ is a strip canonical region.
+Proof.
+We use our above considerations in (I)–(III).
+Let $T:[0,1]\to\mathcal{H}$ be the smooth global transversal through $\mathcal{H}$ and define $Q:=\mathbb{R}\times[0,1]\subset\mathbb{R}^{2}$.
+Then the map $\mu:Q\to\mathcal{H}$, $\mu(s,t):=\Phi(t,T(s))$ is a homeomorphism.
+Moreover, by openness of $\mathcal{H}$, there are no separatrices in $\mathcal{H}$.
+Since $Q$ is simply connected, this theorem is a direct consequence of [10, Theorem I] and [9, Theorem 5.1].
 ∎
-In [3, 4, 10] many examples are illustrating topological properties of the basins of equilibria and global elliptic sectors.
-[10, Example 4.8, Figure 3] demonstrates a heteroclinic region between the two equilibria.
-The following example shows that there can be more than one heteroclinic region between two global elliptic sectors.
+This leads to an interesting observation: global elliptic sectors and heteroclinic regions are both strip regions and thus topologically equivalent in the sense of [9].
+In order to distinguish these two sets we need information about the boundary orbits (separatrices).
+This underlines that the behavior of separatrices is crucial for determining the global phase portrait of (1.1).
+5.2. Illustrating example
+Several examples of basins, sectors, and regions can be found in [4, 5, 19].
+Another example has already been discussed, cf.
 Example 4.7.
-Consider the vector field $F:\mathbb{C}{}\to\mathbb{C}{}$ with
+The vector field $F(x)=x^{3}-x$ has two heteroclinic regions: the left and right open half-planes of $\mathbb{C}{}\setminus\{x\in\mathbb{R}:|x|>1\}$.
+These heteroclinic regions share $\mathrm{i}\mathbb{R}$ (one equilibrium and two separatrices) as their common boundary.
+We now turn to a more complicated example to illustrate the theory developed in the preceding chapters.
+Example 5.8.
+We consider the entire vector field $F:\mathbb{C}{}\to\mathbb{C}{}$ with
 $F(x):=x^{2}(x-1)(x-\mathrm{i})(x-1-\mathrm{i})\text{.}$
-We have $F\in\mathcal{O}(\mathbb{C}{})$ and $F^{-1}(\{0\})=\{0,1,\mathrm{i},1+\mathrm{i}\}$.
-Moreover, $a_{1}=0$ is an equilibrium of order $2$, $a_{2}=1+\mathrm{i}$ is a stable node and $a_{3}=1$ as well as $a_{4}=\mathrm{i}$ are stable foci.
 Figure 5. Local phase portrait of system (1.1) with $F(x)=x^{2}(x-1)(x-\mathrm{i})(x-1-\mathrm{i})$, plotted with Matlab.
-We observe three heteroclinic regions between the two global elliptic sectors: the first between $a_{1}$ and $a_{2}$, the second between $a_{1}$ and $a_{3}$, and the third between $a_{1}$ and $a_{4}$.
+We have $F^{-1}(\{0\})=\{0,1,\mathrm{i},1+\mathrm{i}\}$ (colored green).
+In particular, $a_{1}=0$ is an equilibrium of order $2$, $a_{2}=1+\mathrm{i}$ is an attracting node and $a_{3}=1$ as well as $a_{4}=\mathrm{i}$ are attracting foci.
+We observe three heteroclinic regions (colored blue) lying between the two global elliptic sectors: the first between $a_{1}$ and $a_{2}$, the second between $a_{1}$ and $a_{3}$, and the third between $a_{1}$ and $a_{4}$.
+The red trajectories are the separatrices.
+In sum, we see five strip canonical regions.
 We can verify all the topological properties proven in this chapter as well as Corollary 5.6.
@@
     All heteroclinic orbits of (1.1) tend to $a$ in the same definite direction.
+Appendix A
+Proof of Proposition 3.8.
+Let $x\in\mathcal{S}$.
+Suppose there exists $z_{1}\in\operatorname{Int}(\Gamma(x)\cup\{a\})$ such that $z_{1}\not\in\mathcal{S}$.
+Then there exists $z_{2}\in\partial\mathcal{S}\cap\operatorname{Int}(\Gamma(x)\cup\{a\})\not$$\emptyset$.
+Since the case $a=z_{2}$ is not possible, $\Gamma(z_{2})$ has to be unbounded, cf.
+Theorem 3.7.
+But $\Gamma(z_{2})$ lies in the bounded set $\operatorname{Int}(\Gamma(x)\cup\{a\})$, which is a contradiction.
+Hence we have $\operatorname{Int}(\Gamma(x)\cup\{a\})\subset\mathcal{S}$.
+Additionally, by applying Proposition 3.3, we get
+$\overline{\operatorname{Int}(\Gamma(x)\cup\{a\})}\setminus\{a\}=\operatorname{Int}(\Gamma(x)\cup\{a\})\cup\Gamma(x)\subset\mathcal{S}\quad\forall\,x\in\mathcal{S}\text{.}$
+By Theorem 3.6, for all $x\in\mathcal{S}$ we can choose $y_{x}\in\operatorname{Ext}(\Gamma(x)\cup\{a\})\cap\mathcal{S}$ such that $\Gamma(x)\subset\operatorname{Int}(\Gamma(y_{x})\cup\{a\})$.
+It follows
+$\overline{\operatorname{Int}(\Gamma(x)\cup\{a\})}\setminus\{a\}\subset\operatorname{Int}(\Gamma(y_{x})\cup\{a\})\subset\mathcal{S}\quad\forall\,x\in\mathcal{S}\text{.}$
+These considerations show both equations in (3.1).
+Suppose there exist $y_{1},y_{2}\in\mathcal{S}$, $\Gamma(y_{1})\not=\Gamma(y_{2})$, such that $\Gamma(y_{1})$ and $\Gamma(y_{2})$ are mutually exterior.
+At first, assume that $\{y_{1},y_{2}\}\cap\operatorname{Ext}(\Xi\cup\{a\})\not$$\emptyset$.
+To fix ideas, assume w.l.o.g.
+$y_{1}\in\operatorname{Ext}(\Xi\cup\{a\})$.
+We have $\Xi\subset\operatorname{Int}(\Gamma(y_{1})\cup\{a\})$.
+Since $\Gamma(y_{2})\subset\operatorname{Ext}(\Gamma(y_{1})\cup\{a\})$, we get $y_{2}\in\operatorname{Ext}(\Xi\cup\{a\})$ and thus $\Xi\subset\operatorname{Int}(\Gamma(y_{1})\cup\{a\})\subset\operatorname{Ext}(\Gamma(y_{2})\cup\{a\})$.
+It follows the contradiction $y_{2}\not\in\mathcal{S}$.
+[…]
+The last possible case is $\{y_{1},y_{2}\}\cap\Xi\not$$\emptyset$.
+To fix ideas, we assume w.l.o.g.
+$\Xi=\Gamma(y_{1})$.
+By Definition 3.1, we must have either $y_{2}\in\operatorname{Int}(\Xi\cup\{a\})=\operatorname{Int}(\Gamma(y_{1})\cup\{a\})$ or $\Gamma(y_{1})=\Xi\subset\operatorname{Int}(\Gamma(y_{2})\cup\{a\})$.
+Both results lead to a contradiction.
+∎
+Proof of Proposition 4.4.
+Suppose there exists an isolated point $\tilde{a}\in\partial\mathcal{N}\cap F^{-1}(\{0\})$, i.e. there exists $\rho>0$ such that
+$\partial\mathcal{N}\cap(\underbrace{\mathcal{B}_{\rho}(\tilde{a})\setminus\{\tilde{a}\}}_{=:A})=\left(\mathcal{B}_{\rho}(\tilde{a})\cap\partial\mathcal{N}\right)\setminus\{\tilde{a}\}$$\emptyset\text{.}$
+It follows [^7] The simple but technical proof of $\hat{\partial}(\mathcal{N}\cap A)\subset\partial\mathcal{N}\cap A=\emptyset$, where $\hat{\partial}$ denotes the boundary with respect to the subspace topology on $A$, is omitted.
+that $\mathcal{N}\cap A\not$$\emptyset$ is open and closed with respect to the subspace topology on $A$.
+Since the punctured circle $A$ is connected, we must have $\mathcal{N}\cap A=A$.
+Hence, since $\tilde{a}\not\in\mathcal{N}$, we conclude
+(A.1) $\mathcal{N}\cap\mathcal{B}_{\rho}(\tilde{a})=\mathcal{B}_{\rho}(\tilde{a})\setminus\{\tilde{a}\}\text{.}$
+If $\tilde{a}$ was of order $m\geq 2$, then, by [19, Theorem 4.4], there would be at least one homoclinic orbit in $A$ tending to $\tilde{a}$ in both time directions, i.e.
+equation (A.1) would not be fulfilled and $\tilde{a}$ would not be isolated.
+Thus, $\tilde{a}$ must be a node or focus, cf.
+Lemma 2.6.
+Denote by $\tilde{\mathcal{N}}$ the basin of $F$ in $\tilde{a}$.
+We reduce $\rho$ such that $\mathcal{B}_{\rho}(\tilde{a})\subset\tilde{\mathcal{N}}$ and $\mathcal{B}_{\rho}(a)\subset\mathcal{N}$.
+By (A.1) and Proposition 4.2, we have $\tilde{\mathcal{N}}\subset\mathcal{N}$.
+[…]
+One the one hand, since $C_{2}$ is compact, also $\eta(C_{2})$ is compact in $C_{1}$, cf.
+[28, Theorem 26.5].
+Thus, $\eta(C_{2})$ is closed in $C_{1}$, cf.
+[28, Theorem 26.3].
+On the other hand, by Proposition 4.2, we have $\eta(C_{2})\subset\tilde{\mathcal{N}}\cap C_{1}$.
+Additionally, since all orbits through $\tilde{\mathcal{N}}\cap C_{1}\subset\tilde{\mathcal{N}}$ have to cross $C_{2}$, we also have $\tilde{\mathcal{N}}\cap C_{1}\subset\eta(C_{2})$, i.e.
+$\eta(C_{2})=\tilde{\mathcal{N}}\cap C_{1}$.
+Moreover, by Theorem 4.3, $\tilde{\mathcal{N}}\cap C_{1}$ is open in $C_{1}$.
+We conclude that $\eta(C_{2})\not$$\emptyset$ is open and closed in $C_{1}$.
+Since $C_{1}$ is connected, we must have $C_{1}=\eta(C_{2})$, i.e.
+$\eta$ is surjective.
+Let $y_{1},y_{2}\in C_{2}$ be arbitrary.
+Assume $\eta(y_{1})=\eta(y_{2})$.
+Then, by definition of $\eta$, $\Gamma(y_{1})=\Gamma(\eta(y_{1}))=\Gamma(\eta(y_{2}))=\Gamma(y_{2})$.
+Since all orbits in $\tilde{\mathcal{N}}$ cross $C_{2}$ exactly once, we must have $y_{1}=y_{2}$.
+Hence, $\eta$ is also injective.
+It follows that $\eta$ is bijective and thus a homeomorphism, cf.
+[28, Theorem 26.6].
+We conclude that every orbit in $\mathcal{N}$ and $\tilde{\mathcal{N}}$ crosses $C_{1}$ as well as $C_{2}$ exactly once, i.e.
+we get
+(A.2) $\hskip-19.91692pt\mathcal{H}:=\mathcal{N}\cap\tilde{\mathcal{N}}=\mathcal{N}\setminus\{a\}=\tilde{\mathcal{N}}\setminus\{\tilde{a}\}=\bigcup\limits_{\mathclap{x_{2}\in C_{2}}}\Gamma(x_{2})=\bigcup\limits_{\mathclap{x_{1}\in C_{1}}}\Gamma(x_{1})\text{.}$
+Since $C_{1}\subset\mathcal{H}$ is a Jordan curve, there exists a homeomorphism $h:C_{1}\to S^{1}$.
+Equip the cylinder surface $B:=\mathbb{R}\times S^{1}\subset\mathbb{R}^{3}$ with the subspace topology and define the map $\zeta:B\to\mathcal{H}$ by $\zeta(t,x):=\Phi(t,h^{-1}(x))$.
+By (A.2), we have $\zeta(B)=\mathcal{H}$, i.e.
+$\zeta$ is surjective.
+Let $(t_{1},x_{1}),(t_{2},x_{2})\in B$ be arbitrary and assume $\zeta(t_{1},x_{1})=\zeta(t_{2},x_{2})$.
+Then $\{h^{-1}(x_{1}),h^{-1}(x_{2})\}\subset C_{1}$ and
+$\Gamma(h^{-1}(x_{1}))=\Gamma(\zeta(t_{1},x_{1}))=\Gamma(\zeta(t_{2},x_{2}))=\Gamma(h^{-1}(x_{2}))\text{.}$
+Since all orbits in $\mathcal{H}$ cross $C_{1}$ exactly once, we get $h^{-1}(x_{1})=h^{-1}(x_{2})$.
+Since $h$ is bijective, we conclude $x_{1}=x_{2}$.
+This gives us $\Phi(t_{1},h^{-1}(x_{1}))=\Phi(t_{2},h^{-1}(x_{1}))$ and thus $t_{1}=t_{2}$, since $\Gamma(h^{-1}(x_{1}))$ is not periodic.
+It follows that $\zeta$ is also injective and thus bijective.
+Hence, there exists the inverse $\zeta^{-1}$ of $\zeta$, which is given by $\zeta^{-1}(x):=(-\Psi(x),h(\Phi(\Psi(x),x)))$, $x\in\mathcal{H}$.
+In fact, we easily check that the equation $\zeta(\zeta^{-1}(x))=x$ holds for all $x\in\mathcal{H}$.
+As a composition of continuous functions, $\zeta$ as well as $\zeta^{-1}$ are continuous and therefore $\zeta$ is a homeomorphism.
+By [28, Corollary 52.5], the fundamental group of $\mathcal{H}$ is isomorphic to $B$.
+Additionally, as the circle line $S^{1}\subset\mathbb{R}^{2}$ is a deformation retract of the cylinder surface $B$, we conclude that the fundamental group of $\mathcal{H}$ is isomorphic to the additive group $\mathbb{Z}$, cf.
+[28, Theorem 54.5] and [28, Theorem 58.3].
+But $\mathcal{H}$ has the two holes $a$ and $\tilde{a}$, which is impossible for a set having a fundamental group isomorphic to $\mathbb{Z}$.
+In particular, by using (A.2), we can choose the loop $C_{2}\cup\Phi([0,\Psi(z_{0})],z_{0})\cup C_{1}\subset\mathcal{H}$ with an arbitrary point $z_{0}\in C_{2}$.
+This loop has $a\not\in\mathcal{H}$ as well as $\tilde{a}\not\in\mathcal{H}$ in its interior and thus cannot be transferred to a loop on $S^{1}$.
+All in all, we derive a contradiction and $\tilde{a}$ cannot be isolated on $\partial\mathcal{N}$.
+[…]